A Complete Bibliography of Publications in the Proceedings of the American Philosophical

Society (2000–2049)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

12 June 2023 Version 1.03

Title word cross-reference

11th [Ros06a]. 1769 [BD19]. 1782 [Mul08]. 18th [Nas19]. 1916-6 [BL16]. 1988 [Rig00]. 1991 [DF04, Kot00, McE02, Pin09]. 1992 [Ben04, FP06, Hin01, Kel01, Pal03, Sud00]. 1993 [Cla00, Ram04, Tel02]. 1994 [Ack04, Bar00, RC00]. 1995 [FD06, McC00, Pea01]. 1996 [KHEP08, McK01]. 1997 [BSW06, Con00a, Dow02b, HP06, JB00, Kat00, KK03, Kor00, Lid01, Ost06, OTC00, Pea06, Sco06, SR00]. 1998 [Ada00, And00, Bor01, Bri00, Cha00, Cho00, Con00b, Coo06, Cot00, Cri00, Eag00b, Hux01, Mas00, McK00, Mis00, Nie00, Pat09, Sil04, Wha03]. 1999 [Ber00, Cha01, CT04, Cra01, Dan10, Dar01, Eng01, Fit02, Hac01, Hel05, Kol01, Kos05, Kro01, Mat01, Mon01, Moy01, Pet01, Pin01, Pro07, Pye01, Ree01, SK10, Zem02].

2000 [And03, Bla02, Bro02b, Con03, Cro02, DH07, Deg02, Dri02, Føl12, Geo07, Gle02, GG01, Gra02b, Gra02c, GP02, Hen02, Hun07, Jac03, Ken03, Lam02, Lin03, May02, McC02, MS02, Mos05, Par06, Rei02, Roe08, Sok02, Ste05, Tho02, Wal02, Wal07, Wri07, dA03]. 2001 [Abe03, Ada03, AF03, Bar03b, Bar03c, Bed03, Ben02, Bia04, Bis03, Bur03, Cha03, DT03, EC02, For03, Fra04, Gal03, Gar03, HCC06, Hop14, Kih03, Kir03, Kop06, Lan02, Osm03, Pel03, Sha03, Sil09, Sin04, Sta05]. 2002 [BDN04, Bar05, Bis05, Bla10, Bow11, Bri04, BGS^{+}08, Cas04, Coh04, Con10, Dal05, Doo04, Dot05, DDV06, Got08, Hor14, Jam04, Lan05, Ort04, Pei04, PRP^{+}04, Phi05, Pre05, Put05, Pye05, Rid04, Rud05a, Sab04, Sei05, Sli04, Sol04, Tho04, Tri05, Wol04, Wyn04]. 2003 [BIK^{+}08, Bau05, Bob07, Bom10, Bro05, Cai05, Dav12, Des05, Dys06, Ern05, Fis08, Gin05, Goh05, Gut05, HM06, Hei05, Hol04, Joh06, Kar05, Kat05a, Le 06, May09, MN11, Mit07, Mul10, Mys05, Pap05, Red08, Ris05, San06, Shu06, Sup06, Wat05, Wil05a, Woo06a]. 2004 [Arr11, Boe07, Bok06, Bra06, Bra10, Bri06, BL07, DP07, FS08, Gin10, GA06, GR06, Ham07, Har10, Jor06, Lev07, Lip06, Max07, MS07, Mil06, Mor06, Ost07, Pri07, Ros06a, Ruf06, Sal06, Sed07, Sny07, Ste06, Wu07]. 2005 [Atk08, Awo07, BF13, BK07, Beg07, Bev08, Boc07, Boy07, Bri08, Bur08, Dre07, Eag08, Erk09, Fau07, Fon07, Gou08, Gra08, GC14, LL08, Lie07, Mat07a, Mur07, Nat07, Pel06, Phi07, Ram08, Rim08, Rot09, Sei08, Ste08, Tho07, Tsi07]. 2006 [All09, Cro09, Gar08a, Ham08b, Hin08, Hod12, Jay14, Jon08, Kni08, Lak11, Mar08, Mon08, Mum08, Özy08, Shw10, Tho08, VF08, Zuc08]. 2007 [Alt09, Ber09, Bow09, CR10, Fac09, Fis09, Fox09, Gar13b, Hac09, Hun09, Mat09, Plo09, Pos12, Put09, Rai09, Rho09, RK09, Sim11, SA09, Tre09, Wal09, Wil09, Win09a, Wol11, Wyl11]. 2008 [ALSS10, Ash09, Bai10, Bar10a, Bar16, Bar10b, BG11, Bow10, Bro09, Bro10, CD12, Dys10, Fin10, FEE11, Gou11, Her17, Höl10, Mat10, MGM10, Roo15, Sli10, Ter10, Vac09, VV11]. 2009 [Agr11, Asc11, BL16, Cho12, Cul14, Cun16, Gar12, GH13, Gra11, Jam11, Kle11, Lew11, Mid11, Oki12, Par12, Rot11, Sel11, Sme11, SCL^{+}10, Ste11, Sto11, Van11, VW12]. 2010 [BR12, Bri12, DeL12b, DH12, Dun11, Dut16, Eas18, Fis12, FL13, Gol12, ME13, Nel11, Nor12, Pro12, Rid11, RB12, Slo12, Smi12, Spa12, Vog14, Yak12]. 2011 [Alb14, Aya12, Bai13, Buc13, EDJG12, GTCD13, HKT14, KG14, Kle16, MLT11, MT12, Nec12, Sak17, Sch14, Tha17, TS14]. 2012 [Bas13, Ber13, Cas14, Gar13a, HGG13, Hog14, McC14, Mor13, Pea14, PT13, PO13, SBC^{+}14, She14, Sup14, Wal13]. 2013 [Alt15, Ano13j, Bad17, Bed14, Bir15, Bla16, Fau14, Gut15, Har15, Jef16, Ken16, Mad15, Mar15a, Mar15b, Mid15, Mor16, Olo15, PP14, Rei15, Ric15, Rus16, Tip15, Tri14, Ven15]. 2014 [Ano14j, BWW16, Bro17, Cli15, Fle17, Frè16, FDBH17, Lan16, May17, Not16, Özd17, Pin16, Rak16, Rut15, Sim16, Sor15, Tom15, de 17]. 2015 [Ano15j, Cul17, CSG17, GK17, Kno17, Lig17, McC16, SG18]. 2016 [Cro18, Gar17, JH18, Tho18]. 2017 [Ano17b, Bar18b, Bow18, Hei18, Wor18]. 2018 [Ano18b]. 2019 [Ano19j]. 21st [Gra15]. 275-Year [Van18].

31 [Ano13j, Ano14j, Ano15j, Ano17b, Ano18b, Ano19j].

A. [Ada00, McK00]. Aaron [DH07]. Abeles [Wal07]. Abelson [MS07]. Abortifacients [Sch05]. Abraham [Geo07]. Abram [Des05]. Abrams [CSG17]. Absent [LB13]. Academe [Gra14, Gre10a]. Academic [Atk04]. Accelerating [Kir13]. Accessibility [Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano10c, Ano10d, Ano11a, Ano11b, Ano11c, Ano11d, Ano12a, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano13d, Ano14a, Ano14b, Ano14c, Ano14d, Ano15a, Ano15b, Ano15c, Ano15d]. Accident [Mcc05]. Account [Dir10]. Accuracy [Pre00]. Ackley [Pat09]. Acquire [Ron06]. across [Bri07]. Action [Tho13, VF10]. Acute [Lea12]. Ada [Rus16]. Adam [Gle02, Qua13]. Adams [Bar03c]. Adaptive [Jab12]. Administration [Pie03b, GBG15]. Administrative [Ano19a, Ano19b, Ano19c, Ano19d, Ano20a]. Adrian [Pea01]. Adult [Lea12]. Advanced [Sat03]. Advice [Lea15b]. Aeneid [Put01]. Affair [Hag02]. African [Sha05]. after [Gre04a, SVD11]. Aftermath [Lon11, War15a]. Age [App19, Dud01, Fuc01, Hof13, Rud05b, Var09]. Agnotology [Sch05]. Agriculture [Rut05, Wol05]. Ahmed [Tho18]. AIDS [Hil00]. Aigrain [Sei05]. Akio [Pet01]. Alan [Hux01, Bow18]. Albert [Fac09, Gil14, Rud14]. Alex [Smi12]. Alexander [Par06, Cas11, Cho12, Ham08b, Tre09]. Alexandria [Bag02]. Alfred [Cro18, Mei19, Nas13, Wil09]. Alfvén [FD06]. Alibi [Gos01]. Alien [Rei07]. Alison [McC00]. Alive [Mar18]. Allan [Lie07]. Alligators [Nic05]. Allograft [SVD11]. Allografts [Dub11]. Almond [Pye05]. Alpers [Alt15]. Alphabetic [Bri07]. Alternating [Boo10]. Alternative [Hum07]. Alvin [Zuc08]. Always [Rot12]. Ambiguity [Put01]. America [Hau20, Kai07, Mas07, O’S14, Por11, Tre19, War05, Whi00]. American [Dar19, Pow15, Ber14, Coh09, Dir10, Gre19, Hag02, Kev08, Mas16, Mat07b, Mil15, Nas19, Odo07, Pow14, Pow15, Sel05, SCG^{+}12, Sha05, Spe18, Syl14, Van07, WP15, Wol14b]. Americanist [Dar18]. Americans [Fas15]. Among [Fle15, Win01]. Ancient [Car06, LJ01, Rak19, Rei18, Win09b]. Anderson [Cul17]. André [Bor01]. Andrew [Pea14, MGM10, Sto11]. Animation [Nap05]. Anne [CD12]. Annenberg [Jam04, Jam11]. Anniversary [McK04, Pol04b, Rem06]. Ansley [Pre05]. Anthony [Bla16]. Anthrax [Gui02]. Anthropocene [Rub20]. Anthropology [Dar18, Gre10a]. Antigone [Ful06]. Antikythera [Jon18]. Anxieties [Pip17]. Aotearoa [Poc01]. Apart [Seg17]. Apiculture [Ber14]. Apoplexy [Fin15]. Appleby [JH18]. Apples [Coh11]. Applied [Coh11, Gui02]. Approach [Tra00]. April [Alt09, And03, Ash09, Awo07, BDN04, Bar00, Beg07, Ben02, Boy07, Bri12, Cai05, Cha00, Cha01, CT04, Cla00, Con00b, Cul14, CSG17, Dal05, Dav12, Deg02, Des05, Dow02b, Dys10, Fac09, FD06, Fau07, Fis09, Fra04, Gal03, Gar13b, Gle02, Got08, Gut05, Har10, Hen02, Her17, Hod12, Hog14, Hun09, JH18, Jor06, Ken16, Kir03, Kno17, Kos05, MLT11, MS07, MT12, Mor16, Nel11, Nor12, Olo15, OTC00, Pea01, Pel03, Plo09, Pro12, Rei15, Rim08, Sab04, Sha03, Sil04, Sim16, Ste05, Tom15, Tsi07, Ven15, Wha03, Wri07, Zuc08]. APS [Ano19a, Ano19b, Ano19c, Ano19d, Ano20a, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano20a, Ano20b, Bar18a, Dar18, Sei01]. Archaeology [Bas06]. Archibald [Dys10, Bok06]. Archimedes [EN10]. Architectural [Kin01]. Archival [Sel05]. Archives [Lep16, Tho17, Dar18]. Argentina [Pye02]. Argument [Kit00]. Armand [Bom10]. Armitage [She14]. Arms [Gar01]. Armstrong [HGG13, McC02]. Army [Gra15]. Arnold [Yak12]. Art [Put01, Rem06, Var03, Win19, Mit08]. Arthur [PP14, Ber09, CT04, Coo06, Fle17, Mat09, Reb11]. Arts [All00, Ano00i]. Assaying [Joh16]. Assessment [Wil17]. Asset [Whi00]. Assumption [Von11]. Asteroids [Mor10]. Astronomical [Pre00]. Astronomy [EK11, Gin11]. Asymmetry [Lep16]. Athelstan [Nie00]. Athenian [Tra00]. Atiyah [Wit20]. Atlantic [Man20, Sch05]. Atomic [Gel01, Jon18]. Auburn [Sto06]. Audacity [Sny05]. August [Abe03, All09, And03, Awo07, BIK^{+}08, Bai10, Bar10a, Ber00, Bom10, Bor01, Bow10, Bow11, Bra10, Bri06, Bri08, Bur03, Cho00, Coh04, Coo06, Cro18, Cul17, DeL12b, Doo04, Fau07, FS08, Gin10, Gou08, Gra02b, GR06, Gut05, Ham07, Ham08b, HM06, Hei18, HGG13, Hin01, Jef16, Joh06, Jon08, Kar05, Kat05a, Kle16, Lan02, ME13, MS07, Mid11, Mil06, Mor13, Moy01, Mul10, Mys05, Ost07, Özd17, Par06, PRP^{+}04, Phi05, Red08, RK09, Rim08, Ris05, Roe08, Rot09, San06, Shw10, Sim16, Sok02, Sor15, Sta05, Sup06, Sup14, Tho18, Tre09, Ven15, Wal13, Wil05a, Wyl11, Wyn04]. Augustinus [Pel03]. Austrian [Plo09]. Authentication [Sil16]. Author [Gin09]. Axelrod [Sny07].

B [Bar03c, Ben02, Ern05, Lip06, Sim16, Wal02, Wyl11]. Babcock [San06]. Babies [Sel05]. Bach [Wol01a]. Back [Ano00e, Ano00f, Ano00g, Ano00h, Ano01e, Ano01f, Ano01g, Ano01h, Ano02e, Ano02f, Ano02g, Ano02h, Ano03e, Ano03f, Ano03g, Ano03h, Ano04e, Ano04f, Ano04g, Ano04h, Ano05e, Ano05f, Ano05g, Ano05h, Ano06e, Ano06f, Ano06g, Ano06h, Ano07e, Ano07f, Ano07g, Ano07h, Ano08e, Ano08f, Ano08g, Ano08h, Ano09e, Ano09f, Ano09g, Ano09h, Ano10e, Ano10f, Ano10g, Ano10h, Ano11e, Ano11f, Ano11g, Ano11h, Ano12e, Ano12f, Ano12g, Ano12h, Ano13e, Ano13f, Ano13g, Ano13h, Ano14e, Ano14f, Ano14g, Ano14h, Ano15e, Ano15f, Ano15g, Ano15h]. Backward [Ker16]. Bacterial [Ber06]. Bad [Blo12]. Bahcall [Gou08]. Bailey [Ram08, Tha17]. Baker [Kop06, Sei08]. Baldinger [Rai09]. Baldwin [McC16]. Barbados [CHH13]. Bard [Fit02]. Bardeen [Pin09]. Bariatric [Sug11]. Barnes [Gil14, Rud14]. Barnson [Bri08]. Baron [Kor00, Pea01, Sed07]. Barr [Kni08]. Barton [Cot00]. Bartram [Nic05]. Baruch [MLT11]. Based [Can01, Sca09]. Basel [HO14]. Bases [Fla05]. Basin [Man20]. Bate [Eng01]. Battle [Tat19]. Battlefield [Mar03]. BC [Bas06]. be [Boi16]. Beagle [Key03]. Beams [Gel01]. Bearden [Fin05]. Bearn [Cho12]. Because [Gre04a]. Beck [SK10]. Bedini [Pos12]. Bee [Ber14]. Bee-Setting [Ber14]. Bees [Ber14]. before [Rot12]. Behavior [Ber06]. Behind [Seg17, Von11]. Behold [Hof13]. Being [Sat03]. Bela [Pap05]. Bellah [Tip15]. Bellow [Beg07]. Belonging [Mof20]. Belts [Bru18]. Ben [Gre19, Rem06]. Benedetto [Cla00]. Benedict [Cul17]. Benefits [Vag01]. Benjamin [Ano00i, All00, Ati06, FH08, Fin15, Mul08, Sta06, Tal06, Van18, Von11, WP15]. Benoît [War15j]. Bentley [Erk09]. Benzer [Wol11]. Bergere [VF08]. Bergson [Des05]. Bergstrom [Ham07]. Berlin [BSW06]. Bernard [Gin05, Mor13, Eas18]. Bernhard [Spa12]. Bessborough [Rei15]. Best [Ker09]. Besting [Eve05]. Bethe [Dre07]. Betrayed [Kel00]. Better [Sel05]. between [Dun07, Win19, Woo06b]. Beverly [Bow09]. Bevington [Cox19]. Beyond [Doy19, Gro05, Ree19, WH14]. Bicknell [Joh06]. Biddle [Mur07]. Bifida [Adz13]. Bing [Vog14]. Biologists [Bal12]. Biology [Kea12, Wei06]. Biomedical [Sat03, Vag01]. Bird [Gra09]. Birth [Gin11]. Black [Gly16]. Blackfoot [Miy11]. Blackmun [Gre04b]. Blackwell [Bri12]. Blast [Mil16]. Blitzer [Moy01]. Bloch [Jon08, Ken03]. Block [Van18]. Blumberg [MLT11]. Board [Les04, McK04, Pol04b, WH14]. Boas [Bak10, Dar18, Dir10, Gre10a]. Boasian [Jac10]. Boasians [Boo10]. Bob [Osm03]. Bober [Rid04]. Bodde [Le 06]. Bodian [FP06]. Bogle [NJ20]. Bogorad [HM06]. Bomb [Jon18]. Bond [Sie14]. Bonner [Gra19]. Books [Nul14]. Boorstin [Mor06]. Booth [Bev08, PT13]. Border [Mas16]. Borders [Rot12]. Borel [Bom10]. Born [Tri17]. Boundaries [De 13, Res16, Rot12]. Boundary [Reb11]. Bounded [Lo13]. Bouwsma [Bra06]. Boyd [Mar03]. Braidwood [Wat05]. Brain [Chu11]. Bram [Geo07]. Brée [Bis03]. Brentano [Bis05]. Brian [Kel01]. Bridenbaugh [Hin01]. Brief [Bor17, Ste04a]. Brion [Min20]. British [O’S14]. Britton [Dut16]. Broad [Jac17]. Broadcasting [Tho09]. Broken [Bru18]. Brooke [Bed03]. Brooks [Bra10]. Brothers [Pie03a]. Brown [Pei04, Les04, McK04, Pol04b, WH14]. Bruce [Pal03, Wal09]. Brugt [Lam02]. Bruner [Gar17]. Bruno [Cla00]. Bryan [Bro17]. Bryce [KG14]. Bryozoans [Key03]. Building [Ern00, HM08]. Bullet [Ros07]. Bullock [LL08]. Burkert [SG18]. Burkhardt [Win09a]. Burning [Nic05]. Burns [Sor15]. Buy [WP15]. Byrd [McC14, Ker00].

C [Agr11, Bar10a, Bev08, DDV06, Fin10, MT12, NJ20, Ric12, Ste11, Tho07]. C. [Fau14, Hac01]. Cabaret [Gor08]. Cadet [HD08]. Calvin [OTC00]. cambel [Özd17]. Cambridge [Pea01, Fle15, Mal17]. Cameron [Bow18]. Campaign [Ger15]. Campbell [Bro17]. Can [Boi16, Mal09, Ron06, Vil04]. Candidacy [Tri17]. Cap [Hua06]. Capital [Meg03, Pie03b, Tab03]. Carew [Sco06]. Caricature [Bro01]. Carl [Bar00, Fis12, Hin01, JB00, Rak16].

Carleton [Roo15]. Carlo [Roe08]. Carlos [dA03]. Carrier [BGS^{+}08]. Carrington [Gie15]. Carter [DDV06, Pei04]. Caryl [Abe03]. Case [Ker00, Mil16, Reb11, Rei07]. Casimir [Lam02]. Caspersson [KK03]. Cassandra [War15c, Wei04]. Catalyst [Sei04]. Catholic [App19]. Cavendish [Sei04]. Celebration [Bar18a]. Cell [Bas10, Jac17, Qua13, Sca09]. Cell-Based [Sca09]. Cells [Wei06]. Census [And19]. Centennial [Wil17]. Center [Dar19]. Centuries [Syl14, Bas06]. Century [KRV12, Nor02, Ste03, Fin15, Gra15, Ham08a, Kev08, Nas19, RC07, Rid05, Sch05, War05, Was06]. Chadwick [Bro10]. Chagas [dA03]. Chains [Bru18]. Chairs [GBG15]. Challenge [Eva10, Ros00]. Challenges [Rak19]. Chance [Dut16]. Chandler [Wil09]. Change [Anc17, Coh10, Eva10, KRV12, Kit00, Par17, Tho10, Was17]. Changed [War12]. Changing [De 03, Wol00]. Chapman [Kat05a]. Chargaff [Coh04]. Charity [Rim01]. Charles [PT13, Ash09, Bir15, Con00b, Fis08, Key03, Kot00, Mar12, Moy01, Nic05, RB12]. Chemical [Sei04]. Chemistry [ISO^{+}07, MB07]. Chern [Wu07]. Chief [Mar09]. Chien [Lid01]. Chien-Shiung [Lid01]. Children [Fas15]. Childs [Cha01]. China [Ham08a, LB13, Mit08, Sei02]. Chinese [Hua06, Mal09, Pom08, Yu07]. Chips [Van18]. Choices [Wol00]. Christendom [Meg03]. Christoffel [Bla02]. Christopher [PT13]. Chronology [Gar18]. Cipolla [Roe08]. Circe [Wei04]. Citizen [Joh16, Tri17]. Citizens [Ger15]. Civilian [Reb11]. Civilization [Car06]. Clagett [Gra08]. Clara [Nor05]. Clark [Eag00b, Sim11, Pre00]. Clarke [Bro17]. Classic [Hea04]. Claude [Gal03]. Clause [Tri17]. Claw [Tho11]. Clerk [Sei01]. Clifford [Shw10]. Climate [Anc17, Bro02a, Coh10, Eag00a, Eva10, KRV12, Ken15b, Par17, Tho10, Was06, Was17]. Clinics [Bar09]. Clinton [O’S14]. Clothes [Kat05b]. Co [GBG15]. Co-Chairs [GBG15]. Coale [Pre05]. Cochran [Cha01]. Code [Les07]. Coeducation [Mal17]. Cognition [Tat16]. Cohen [Gin05]. Colin [KHEP08]. Collaboration [Ful06]. Collapse [Ber14, Sta03]. Collections [Pow15]. Colleges [Del12a]. Collision [Ern00]. Colony [Ber14]. Commission [GBG15]. Common [Coh11, Mar01b]. communi [Vir02]. Communication [Bas10]. Communities [Eva10]. comodo [Vir02]. Company [Vag01, Van18]. Complexity [And02, Ken15b]. Component [Yod02]. Compounds [Cla07]. Computer [Was06]. Concepts [Bej19, Blu11]. Concerns [Wal06]. Condition [Ste09]. Conduct [O’S14]. Cone [BL07, ISO^{+}07]. Confessional [HM08]. Conflict [Arr14, Wol14a]. Connection [Hua06]. Connections [Boo10]. Consistency [Gib14]. Conspiracy [Bak10]. Constitution [Mil15]. Constitutions [Mil15]. Constraints [Rut05]. Construction [Mil16]. Contagion [Fle15]. Contemporary [Jac10, MB07]. Contents [Ano13i, Ano14i, Ano15i, Ano17a, Ano18a, Ano19i]. Context [Tat16]. Contexts [Ada05, All07]. Continent [WP15]. Continental [Ern00]. Continuities [Bri07]. Contrast [Dun07]. Contribution [Fin01]. Contributions [Car06, Par17]. Control [Gar01]. Controversy [Gui02].

Conundrum [Sei02]. Convention [Doy19]. Conversation [Dar19, GBG15, Win19]. Cooper [Coh03]. Copernican [Swe04]. Coping [KRV12]. Copley [Bur01]. Copyright [Gin09]. Corinth [Rom06]. Coronal [Gie15]. Corporate [Gom14]. Corporation [Sto14, Syl14]. Corporations [Wol14b]. Correction [Ano00i]. Corrections [Sta06]. Correspondence [Coh03]. Cortex [SVD11]. Cosmic [Lun09]. Costs [Bog14]. Cotton [Fac09]. Council [O’M12]. Count [Jor11]. Countess [Rei15]. Counting [Gar08b]. Countries [Eva10]. Course [Pow15]. Court [SGTG10]. Courthouses [RC07]. Courts [Wal06]. Covenant [AQ03]. Cox [Bok06]. Craig [Ste08, Cro09]. Crawford [KG14, Bas13]. Creation [Hof13]. Creek [HD08]. Crick [Ste06]. Crime [Ste03]. Criminal [Wal06]. Crisis [Ber14, Mas07, SCG^{+}12]. Critical [Kit13]. Critique [Kit00]. Cronbach [Sha03]. Cronkite [Cun16]. Crosby [Mei19]. Cross [Gol12]. Crozier [Mar15a]. Cruz [Tri17]. Cult [Bak10]. Cultural [Gre10a, Pye02, Sch05]. Culture [Gra13, Mit08, Yu07]. Curculionites [Pre13]. Cure [Ric12]. Curious [Meg03]. Curtin [Van11]. Custody [Bru18]. Cutter [Sha05].

D [Buc13, Gra02b, McK01, Sta05, Van11, Wil09]. D. [Roo15]. d’être [Lea12]. Dahl [May17]. Dahrendorf [Sme11]. Dainton [Kor00]. Dame [Zio20]. Damned [Mal17]. Daniel [Bau05, CR10, Dan10, Mor06, Pin01]. Dark [Ger15, Gor08]. D’Arms [Con10]. Darwin [ME13, Bro01, Bur01, Key03, Nic05, Win01]. Darwinism [May01]. Dating [Tra00]. David [Doo04, FDBH17, HD08, Hel05, Sin04, Atk08, Bri12, Bur08, Cox19, Eag08, FP06, Hog14, Hor14, Joh06, Kno17, Min20, Ram08]. Davidson [Sup06]. Davis [Con00b, Min20]. Davison [Olo15]. Davy [Tho13]. Day [War15u]. Dead [Bar19]. Deal [Gar16]. Dear [Ano00i, All00]. Death [Pie03a]. Deaths [Ano13j, Ano14j, Ano15j, Ano17b, Ano18b, Ano19j]. deb [Gar13a]. Debates [Jac10]. Debreu [Arr11]. Decade [Rot12]. December [Ada00, Alb14, Ano13j, Ano14j, Ano15j, Ano17b, Ano18b, Ano19j, Arr11, Atk08, BIK^{+}08, Bir15, Boe07, Bri00, Con00a, Cro09, Cul17, Dar01, Føl12, Fra04, FL13, GTCD13, Gou08, Gra11, Gut15, Hac01, Har15, HM06, Hen02, Höl10, Hun07, Hux01, Jac03, JH18, KK03, Kle11, Kor00, Kro01, Lak11, Lan16, LL08, Mar08, Max07, May17, McK01, Mit07, Oki12, Olo15, Pap05, Phi05, Pri07, Pye05, RC00, Rak16, Ram08, Ree01, Rid04, Roo15, Rot09, SBC^{+}14, Sed07, Sin04, SCL^{+}10, Sny07, SR00, Sud00, Tel02, Ter10, Tri14, VW12, Wu07]. Decide [Pol04a]. Decision [Arr14]. Decline [Mar01a]. Deeds [Hea04]. Defense [Bre05, Can01, Fin01]. Deferential [Cor17]. Defining [Mar01b, Wol01a]. Degler [Rak16]. Degree [Gor08]. Deliberation [Fis19]. Demise [Jam15]. Democracy [Fis19, Rak19]. Democratic [AP06, Fis19]. Demographic [Dun07, Von11]. Demography [Coh09]. Dennis [Boy07, Ree01]. Depends [Sab07]. Derek [Cot00, Jon18]. Derk [Le 06]. Design [Rub20]. Destroy [Bak10]. Detection [Rag13]. Developed [WP15]. Developing [Her05, NM05]. Development [Pom08, Qua13, Wol00].

Developments [Kit13, Pow15]. Devonshire [Ros06a]. D’Harnoncourt [CD12]. Diabo [Rei07]. Diamonds [Ern00, Ker09]. Dicke [HP06]. Dickinson [Lan02]. Did [Pol04a, Wil17]. Diet [FH08]. Dietz [FS08]. Digby [Pri07]. Digital [Dud01, Fuc01, Mar01b, Tho09]. Dil [SR00]. Dilettante [Eas00]. Dillon [Cha03]. Dimitri [Fra04]. Dinner [Gre19]. Dionysus [Ols19]. Disabilities [Gly16]. Disappear [Fur15]. Discontent [Hag02]. Discourse [Jam15]. Discoverer [Tho13]. Discoveries [EK11]. Discovery [Car06, Cla07, Lon11, Mar17, MB07, Sch11]. Discussion [Bor17, SCG^{+}12, SGTG10]. Disease [Kil04]. Dispossessing [Bru18]. Disremembered [Hol16]. Dissemination [Bro01]. Divided [Fis13]. Dmitri [Mat01]. DNA [Cof04, Rei18]. Do [Gre10b]. doctorum [Vir02]. Does [Kat05b]. Dogs [Hor15]. Domenico [Asc11]. Domicile [War15s]. Donald [Bed14, PP14, Ris05, Sup06, Wyn04]. Doolittle [KFW20]. Doty [Alb14]. Doubling [Von11]. Downward [Fis13]. Dream [Dub11, Hag09]. Dreams [Bag02]. Drickamer [Sli04]. Drugs [Kea12]. Dudley [GHX19]. Due [Tab03]. Duke [Ros06a]. Dumas [Coh03]. Duncan [Sup14, GHX19]. Dunlop [Shu06]. During [Ken15a, Hag02]. Dusen [Max07]. Dust [Bro02a]. Dworkin [Tri14].

Each [All00, Ano00i]. Eakins [Bar09]. Early [Man20, Wol01a, Sha05]. Earnings [Sie14]. Earth [Mor10, Sch02]. Earths [Mar10a]. East [Ulr20, Win09b]. Eastburn [Eag08]. Ebert [Sin04]. Eccles [Red08, Sco06]. Economic [AP06, Bli19, Mal09, Nas13, Qua13]. Economists [Win01]. Economy [Ros00]. Edelman [Rut15]. Edinburgh [Ati06]. Edmund [Mid15]. Edsall [Dot05]. Education [AP06, Gut14, Les04, McK04, Pol04b, WH14]. Educational [Gra02a]. Edward [BL07, DDV06, Gra02c, Ken16, Lea15b, Lip06, Mon08, Pec18, Woo06a]. Edwin [Bur08, Buc13, McE02, Ter10]. Edzard [FS08]. Effect [PW01]. Effective [Kat05b]. Effects [LHM05]. Egregiously [Tab03]. Egypt [Ham08a]. Eight [Tal06]. Eighteenth [Fin15, Sch05, War05]. Eighteenth-Century [Sch05, War05]. Einstein [Wil17]. Eisenstadt [BR12]. Ejection [Gie15]. Eldridge [Mat10]. Elected [Ano13p, Ano14p, Ano15p, Ano17d, Ano18d, Ano19l]. Election [GBG15]. Elfriede [Rid11]. Elinor [Wal13]. Elizabeth [Bar03b]. Eloquence [Mul08]. Elwood [Gal03]. Emancipation [Hol16]. Emanuel [GG01]. Emblematic [Mar13]. Emeneau [Bri08]. Emergence [Tat16]. Emil [GH13]. Emily [Lan02, Mas00]. Empire [Mat07b, HM08]. Enabled [Bal12]. End [Rot12]. Endless [Ste04a]. Enduring [Var07]. Engagement [Fla05]. Engine [Kit13]. Engineered [Rot12]. Engineering [Kea12, Vac07]. England [Kit00]. Enigmatic [Ols19]. Enlightened [Jac20]. Enlightenment [Rud05b, Ati06]. Enterprise [Wil05b]. Entrepreneur [LB13, Nas13]. Entrepreneurship [CHH13, Sch13]. Environments [Des10]. Envy [Fis13]. Ephraim [Sel11]. Epidemic [Gui02]. Epigenetics [Gra16]. Epistemologic [Sta03].

Eradication [Ric12]. Erasable [Sta06]. Erica [Rot09]. Ernest [Kih03]. Ernst [Boc07, Hop14, Mit07]. Erwin [Coh04]. Esmond [Mul10]. Essay [Swe04]. Essays [Dun08]. Essential [Lea12]. Establishing [Tat16]. Ethical [Sat03]. Ethics [Wei06]. Ethnography [Jac10]. Eugen [Hun09]. Eugene [Sil04, Sud00, Bai10]. Eugenics [Sel05]. Eugenio [Boe07]. Europe [Jos18]. European [Gar18, Kai07, Rei18, Van07, Yu07]. Europeanization [Jos18]. Eve [Cul14]. Evelyn [McC14]. Even [Bar19]. Everett [Fau14]. Everyday [App19, Lea15a]. Evidence [Des10]. Evident [Col04]. Evolution [Bro01, Cof04, Fin12, Gin12, ISO^{+}07, Jab12, Kil04, Mor10]. Evolved [Syl14]. Evolving [Arm10, Rim01]. Evon [Bri06]. Example [Jab12]. Examples [Mid06]. Executions [Ste03]. Existence [Nap05]. Exoplanet [Bor17]. Exotic [Sch05]. Expanding [Fin05]. Expectancy [Fin12]. Expert [Sto15]. Exploited [Jam19]. Exploration [Ron06]. Exploring [Des10, Pet09]. Extant [Pre13]. Extending [Bre00]. Externalities [Smi01]. Extravagant [Von11].

F [Bai10, De 03, Got08, Hin08, Lev07, Mat07a, PO13, Wol00, Wor18]. Fabrication [Sil16]. Face [All07, Dub11, Lan11, SB11]. Facing [Lan11]. Fact [Jam15]. Factionalism [Nor05]. Factors [FH08]. Fagles [Bro09]. Fairmount [Mar10b, Mar13]. Fake [Boi16]. Fallen [Win01]. Families [Sel05]. Family [Lea03, Mar12]. Fandango [Lin08]. Faraday [Tho06]. Farnsworth [Lie07]. Fascination [Seg17]. Feasible [Rot12]. February [Ada00, Alt09, ALSS10, Atk08, AF03, Ber00, Bev08, Bla10, Bob07, Boc07, BG11, Bri08, Bro05, Bro17, Buc13, Cro02, DH07, Dys06, Fac09, Fau14, Fis12, Fon07, Gal03, Gol12, Gou11, Gra11, Hei05, Hel05, HKT14, Hol04, Hux01, Jam11, Jef16, Kop06, Kos05, Lan02, Lid01, Mat07a, Mat09, May17, McC16, McC00, MN11, Mon08, Mor06, Moy01, Mum08, Not16, Ort04, Özy08, Pap05, Pat09, Pin16, PP14, Pro07, Put05, Rak16, SK10, She14, SG18, Tho04, TS14, Tho18, Tip15, Tri14, Vac09, VF08, Wal09, Wha03, dA03]. Federal [Kel00]. Feed [NM05]. Feet [Tal06]. Fetal [Adz13]. Field [Bal12]. Fielding [Pea14]. Fifteenth [Ham08a]. Fifth [Bas06]. Fiftieth [McK04, Pol04b]. Figaro [Lin08]. Filho [dA03]. Film [Gor08]. Final [Gre04a]. Finance [Ger15]. Finding [Jon18, Lea15b]. Findings [Key03]. Fire [War15d]. Firm [Nas13]. First [RC07, Ste03, Swe04, Was06, Mar09, Rid05]. Firth [Ort04]. Fish [ISO^{+}07]. Fish-Hunting [ISO^{+}07]. Fishing [Dow02a]. Fit [Tal06]. Fitch [Aya12]. Fitness [Ste09]. Fitter [Sel05]. Five [Miy11, Von11]. Fixe [War15s]. Floyd [Cha00, Zem02]. Flux [Ful06]. Foam [Wea01]. Folding [Wol01b]. Folkman [Vac09]. Food [Dys01, Her05]. Forced [Sat03]. Ford [BIK^{+}08]. Foreign [Van07]. Forgery [Jon16, Rag13]. Forging [Wil16]. Forty [Cer14, LHM05]. Forty-One [LHM05]. Forty-Year [Cer14]. Fossils [Cof04]. Foster [Kle16]. Found [Woo06b]. Foundation [Rud14]. Foundations [May01]. Fourth [Ros00]. Fowler [SPF03]. Fragonard [Ros06b]. Framed [Dar19]. Framing [Dar15]. Francis [Bow10, BGS^{+}08, Wit20, Gut15, Sim11, Sta05, Ste06, Ste05]. Francois [Mor16]. Frängsmyr [Hei18]. Frank [Fac09, Fis09, Goh05]. Franklin [Ano00i, Lew11, All00, Ati06, BIK^{+}08, FH08, Fin15, Gre19, Hei06, Mul08, Nas06, Rem06, Ros06b, Sta06, Tal06, Tho06, Van18, Von11, WP15, Zew06]. Frantz [McC00]. Franz [Bak10, Dar18, Dir10, Gre10a, Gut05, HO14]. Fraser [Rud05a]. Fred [Bur03, Gin10]. Frederic [Hei05]. Frederick [Bed14, Bar03c, Fon07, Kat05a, Kor00, Mar10b, Sli10, Win09a]. Fredrickson [Wyn04]. Free [Ste04b]. Freedman [Gar08a]. Freedom [Atk04, Bej19, Gly16]. Freedoms [Hum07]. French [Ful06]. Friedman [GP02]. Frieman [Ken16]. Friend [Ker09]. Friendship [Hag09]. Front [Ano00j, Ano00k, Ano00l, Ano00m, Ano01i, Ano01j, Ano01k, Ano01l, Ano02i, Ano02j, Ano02k, Ano02l, Ano03i, Ano03j, Ano03k, Ano03l, Ano04i, Ano04j, Ano04k, Ano04l, Ano05i, Ano05j, Ano05k, Ano05l, Ano06i, Ano06j, Ano06k, Ano06l, Ano07i, Ano07j, Ano07k, Ano07l, Ano08i, Ano08j, Ano08k, Ano08l, Ano09i, Ano09j, Ano09k, Ano09l, Ano10i, Ano10j, Ano10k, Ano10l, Ano11i, Ano11j, Ano11k, Ano11l, Ano12i, Ano12j, Ano12k, Ano12l, Ano13k, Ano13l, Ano13m, Ano13n, Ano14k, Ano14l, Ano14m, Ano14n, Ano15k, Ano15l, Ano15m, Ano15n]. Frontier [Arm10, Pet09]. Frustrating [Hil00]. Fruton [SA09]. Frye [Pel06]. Fuels [Kea12]. Fundamental [Gut14]. Funnels [Wol01b]. Further [Sim09]. Future [Anc17, Dar18, Dud01, Fin12, Gin09, Gre09, Her14, Kov09, Lan11, Pil18, Rut05, Sie14, SS18, Tau19, Tho09, Tho17, Wil05b]. Futures [Gre10a].

G [Cro18, Kop06, Sli04]. G. [May02]. Gabriel [Pye05]. GAFOH [Dun08]. Gain [O’S14]. Gajdusek [Roo15]. Galbraith [Hod12]. Galileo [Gin11]. Game [Por11]. Gardner [ALSS10, Pat09]. Garin [Boe07]. Garrett [Cai05]. Geertz [Shw10]. General [O’S14]. Generation [Bal12]. Generational [Boo10, Fas15]. Genes [Bea04, Dow02a, FH08]. Genetically [Rot12]. Genius [NM05, Wol01a]. Genome [Bal12, NM05]. Genome-Enabled [Bal12]. Genuine [Boi16]. Geodynamics [And02]. Geology [Yod02]. Geomorphic [LHM05]. Georg [SK10]. George [Ham08b, Sed07, All09, BGS^{+}08, Cra01, Dow02b, Eag00b, Hag02, Mat07a, McK01, Mum08, She14, Tho07]. Georges [Bro05, de 17]. Gerald [Cas04, Rut15]. Gerard [DP07, Arr11]. Gerhard [Lam02, Kro01]. Germaine [Bis03]. German [Mar01a]. Gerry [BWW16]. Get [Wil17]. Ghosts [And19]. Gilbert [BF13, Hin08]. Girl [Ker09]. Gita [Hij00]. Giuseppe [Gos12, Tel02]. Glacier [Tho17]. Glaser [Ber13, PP14]. Glass [Erk09]. Glasses [Wol01b]. Glaucoma [War15e]. Gleason [MGM10]. Glenn [Cha00, Kos05, Sei04]. Global [And02, Fas15, Ken15b, Qua13, Sch13, Tho10]. Globalized [Hen04]. Go [Bea04, Ker16]. Goals [Gom14]. Gobind [Sak17]. God [Coo03, Böw03]. Goetzmann [Slo12]. Goheen [Bow10]. Gold [Sal06]. Gol’danskii [DT03]. Golden [Rho09, Fut12]. Goldstine [GA06]. Gombrich [Hop14]. Good [Blo12, Mar01b]. Goodenough [Mar15b]. Gordimer [Cli15]. Gordon [Cho12, Deg02, Rud05a, Sab04, Ste08]. Got [Ham08a]. Gout [FH08].

Governance [Nor05]. Grabar [Nec12]. Graff [Mar10b]. Graffiti [War15f]. Graham [Mon08]. Grammatology [Ham08a]. Grand [Nas13]. Great [Isa06, Tho11]. Great-Claw [Tho11]. Greatest [Sch02]. Greek [Bas06, LJ01, Rid05]. Green [Eve05]. Greenberg [Sil09]. Greenewalt [Bas13]. Greenhouse [PW01]. Greenough [Bow09]. Greenstein [Tri05]. Gregory [DF04]. Griffin [Ris05]. Group [Smi01]. Groves [Gra14]. Growth [Nas13]. Gunther [Cas04, Bar10b]. Gustav [Alt09, JB00, Rei02]. Güterbock [Rei02]. Gutowsky [Dri02]. Guyford [Pro12].

H [App19, Bas13, Bri12, Con10, Cot00, Fis09, HKT14, Jam04, Mar15b, Max07, SBC^{+}14, Sil09, Slo12, TS14, Tho18, Win09a]. H. [Erk09, Kop06, Pro12, Ste11]. Haas [MT12]. Habakkuk [Lan05]. Habitable [Des10]. Hackerman [Fox09]. Hackers [Jam19]. Hackney [Gut15]. Halet [Özd17]. Half [KRV12]. Half-Century [KRV12]. Hallam [Kor00]. Hamilton [Gra02b, HCC06, Ost07, Tre09]. Hand [Dub11, SVD11]. Handlin [Bai13]. Hannes [FD06]. Hans [Hop14, Dre07, Rei02, SK10]. Hans-Georg [SK10]. Happening [Del12a]. Happiness [Wol14a]. Har [Sak17]. Hardin [Cai05]. Harold [Kir03]. Harrison [McC14]. Harry [Sli04, Wil05a]. Hartshorne [Mas00]. Harvey [McK00, Bra10]. Haskell [Ste05]. Haskins [Abe03]. Hastings [Mar03]. Hatten [Mys05]. Hauge [MS07]. Health [All07]. Heaney [Ven15]. Hearts [O’S14, Pie02]. Heckscher [Ada03]. Heeschen [Hog14]. Hegemony [Odo07]. Heiko [Pel03]. Heine [GA06]. Heinrich [Orc19]. Heinz [Wal07]. Helen [PO13, Rot11]. Hempel [JB00]. Hendrik [Bla02, Lam02]. Hennessy [RC00]. Henry [HP06, O’S14, Pin01, Bro10, Dav12, DeL12b, GC14, Sei04]. Hepatitis [Ric12]. Her [Lea03]. Herbert [McE02, Ste05, AF03, Dri02, GP02, Hen02, Hun07, Jon08, Tha17]. Herman [GA06, Wal02]. Hermann [Gar13b]. Hero [Kit13]. Heroic [Mil16]. Herring [GR06]. Herzberg [Kro01]. Hewlett [Gar03]. Hiatus [Ken15b]. Hidden [Tat19]. Hieroglyphic [Bri07]. Hieroglyphs [Ham08a]. HigginBotham [Ada00]. Higher [AP06, Gut14]. Highest [Tho10, Tho17]. Hilgard [Kih03]. Hilleman [Fau07]. Hindle [Bed03]. Hiroshi [Tho02]. Hirsch [Gra02c]. Historical [Mcc05]. History [AQ03, Bog14, Bra08, Gil14, Gos01, Man20, Meg03, Pye02, Sel05, Smi04]. Hitchcock [Pip17]. Hoaxes [Mul08]. Hodgkin [Hux01]. Hoenigswald [Dav12]. Hoffmann [GK17]. Holmes [Hei05, LL08, Pin01]. Holy [HM08]. Homer [McC02]. Homeric [Wes11]. Honest [Coo03]. Honey [Ber14]. Honor [Bar03a]. Hope [Lew11]. Horace [San06]. Hornig [Bed14]. Horticulture [Kev08]. Hot [Mil16]. House [BD19, Lea15b]. Howard [CSG17, VW12]. Howell [Sim11]. Hoyle [Bur03]. Hugh [Cro09, Par12]. Hughes [Pin16]. Hulst [Bla02]. Human [AQ03, Fin12, Gre10a, Jab12, SVD11, Ste09, Ste04b, Tat16, Vil04, Wol02]. Humanities [Blo12]. Humanizing [Bil01]. Humboldt [Cas11]. Hume [Pea01]. Humor [Mid06]. Humphry [Tho13]. Hundred [Gre04a, Plo18, Rem06]. Hunting [ISO^{+}07]. Hurst [Kat00]. Hutchison [Pal03]. Huxley [Pea14]. Huxtable [Rus16]. Hyde [Red08].

I. [Gin05]. Iconography [RC07]. Idea [HM08]. Ideas [War12]. Identity [Ful06, Mof20, Sha05, Tre19]. If [War15b]. Ignorance [Sch05]. Igorevich [Yak12]. Ihor [Gra11]. II [DDV06]. Illegal [Rei07]. Illuminated [Tri17]. Image [Wol01a]. Images [Win09b]. Imagination [Mas16, Yu07]. Imaging [Bra08]. Imbrie [Gou11]. Immediate [Eva10]. Immigrants [Por11]. Immigration [Mas07, Rei07]. Impact [Gin11, Jac17, Wol02]. Impacts [Eva10, Lea15a, Mor10]. Implementation [Tab03]. Implications [Jam15, Sch13, Tau19]. Importance [De 13]. Impostors [Jac20]. Inaugural [HO14]. Inconstant [VF06]. Index [Ano13o, Ano14o, Ano15o, Ano17c, Ano18c, Ano19k]. Indian [HD08, Kel00, Mil15, WP15]. Indigenous [Dar19, Pow14, Tre19]. Indignation [Coo03]. Individual [Smi01]. Indo [Gar18, Jos18, Rei18]. Indo-European [Gar18, Rei18]. Indo-Europeanization [Jos18]. Industrialization [LB13]. Infinite [EN10]. Inflammatory [Fin12]. Inflation [Sie14]. Influence [Mil15]. Influenza [Bar19, Pal19, Tau06, Tau19]. Information [Ano00n, Ano01n, Ano02m, Ano03m, Ano04m, Ano05m, Ano06m, Ano07m, Ano08m, Bil01, Gre09]. Infraspecific [SPF03]. Ingalls [Pin01]. Inkeles [Smi12]. Innovative [LB13]. Inose [Tho02]. Inscriptions [Tra00]. Insect [Cos02]. Insects [Lea12]. Insights [Tau19]. Intellectual [Her05, Jac10, Kev08]. Intelligence [Lo13]. International [Doy19, Pet09, Wal06]. Internet [Var09]. Interpretation [Tat19]. Intrigue [Ken15a]. Introduction [Arm10, Bar03a, Bar18a, Dar15, Dun08, EK11, Gre09, Gre10a, Jos18, Kil04, Mat07b, MB07, Rak19, Sil16, Ste04a, VF10, War15g, Was17, Wol05]. Introductory [McK04, Pie03b, Tat10]. Invention [Kit13, LB13, MB07, Mil16]. Inventor [Tho13]. Investment [Bog14]. Invisible [Gly16]. Ion [Gel01]. Iran [Gar16]. Iraqi [Ada05]. Iron [Mil16]. Ironworker [Rei07]. Iroquois [Sta08]. Irving [For03]. Irwin [Mil06]. Isaac [Rig00, Bia04]. Isaiah [BSW06]. Isidor [Rig00]. Islam [Sim09]. Issues [Jos18, Sat03, Wol05]. Italian [Gos12]. Ivanovich [Gra11]. Ivy [Mal17].

J [Bra06, Hij00, Mor06, Nat07, Sha03, SCL^{+}10, Tho08]. J. [Kar05, Pei04, SR00, Wol01a]. Jack [Phi07]. Jackson [Eng01]. Jacob [Mor16, War15j]. Jacqueline [FL13]. Jamaica [CHH13, Dun07]. James [Cri00, Rud05a, BF13, Bar10a, Cro02, Fla05, Gar08a, Kat00, Kni08, Mid11, Mur07, Sei01, Sin04, Sol04, Sor15]. Jameson [Ost07]. Jan [Mar08]. Janet [Olo15]. January [Bar03c, Bar05, BK07, Bed14, Boy07, Bur08, Con00a, Con10, Cra01, Cri00, Deg02, Dri02, DT03, DDV06, Eag00b, Eag08, Erk09, EDJG12, Fis09, Gar03, Gar13a, Gra08, Ham07, Har10, Hel05, Hin01, JB00, Kar05, Kel01, Ken03, Ken16, Lew11, Lie07, May02, MS02, Mid11, Mid15, Mis00, Mit07, Nec12, Nel11, Nor12, Oki12, OTC00, Özd17, Pel06, Pet01, Pin09, Plo09, Pos12, PO13, Pye05, Rai09, Ram04, RK09, Rig00, Ros06a, Rot11, Rus16, Sak17, Sco06, SCL^{+}10, Ste05, Sto11, Tho02, Vac09, Wal07, Wat05, Wil05a, Wol04, de 17]. Japan [Ros00]. Japanese [Nap05]. Jaroslav [Mar08]. Jayne [Bar09, Hea04, Key03, RC07, Vac07]. Jean [And00]. Jefferson [Mon08, Cas11, Coh03, Hag09, Hal18, Kel00, Pow15, Ron06, Spe18, Tho11]. Jeffersonian [Dar19]. Jencks [RK09]. Jeremy [Her17]. Jermy [Löv19]. Jerome [Gar17]. Jesse [Tri05]. Joan [Ric15]. Johanna [Özy08]. John [Lan05, RC00, Wat05, BDN04, Ben02, Bro02b, Con10, Dot05, Dys10, FEE11, Gou08, Gra19, Har10, Hen04, Hod12, Lew11, Mar09, McC16, MT12, Mos05, NJ20, Par06, Pin09, Pol04a, Put05, Ruf06, Shu06, TS14]. Johns [Pri07]. Johnson [Pre05, VW12]. Joining [Mar13]. Jonathan [Bar03a, Bar05]. Jones [Par12]. Joseph [Mil06, Nas13, Sil09, SA09, Tom15]. Joshua [BG11]. Journey [Hil00]. Joy [Ber13]. Joyce [JH18]. Jr [Ada00, Bai10, Bar03c, Bas13, Bri06, Bri00, CR10, Cun16, Dun11, EC02, Fin10, Jay14, KG14, Mat09, Max07, Mys05, Ost06, SCL^{+}10, Tha17, Wil09, Wol00, Wor18]. Judah [Vac09]. Judges [Gre10b]. Judgment [O’M12]. Judicial [Gre04a]. Judith [Ben04]. Juggler [Zio20]. Julesz [Pap05]. Julian [Mar03]. Julius [Sny07]. July [Ack04, All09, Arr11, Asc11, Bar16, Bia04, Bla02, BL16, Boc07, Bow09, Bow18, Bri04, Bri12, Bri00, Bro02b, Buc13, Cas04, CR10, Cli15, Con03, Cri00, Cro18, Cro09, CSG17, Cun16, DF04, DH12, Dre07, Dun11, Dut16, Dys10, EC02, Eas18, Eng01, EDJG12, Fis08, For03, Frè16, Geo07, GH13, GG01, Har15, Hei18, Hol04, Hun07, Kel01, Kih03, Kno17, Kop06, Lam02, Lev07, Lig17, Lip06, Mad15, McC16, McK00, MLT11, Mid15, Mis00, Mon08, Mos05, Mur07, Not16, Pel06, PT13, Pin01, Red08, Rid11, Rig00, Rud05a, Ruf06, Rut15, Sed07, Sei08, She14, Shu06, SA09, Sli10, Slo12, Smi12, Sor15, Ste06, Sto11, Tha17, Tho07, Tip15, Tre09, VW12, Wat05, de 17, War15h]. June [Alb14, AF03, BF13, Bar10b, Bas13, BK07, Bed03, Beg07, Ber13, BSW06, BL16, Bro17, Bro10, Bur03, Coh04, CD12, Dar01, Dav12, Dot05, DT03, Eag00b, Fin10, Fle17, Føl12, Fon07, Fox09, FS08, Gar12, Gar17, Gin05, Goh05, GA06, Gra02c, GP02, Hac09, HCC06, Jay14, Joh06, Kat00, Lie07, ME13, Mar15b, Mat10, May09, McC14, McK01, Mon01, Mor16, Mos05, Nat07, Ost06, Pal03, Pat09, Pea06, Pei04, PT13, Phi07, Pro07, Put09, Pye01, Ric15, Rid11, Sal06, Sme11, Ste06, SR00, Tha17, Van11, Wal02, Wal13, Wal07, Wol04, Wri07, Wyn04, Yak12, Zem02]. Jurisdiction [Res16]. Justice [Pie03b, RC07, Mar09]. Justin [Ven15].

Kamen [Doo04]. Karl [Pye02]. Katchalski [Sel11]. Katz [Sim16]. Katzenbach [Gar13a]. Katzir [Sel11]. Katzir-Katchalski [Sel11]. Kaysen [Fis12]. Keating [App19]. Keep [Mal17]. Keith [Bar18a, Pea06]. Kennan [Mat07a]. Kenneth [Hod12, Con00a, HKT14, Jef16]. Kepler [Bor17]. Kerman [Tom15]. Kety [Sok02]. Keynes [ME13]. Khorana [Sak17]. Kilbourne [Buc13]. Kilby [Phi07]. Killing [Kuk14]. Kindleberger [Fis08]. Kirk [Kar05]. Kitzinger [Mit07]. Klein [MO20]. Knauer [Rid11]. Knolege [Ron06]. Knopoff [EDJG12]. Know [Gre10b]. Knowledge [Dar18, SGTG10]. Knowles [Her17]. Knox [Eas18]. Knudson [Cro18]. Konrad [Ern05, Ken03]. Koop [Fau14]. Kornberg [Ber09]. Korsmeyer [Nat07]. Koshland [CR10]. Kosofsky [Cul14]. Krauskopf [Ern05]. Krautheimer [Ack04]. Kristeller [Mon01]. Kudos [Wal06]. Kuhn [Swe04]. Kurt [Gar13b, Rai09]. Kusch [Ram04].

L [BIK^{+}08, Cra01, Dun11, GH13, Ham08b, KG14, Lan16, Tri05]. Lack [Tab03]. Lamont [Gol12]. Lamppost [Bli19]. Lamprecht [Pye02]. Land [Gin12, McE02]. Landau [Jor06]. Landon [Ste11]. Landscape [De 03, Man20, Sto06]. Landscapes [Wol01b]. Lane [Awo07]. Lang [DH12]. Langfitt [Rim08]. Language [Lea03, Sto06, Tre19]. Languages [Rei18, Woo06b]. Lardy [DeL12b]. Last [Mal09, O’M12]. Late [Kev08]. Laudise [Cho00]. Laughter [Gor08]. Laundering [Ber14]. Law [Res16, Ros07, Sto14]. Lawrence [Gin10, Hei05, Dar01, HM06, MO20]. Lay [Joh16]. Lead [FH08]. League [HM08, Mal17, Sta08]. Lear [Pec18]. Learning [Lan11]. Lecture [HO14, Bar09, De 03, Hea04, Key03, RC07, Vac07, Wol00]. Led [Kit13]. Lederberg [BG11]. Ledyard [May02]. Lee [Sha03]. Legacies [Hol16]. Legacy [Dar18, Dar19, Gre10a, Mar18, Mit01]. Leighton [Tre09]. Leland [Cun16]. Length [Ven01]. Leon [Ada00, EDJG12]. Leonard [Tho08]. Leonore [Jam11]. Leontief [Ber00]. Leopold [VF08]. Leprince [Jac03]. Leprince-Ringuet [Jac03]. Leray [And00]. Leslie [Sed07]. Lessons [Bar19, Bog14, Rak19]. Levi [Gra02c, Ang19]. Levi-Montalcini [Ang19]. Levy [Jef16]. Lewis [Bla16, Lip06, Hag02, Max07, Pre00]. Liberty [BD19]. Libraries [Var09]. Library [Bag02, Ano19e, Ano19f, Ano19g, Ano19h, Ano20b, Van18, Vir02]. Life [App19, Des10, Dun07, Fin12, Lun09, Mar10a]. Lifespan [Fin12]. Lighthill [Cri00]. Like [Jon18]. Likhachev [Mat01]. Limb [SB11]. Limits [Fla05, Ste04b]. Lin [Wal13]. Lindauer [Höl10]. Lindzey [ALSS10]. Link [Coo06]. Linnaean [Wil05b]. Lipset [Lak11]. Litchfield [Lig17]. Literary [Bri07]. Litz [Fle17]. Living [Gar08b]. Lloyd [Par12]. Lloyd-Jones [Par12]. Load [Fin12]. Locating [BD19]. Long [Pal19, Pom08]. Long-Run [Pom08]. Longevity [Hir00]. Look [Fel04, Sab07]. Looks [Wea01]. Looting [Ada05]. Lord [Phi05]. Lose [Odo07]. Lost [Bru18]. Lou [WH14]. Louis [PRP^{+}04, Jac03, SBC^{+}14]. Louise [DH12, Rus16]. Louisiana [Kel00]. Lounsbury [Cha00]. Lovell [Mor13]. Löwdin [Lin03]. Lowlands [Sab07]. Luce [Sup14]. Lucian [VV11]. Luck [Mcc05]. Luddenham [Phi05]. Lullabies [Miy11]. Lumumba [Kuk14]. Luna [VF08]. Lundquist [And03]. Lyman [Cas14, Ost06].

M [Bir15, Eas18, Jay14, Roe08, Zuc08]. M. [Vac09]. Mabel [DH12]. Mac [Awo07]. Maccready [Wyl11]. MacDonald [Rud05a]. Macgregor [Sor15]. Machteld [Özy08]. Mackey [Mum08]. Maclyn [BK07]. Madison [Hag09, Kra04, Pol04a]. Maffei [Asc11]. Magic [Jac17, Saa16]. Mahoney [GTCD13]. Mainstream [Fin05]. Make [Vil04]. Malcolm [Kno17]. Malone [Coh03]. Malthus [Smi01, Eve05]. Malthusian [Dys01, Kit00]. Man [Cas11, Fut12, Lea15b, Tho13, War15i]. Mandela [Har15]. Manfred [Sch14]. Manning [Wal09]. Manuscript [Coh03]. Marbury [Fel04, Gre04a, Kra04, Pol04a]. March [Aya12, BDN04, Bar03c, Bar18b, Bar10b, Bau05, Bed14, Ber09, Bla16, Bla10, Bow10, Bow11, Bow18, Bra06, Bro09, BGS^{+}08, Cha03, Cho12, CR10, Coo06, Cot00, Cro02, DP07, Dre07, Fis12, Fox09, FL13, Gar08a, Gin05, Gle02, Gra02b, Gra02c, HP06, Hei05, Hog14, Hop14, Jac03, Jam04, Jam11, KHEP08, Kir03, Kni08, Kro01, Lak11, Le 06, Lev07, Lew11, Mat07a, MS02, McE02, Mum08, Mur07, Mys05, Nat07, Nie00, Ort04, Osm03, Ost06, Pye01, Ram04, Rei15, Rei02, RB12, Ruf06, Rus16, Sab04, Sim11, Sli10, Smi12, Sol04, Ste11, Sup06, SG18, Tho07, Tom15, Wal02]. Margaret [GTCD13]. Marianne [Lea03]. Mark [Mid06]. Market [Her05, Sto15]. Marler [Not16]. Marriage [Fur15]. Mars [Des10]. Marshak [Sud00]. Marshall [Dys06, Gra08, Hen04, Mar09, Nas13, Nel11, Pol04a]. Martí [Bad17]. Martin [Bri00, Bro02b, Cox19, Lak11, Doo04, Hac09, Höl10, Lig17, Nor12]. Marvels [Ber06]. Marvin [Sil04]. Mary [Red08, Rei15]. Mass [Gie15, Hir00]. Material [Pre13]. Materials [Pow14]. Mathematics [Coh11, Gra13]. Matter [Ano00e, Ano00f, Ano00g, Ano00h, Ano00j, Ano00k, Ano00l, Ano00m, Ano01e, Ano01f, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02j, Ano02k, Ano02l, Ano03e, Ano03f, Ano03g, Ano03h, Ano03i, Ano03j, Ano03k, Ano03l, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano05j, Ano05k, Ano05l, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano09e, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l]. Matter [Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13e, Ano13f, Ano13g, Ano13h, Ano13k, Ano13l, Ano13m, Ano13n, Ano14e, Ano14f, Ano14g, Ano14h, Ano14k, Ano14l, Ano14m, Ano14n, Ano15e, Ano15f, Ano15g, Ano15h, Ano15k, Ano15l, Ano15m, Ano15n, Hau20, Ven01]. Mattered [War05]. Matters [Var03]. Maurice [Ada03, Fau07]. Max [Dav12, Tho04, War15j]. Maxwell [Sei01]. May [Ada03, Alt15, Ash09, Aya12, Bad17, Bar03b, Bar05, Bar10a, Bar00, Bas13, Bis05, BG11, Boe07, Bok06, Bom10, Bor01, Bow09, Bra10, Bri06, Bro05, BGS^{+}08, BL07, Cas04, Cas14, Cha00, Cha01, Cho12, CT04, Cul14, Dun11, Eng01, Ern05, FD06, FP06, Fit02, FEE11, Gar03, Gar13a, Geo07, GH13, Goh05, GG01, Ham08b, HP06, Hor14, Hun09, Jor06, Kle11, Lam02, Lan05, LL08, Lid01, Lip06, Mar15a, Mar15b, Mar08, Mas00, Mat10, McC02, McE02, McK00, MN11, Mil06, MT12, Mon01, Pea06, Pea14, Pin09, Pin01, Rei02, Rid04, Ros06a, Rud05a, Rut15, Sal06, SK10, SBC^{+}14, Sco06, Sel11, Sil09, Sin04, SA09, Sli04, Sme11, Sny07, Sok02, Spa12, Sup14, Tho04, TS14, Van11, Wil09, Zem02]. Maya [Sab07, Blu11, Bri07, Ham08a]. Maynard [Har10]. Mayr [Boc07]. Mayrhofer [Sch14]. McCarty [BK07]. McChesney [Bri00]. McElroy [Hel05]. McGehee [McK00]. McGhee [Tho07]. McKusick [Bar16, Lan16]. McLaren [Pri07]. McNamara [BL16]. Mcneil [Dun11]. Mead [Alb14]. Mean [Kat05b]. Meaning [Fel04, Wol14a]. Mechanical [Pie02]. Mechanism [Jon18]. Medal [Bur01]. Media [Jam19, Tho09]. Medical [Bar09, Bra08, Sat03]. Medicine [Cer14, Fin15, Sca09, Vac07]. Medieval [Smi04, Kit00, LB13]. Medievalizing [Zio20]. Mediterranean [War15k]. Meeting [Ano01m]. Mellink [Özy08]. Mellon [Pro07]. Melvin [OTC00]. Member [Sei01]. Members [Ano13j, Ano13p, Ano14j, Ano14p, Ano15j, Ano15p, Ano17b, Ano17d, Ano18b, Ano18d, Ano19j, Ano19l, WP15]. Membership [Coh09]. Memorial [Wol00]. Memory [Sei04]. Men [Nul14]. Menus [Gro05]. Merck [Vag01]. Meritocracy [Del12a]. Merton [Hol04]. Metaphor [Wol01b]. Metzger [Wal09]. Mexico [Ham08a, Mas16, Nor05, Reb11]. Meyer [CSG17]. Meyerson [Hac09]. Mice [Gro05]. Michael [Bla10, Ost07, Sim16, Wit20]. Michel [Mar15a]. Michelangelo [O’M12]. Michelet [Gos01]. Mid [Fin15]. Mid-Eighteenth [Fin15]. Migrants [Doy19]. Migration [Res16]. Migrations [Rei18]. Military [Hag02]. Miller [Bow11, She14, Wor18]. Minds [Isa06, O’S14]. Minus [Hei06]. Minutes [Gar16]. Miracle [Mal09]. Mirrors [Por11]. Misnomer [Sch02]. Missile [Can01]. Mission [Bor17]. Mistral [War15l]. Mixing [Coh11]. Mobility [Doy19]. Model [Doy19]. Modeling [Was06]. Models [Cos02]. Moderator [Pol04b]. Modern [Gin11, Man20, Rak19, Ros00, Tat16, Var03]. Modernity [Zio20]. Mohawk [Rei07]. Moment [War15m]. Moments [Hol16]. Money [Ger15]. Moniac [HD08]. Monroe [Aya12]. Monster [Gil08]. Montalcini [Ang19]. Moon [Wil16]. Moore [Sta05, Lea03]. Moors [Kor00]. Moral [Sat03]. Morgan [Mid15]. Morita [Pet01]. Morrison [Tsi07]. Moses [Sha05]. Mosse [Cra01]. Mote [Fon07]. Motor [SVD11]. Mount [Sto06]. Mountain [Ern00]. Mountains [Tho10, Tho17]. Movement [Sel05]. Moving [Res16]. Moynihan [Bau05]. Mozart [Lin08, Ric08]. Mud [War15n]. Mudd [Mas00]. Multidisciplinary [Yod02]. Munk [Orc19]. Murray [Bri08]. Muscle [Ste09]. Museum [Fut12, Ano19a, Ano19b, Ano19c, Ano19d, Ano20a]. Museums [De 03]. Mushat [Pel06]. Music [Ker00]. Mutations [Ste09]. Myron [Pye01]. Mysore [Kol01]. Myth [Dub11].

N [BR12, Dys06, Rak16]. Nadine [Cli15]. Name [Ham08a]. Names [Gra13]. Narasimhachar [Kol01]. NASA [Her14]. Nathans [Dan10]. Nation [HM08]. Nation-Building [HM08]. Nations [Tre19]. Native [Dar19, Pow15, Tre19]. Natural [Gos01, Sei02, Tho13, Tri17]. Nature [Bre00, Cla07, Gos01, Pec18, Rud05b, Saa16]. Neal [Bow11]. Near [Win09b]. Neel [Cro02]. Neglected [All07]. Neil [HGG13]. Neilson [Mil16]. Nelson [Wri07, Har15, Tho08]. Neo [Dys01, Kit00]. Neo-Malthusian [Dys01, Kit00]. Ness [Osm03]. Net [Ken15a]. Neugebauer [BWW16]. Neustadt [May09]. Never [Bar19]. Newcomen [Kit13]. Newell [Boy07]. News [Kov09]. Next [KRV12]. Nicholas [Gar13a]. Nicola [Hag02]. Nierenberg [DH07]. Nineteenth [Kev08]. Nirenberg [Nel11]. Nisse [Ben04]. Nitric [Mar17]. Nobel [Kir13, Nor02]. Non [Bog14]. Norman [Boy07, Fox09, Kle16]. Normative [Gib14]. Norms [Wol14a]. North [PO13, War05]. Northern [Sab07]. Not-for-Profits [Whi00]. Notre [Zio20]. Notropis [SPF03]. Novel [Dow02a, SVD11]. November [Ack04, Agr11, And00, ALSS10, BF13, BSW06, Bis05, Bla02, Bra06, Bur08, Cla00, Cli15, Con10, Con03, Cun16, Dan10, DH12, Dot05, Dow02b, Dri02, Dut16, EC02, Eas18, Ern05, Fit02, Gin10, Gou11, GK17, GC14, Hac01, Hac09, HCC06, Hin08, Höl10, Hop14, JB00, KG14, Kle16, Kol01, Kor00, Kot00, Lan05, Le 06, Mar15a, Mat01, Max07, MGM10, McC14, Mul10, Nec12, Nie00, Par06, Pea14, Phi07, Pos12, Pre05, Put05, Rai09, Ram08, Ree01, Ris05, Rot11, Sak17, Sil04, Sim11, Sli04, Spa12, Sta05, Ste08, Ste11, Ter10, Tsi07, Vog14, Wol11, Woo06a]. nozze [Lin08]. Nuclear [Fin01, Gar08b, PW01]. Nuclei [Gel01]. Number [Saa16]. Numbers [Von11].

O [Gar08a]. Oakes [Ben02]. Oberman [Pel03]. Obesity [Sug11]. Objects [Gre10a]. Obolensky [Fra04]. O’Brien [Kel01]. Observation [LHM05]. Observations [Pre00, Wol02]. Observatory [BD19]. Observed [Was17]. Occam [And02]. Occhialini [Tel02]. October [Abe03, Ada03, Agr11, Alt15, Bar16, Ber09, Bev08, Bis03, Bob07, Bri04, Bro05, BL07, Cas14, Dan10, DF04, DDV06, Eag08, Fau14, Fin10, Fis08, Fle17, Frè16, FDBH17, Gar12, Gar17, GTCD13, Gra08, GR06, Hin08, Hod12, Jam04, Jay14, Kat00, Ken03, KHEP08, Kih03, KK03, Kni08, Lan16, Lin03, Mat09, May09, MGM10, Mor06, Ost07, Özy08, Par12, Pea01, Pei04, Pel03, Pet01, Pro12, Put09, RC00, Rho09, Ric15, RB12, Sch14, Sei05, Sei08, Sha03, Shu06, Shw10, Ste08, Sud00, Tho02, Tho08, Tri05, VV11, VF08, Vog14, Wol11, Wor18, Wu07, Zuc08]. Ocular [War15o]. Odhiambo [MN11]. Odom [Mat10]. Odysseus [Men17]. Odyssey [Cer14]. off [Van18]. Offices [Ano19a, Ano19b, Ano19c, Ano19d, Ano20a]. O’Gorman [Cul17]. Oil [CHH13]. Old [Cor17, Pie02]. Oldham [JH18]. Oleg [Nec12]. Olfactory [Lea12]. Oliver [Sei08]. Olov [And03, Lin03]. Olsen [HKT14]. One [LHM05]. Opening [Jon18]. Operatic [Ful06]. Opinion [Sto15]. Oppenheimer [Hij00]. Opportunity [AP06, Pet09]. Optics [Smi04]. Option [Pie02]. Oranges [Coh11]. Organ [Sta03, SB11]. Orientalism [Mar01a]. Origin [Lo13, Tau06]. Origins [Gre19, ISO^{+}07, Kel00, Sta08]. Orman [Føl12]. Ornithology [Hal18]. Oscar [Bai13]. Oskar [Mon01]. Ostrom [Wal13]. Ostwald [Nor12]. Other [All00, All07, Ano00i, Ber14, Mar10a, Spe18, WP15]. Otis [GHX19]. Otto [FS08]. Our [Cof04, Del12a, Hir00]. Ourselves [Cos02]. Outrage [Jac20]. Overbeck [HO14]. Overgrowth [Gra14]. Overseas [Les04]. Owen [FDBH17]. Owns [Wol14b]. Owsei [Bri04]. Oxford [Mal17]. Oxide [Mar17].

P [Eag08, Fis08]. Page [Gos16]. Paine [Hof13]. Painter [Mar12]. Paintings [Rag13]. Pais [Geo07]. Pakistan [Sim09]. Palay [PRP^{+}04]. Paleoclimate [Tho10]. Paleontology [Kil04]. Palimpsest [EN10]. Palladianism [Mar10b]. Palmer [Wol04]. Pamphlet [Pre13]. Pandemic [Bar19, Tau19]. Panel [SCG^{+}12, SGTG10]. Panofsky [Gar13b]. Papers [Gre04b]. Paradox [Ken15b]. Paradoxes [Rud14, Ste04b]. Parke [Pin16]. Parker [Abe03, Wri07]. Parties [Ger15]. Passy [Mul08]. Past [Anc17, And19, Bar19, Fin12, Ker16, Lan11, Mar13, Pil18, Tho17, Wil05b]. Patents [Kev08, Mil16]. Pater [Mar12]. Patrick [Bau05]. Patron [Ron06]. Paul [FEE11, Agr11, Alb14, Alt15, Kle11, Mon01, Pro07, Rei07, Wor18, Wyl11]. Peace [Sch13]. Peale [Mar12, Mar12]. Pejorative [Cor17]. Pelikan [Mar08]. Pendleton [GR06]. Penrose [De 03, Wol00]. People [Fis19, Fut12, Res16]. Per-Olov [Lin03]. Perfect [Mun13, VF15]. Performing [Lin08]. Period [Sab07, Von11]. Persistent [McA03]. Personal [Dir10]. Persons [Men17]. Perspective [Pom08, Tho10]. Perspectives [Gar18, Sab07]. Perutz [Tho04]. Pet [Jor11]. Peter [Bla10, Not16]. Pfaffmann [Bar00]. Phases [Fin12]. Philadelphia [Cas11]. Philanthropist [Kat05b]. Philanthropy [Bre05, Kat05b, Kel00, Rim01]. Philip [MS07, Tsi07, Van11]. Philomorph [Wea01]. Philosopher [Tho13]. Philosophical [Coh09, Gre19, May01, Nas19, Pip17, Pow14, Pow15, Spe18, Van18, WP15]. Philosophy [Win19]. Phone [Jac17]. Phones [Qua13]. Phrenology [Cor17]. Phyllis [Rid04]. Phylogeny [Gar18]. Physico [Hof13]. Physico-Theology [Hof13]. Physiocratic [Hag09]. Picasso [Boi16]. Picturing [Rud05b]. Piel [DP07]. Pierce [BDN04]. Pierre [Sei05]. Pigmentation [Jab12]. Pigs [Rot12, Wei04]. Pingree [Bur08]. Piore [GG01]. Pittendrigh [KHEP08]. Pitzer [Con00a]. Place [Gin09, Meg03, Sto06]. Plain [Tat19]. Planet [Sch02]. Planner [Ron06]. Platonists [Fle15]. Platt [RK09]. Pleasures [Var07]. Pluralism [Fla05]. Plurality [Wol14a]. Plus [Hei06]. Poet [Tho13]. Poetry [Coh11, Nic05, Ven01, War15p, Yu07]. Point [HD08]. Poisoning [FH08]. Policy [Bli19, Kel00, Van07]. Political [Gui02, Jam15, Nor05, Ros00, Wei06, Win01]. Politics [Hag09, Ken15a, Ker00]. Pollak [SBC^{+}14, WH14]. Polykarp [Ram04]. Pont [Hag09]. Poor [Hua06]. Poorest [Eva10]. Pope [RC00]. Pope-Hennessy [RC00]. Popular [Mit08]. Popularisation [Bro01]. Population [Coh10, Ros00]. Porta [War15q]. Porter [Pea06, Phi05]. Portese [War15q]. Portrait [Win09b]. Possessing [Bru18]. Possibilities [EN10, Fuc01]. Possible [FH08, Lit10]. Postclassic [Sab07]. Poverty [Sch13]. Power [Gra13, Her05, Ken15a, McA03, PW01]. Powers [Tat19]. Practical [Sta03]. Practice [Wee04]. Pray [Rid04]. Prelog [Mis00]. Present [Anc17, Fin12, Her14, Mar13, Tho17, Wil05b]. Presidency [Spe18].

Presidential [GBG15]. Press [Mul08]. Previously [Coh03]. Price [CHH13, Jon18]. Primate [Tat16]. Primordial [AQ03]. Principle [Sny05]. Principles [Mil16]. Printed [Sta06]. Printing [War12]. Privacy [All07, Les07, Ros07]. Private [Mar12]. Privileged [Meg03]. Privileges [Kev08]. Prizes [Nor02]. Pro [Vir02]. Problem [Ger15, Nap05, Rub20]. Problem-Solving [Rub20]. Problems [Ber14]. Process [Nor05, Tab03]. Production [Mil16, Rut05, Sch05]. Professions [SS18]. Professor [Sei02]. Profiles [Sha05]. Profits [Whi00]. Program [Ano01m]. Propaganda [Mit08]. Property [Her05, Jac10, Kev08]. Prospect [Dys01]. Prospects [Fuc01, Rut05]. Prosperity [Sch13]. Protecting [Mor10, Ros07]. Protections [Kev08]. Protocol [WP15]. Protocols [Pow14]. Pseudo [Cor17]. Psychology [Win19]. Public [Mar12]. Publishing [Var09, War12]. Puck [Ulr20]. Pueblo [Nor05]. Pulsars [Lon11]. Punishment [Meg03, Pie03b, Ste03, Tab03]. Purchase [Kel00]. Pursuit [Nas13]. Pye [VV11]. Pyramid [Jon18].

Quality [Tho09]. Quasars [Sch11]. Quatt [Sed07]. Question [Wes11]. Quine [Føl12]. Qur’ān [AQ03, Böw03, McA03].

R [BDN04, Cot00, Fau07, MO20, Wol04]. R. [EC02, Sup14]. Rabi [Rig00]. Race [AP06, Bak10]. Radioactive [Gel01]. Raise [Fas15]. Raison [Lea12]. Ralf [Sme11]. Ralph [Jor06]. Ramsey [Kle16]. Randall [Her17]. Randolph [Sab04]. Rapid [LB13]. Rathbun [MS02]. Rather [Wol14a]. Rationality [Lo13]. Ratliff [Zem02]. Rats [War15r]. Rauch [EC02]. Raup [Kno17]. Rawls [Put05]. Ray [FDBH17]. Raymond [Ort04]. Razor [And02]. Re [Hol16]. Re-reading [Hol16]. Reading [Hor15, War15p, Hol16]. Real [Bog14, Ger15]. Reality [Dub11]. Really [Ker09, Smi04]. Reared [Seg17]. Reason [App19, Ste04b, Hof13]. Reconsidered [Hal18]. Reconstruct [ISO^{+}07]. Recorded [Ano13j, Ano14j, Ano15j, Ano17b, Ano18b, Ano19j]. Redefining [Lit10]. Redfield [Ris05]. Redington [Gar03]. Reducing [Eva10]. Redux [Nul14]. Reference [Pre13]. References [Cor17]. Reflections [Tri17, Wol01a]. Reform [Gra02a]. Refugees [Doy19]. Regenerative [Vac07]. Regge [Frè16]. Regina [Rid11]. Reiner [Rot09]. Rejection [Gil14]. Relation [Fin12]. Relations [Fas15]. Relationship [Van18]. Relevance [And19]. Religion [LJ01]. Remarkable [Pec18]. Remarks [McK04, Pie03b, Pol04b, Tat10]. Remembrance [Sei02]. Removal [Kel00]. Renaissance [RC07]. Renewal [Fis19]. Repossessing [Bru18]. Representations [SVD11]. Representing [RC07]. Reproduction [Lea12]. Republic [Sha05]. Republicanism [Mit01]. Repugnant [Fel04]. Research [Atk04, Mcc05, SCG^{+}12, Tho03, Vag01, Dar19]. Resident [Coh09]. Resources [Sel05]. Results [Bor17]. Retrospecting [Sta08]. Retroviruses [Cof04]. Return [Ano19m, Ano19n, Ano19o, Ano19p, Ano20c, Bog14]. Returns [Sie14]. Reveal [Anc17]. Review [Gre04a]. Revisited [Fin05].

Revolution [Bil01, Sei04, Swe04, Eve05, Fas15, Swe04]. Revolutionary [Hag02, Mit08]. Revolutions [Ker16]. Rhoads [Bar05, Bar03a]. Rice [Bai10]. Richard [Cul17, Ack04, Cas14, De 03, Hua06, ME13, May09, Pea01, Put09, Vog14, Wol00]. Richardson [SR00]. Rider [de 17]. Riesman [Hor14]. Right [GBG15, Men17, Tat15, Wil17]. Rights [Gre10a, Pil18]. Rings [Anc17]. Ringuet [Jac03]. Ripley [Cha03]. Riquer [Bad17]. Risk [FH08]. Risley [MN11]. Risorgimento [Gos12]. Rita [Ang19]. Ritual [Gil08]. Rivalry [Ken15a]. Road [Pal19, WH14]. Roaring [Nic05]. Robbins [Kat05a, Ste11]. Robert [Dal05, Hij00, Ber13, Bis05, BL16, Bow10, Bro09, Cho00, Dun11, Fin10, Gou11, HP06, Hol04, Jor11, May17, MS02, Oki12, Plo09, Sud00, Tip15, Wal07, Wat05, Wol04]. Roberts [Eag00b, Pea06]. Robertson [Osm03]. Robin [Jay14]. Rodgers [Ruf06]. Roland [Pel06]. Role [Eag00a, Mat07b, Rim01]. Rolihlahla [Har15]. Roman [HM08, Kin01, Mit01, Rom06]. Romare [Fin05]. Romilly [FL13]. Ronald [Tri14]. Room [Ste09]. Ropiequet [Kih03]. Rorty [Put09]. Rosenbluth [Dys06]. Rosenmeyer [Alt09]. Rosenthal [Gut05]. Rosetta [Bro02a]. Rossi [Cla00]. Rostow [Bob07]. Rough [Ern00]. Row [Pie03a]. Rowley [Olo15]. Roy [Ram08]. Royal [Ano00i, All00, Win09b]. Ruben [GG01]. Run [Pom08]. Runciman [Con03]. Rupert [Bro02b]. Rural [Car06]. Rush [Mar13]. Russell [Bar03b, KFW20, Lev07]. Russia [Gra16]. Russian [Jam19]. Russification [Wee04]. Rutherford [Osm03]. Ryland [Gar12]. Ryskamp [RB12].

S [Bar03b, Bar10b, BL16, For03, KHEP08, MLT11, Mys05, Phi07, SA09, Sok02, Tha17, Wol01a]. S. [Cha03]. Saga [Van18]. Said [Woo06a]. Saint [War15j]. Salpeter [Ter10]. Sam [Fit02]. Same [Van18]. Samuelson [Kle11]. Sanative [Fle15]. Sanders [Dri02]. Sanford [PRP^{+}04]. Sans [War15s]. Santa [Nor05]. Satellite [Par17]. Saturn [Lun09]. Saul [Beg07]. Saunders [Awo07]. Savage [Mul08]. Saxon [Atk08]. Say [Kat05b, Pre13]. Scale [Cos02]. Scales [Ven01]. Scarman [Sed07]. Schawlow [CT04]. Schlesinger [Mat09]. Schmalkald [HM08]. Scholarship [Vir02]. Schooling [HD08]. Schools [Hau20]. Schott [Reb11]. Schultz [Wha03]. Schumpeter [Nas13]. Schuylkill [Mar10b]. Sciama [Ree01]. Science [App19, Gui02, Her14, Ken15b, Nic05, Pet09, Seg17, Sei02, Sny05, Tho11, Var09]. Sciences [Car06, Yod02]. Scientific [Joh16, Kit13, Rag13, Rut05, Swe04, Tau19]. Scorn [Fis13]. Scranton [Mad15]. Sculpture [Rid05]. Sea [Cer14, Gin12]. Seaborg [Kos05, Sei04]. Seamans [Fin10]. Seamus [Ven15]. Search [Mar10a]. Sears [Mid15]. Seasons [Fut12]. Second [Gor08, Pea01]. Security [Ros07]. Sedgwick [Cul14]. Seitz [Sli10]. Self [Chu11, Col04, Nor05, Wol01a]. Self-Evident [Col04]. Self-Governance [Nor05]. Self-Image [Wol01a]. Selfish [Ken15a]. Separation [Tat19]. September [Asc11, Bad17, Bai13, BWW16, Bed03, BR12, Ben04, Ber13, Bia04, Bir15, Bis03, Bro09, Bro02b, Cai05, Cha03, Cho00, Cot00, Cra01, CD12, Dal05, DP07, DH07, Dys06, FP06, For03, FDBH17, Gar08a, Gar13b, GA06, Got08, GK17, GP02, Gut15, Hor14, Jon08, KG14, Lig17, Mas00, Mat01, McC00, McC02, Osm03, Pal03, Par12, PRP^{+}04, Pin16, PP14, Roe08, Roo15, San06, Sch14, Sei05, Slo12, Tho08, VV11, Wil09, Win09a, Woo06a, Wor18, Wyl11, dA03]. Serendipity [Mar17, Mcc05]. Sergeyevich [Mat01]. Service [Vir02]. Serviceable [All00, Ano00i]. Session [Bar18a]. Set [Pow15]. Setting [Ber14]. Sevcenko [Gra11]. Seven [Saa16, Mal17]. Severe [Sug11]. Sewell [HCC06]. Seymour [Gar17, Lak11, Sok02, Wol11]. Shackleton [Ram08]. Shadow [Ger15]. Shakespeare [Gos16]. Shannon [Gal03]. Shapiro [For03]. Shared [SB11]. Sharp [Wyn04]. Sheldon [Gut15]. Shen [Wu07]. Shiing [Wu07]. Shiing-Shen [Wu07]. Shipwrecks [Bas06]. Shiung [Lid01]. Shklar [Ben04]. Shmuel [BR12]. Shock [CHH13]. Side [Gor08]. Sight [Tat19]. Significant [Yod02]. Silhouettes [Sha05]. Sills [Bow09]. Silver [Ros07]. Silvio [Pos12]. Simon [AF03]. Simple [War12]. Simplicity [And02]. Simpson [Par06]. since [Cor17, Ham08a]. Sinfelt [TS14]. Sir [O’S14, BSW06, Bur03, Con03, Cri00, Dal05, Fra04, Hux01, Lan05, Mor13, Ort04, Osm03, Par12, Pea14, PT13, RC00, Sco06, Tho13]. Sisters [Mal17]. Six [Gar16]. Sixth [Bas06]. Sixty [Gar08b, Sim09]. Sizer [Gar12]. Skin [Jab12]. Slave [Dun07]. Slavery [Nas06]. Slow [LB13]. Smith [GH13, Gou11, Har10, Qua13]. Smyth [Cro09]. Snails [ISO^{+}07]. Snell [McK01]. Sociability [App19]. Social [Vag01]. Societal [Jac17]. Societies [Cos02, Mof20]. Society [All00, Ano00i, AP06, Hir00, Coh09, Gre19, Nas19, Pow14, Pow15, Spe18, Van18, WP15]. Software [Wol01b]. Solar [Gra15, Lea15a, VF15]. Solla [Jon18]. Solving [Rub20]. Some [Ree19, Wal06, Wol02]. Something [Cor17]. Songs [She15]. Sources [Bog14]. South [Kor00]. Space [Arm10, Can01, Fin01, Gar01, Her14, Pet09]. Space-Based [Can01]. Spain [Gil08]. Spanish [Tau06]. Speak [She15]. Special [Pre13]. Speech [Bej19]. Spilhaus [Nie00]. Spina [Adz13]. Spitzer [Ost06]. Spolia [Kin01]. Spread [Qua13, Rei18]. Square [Tal06]. Srinivas [Kol01]. Stability [Can01, Fin01]. Stage [Gos16, Ric08]. Stagnant [Ros00]. Stanley [Coo06, GK17, Nat07]. Starzl [Bar18b, Mar18]. Stassen [Kir03]. State [BD19]. States [Mil15, Reb11, Mat07b, Tab03]. Station [Pet09]. Statuary [Mar13]. Status [AP06, Fis13, SPF03]. Statutes [Fel04]. Statutory [Tat19]. Steam [Kit13]. Stebbins [May02]. Steering [Pow15]. Stem [Sca09, Wei06]. Stent [Bar10b]. Stephen [Atk08]. Steppe [Rei18]. Stern [Bia04]. Steven [Con03]. Stever [Pro12]. Stewart [Bar18a, EC02]. Stig [And03]. Still [Col04]. Stock [Sie14, Wol14b]. Stone [Dar01, Bro02a]. stonei [SPF03]. Stones [FH08]. Storm [Mun13, VF15]. Storms [Lea15a]. Strategic [Bre05, Fin01]. Strategy [Rim01]. Stream [War15i]. Strict [Lun09]. Struggle [Mal17]. Studies [Par17]. Study [Bro01, EN10, Rid05]. Subdue [O’S14]. Subjects [Gre10a]. Successful [Vag01]. Sufficiently [Col04]. Suffrage [Ulr20]. Suggests [Rei18]. Sum [Hei06]. Sun [VF06]. Sune [Ham07]. Super [Tab03]. Supreme [SGTG10]. Surgery [Adz13, Sug11]. Surgical [Tho03]. Surprise [Kir13]. Surveyor [Reb11].

Surveyors [Rom06]. Sutphin [Hog14]. Suzman [Rot11]. Sydney [Kor00]. Symposium [McK04, MB07, Pie03b, Pol04b, Was17, Wol05, Bar03a]. Syntax [Jon16]. System [Lea12]. Systems [Fin01].

T [Dot05, Fut12, Gra19, Kar05, Rho09, Sei04]. Take [Lea15b]. Tales [CHH13]. Taplin [Goh05]. Tarasca [Gil08]. Tashlich [War15t]. Taube [GC14]. Teach [Cos02]. Technical [Rut05]. Techniques [Bas06]. Technological [Kit00, Kit13]. Technologies [Sat03]. Technology [Ros07]. Telecommunications [Win15]. Telescope [Gin11]. Temkin [Bri04]. Ten [EN10]. Test [Lun09]. Their [And19, Fas15, Ham08a, Lea15a, Lea03, Mil15, Tau19, Van18]. Theodore [Gar12, Gol12, LL08, Wha03]. Theology [Hof13]. Theory [Bli19]. Things [Sab07]. Thinking [Fis19, Hei06, Tat16]. Thirsk [Ric15]. Thomas [Alt09, Bar09, Bar18b, Cha01, Coh03, Hof13, Kel00, MN11, Pin16, Pow15, Pre13, Rim08, Ron06, Sal06, Spe18, Swe04, Tho11]. Thompson [Bar10a, McC02]. Thomson [Bar18a]. Thorpe [Mid11]. Thought [Ste09]. Thoughts [De 13, Del12a, Ree19]. Thousand [Tal06]. Three [Jac20, Rem06, Rud14]. Thule [Cas11]. Thyrsus [Ols19]. Tibby [Bar03b]. Tibor [Löv19]. Till [Con00b]. Tilly [Ash09]. Time [Fut12]. Times [Blo12]. Tissue [Vac07]. Titan [Lun09]. Title [Hea04]. Tobin [Sol04]. Today [Kai07, Mar18, Mat07b, Rak19, Wes11]. Together [Seg17]. Tom [Mar18]. Toner [BL07]. top [Ano19m, Ano19n, Ano19o, Ano19p, Ano20c]. Torbjörn [KK03]. Tore [Hei18]. Townsend [Lan02]. Tradition [Hof13]. Trail [SB11]. Training [Bal12]. Transformation [Bri07]. Transforming [Sel05]. Transforms [Hir00]. Translating [Hea04]. Translation [Woo06b]. Translational [Cer14]. Transplant [Lan11, Tho03]. Transplantation [Mar18, Sta03, SB11, Wei06]. Travels [Woo06b, Yu07]. Treaties [Her05, WP15]. Treatment [Pow14]. Tree [Anc17]. Treiman [Fit02]. Trends [Dys01]. Trent [O’M12]. Trepan [Tho03]. Trials [Adz13]. Tribulations [Adz13]. Tribute [Fut12]. Triumph [Ful06, Sta03]. Trojan [Wei04]. Trolls [Jam19]. True [Cor17, War05]. Truman [Joh06]. Truths [Col04]. Tsunami [Key03]. Tukey [Mos05]. Tullio [Frè16]. Twain [Mid06]. Twelfth [War15u]. Twentieth [Was06]. Twentieth- [Was06]. Twenty [Plo18, Ste03, Von11, RC07, Rid05, Was06]. Twenty-First [Ste03, Rid05]. Twenty-First-Century [RC07, Was06]. Twenty-Five-Year [Von11]. Twins [Seg17]. Two [Bej19, CHH13, Gre04a, Plo18, Syl14, Van18]. Tygers [Nic05]. Type [Pre13].

U.S. [Jam19]. Ubiquity [Lun09]. Ulam [Gle02]. Ultima [Cas11]. Underdeveloped [Whi00]. Undergrowth [Gra14]. Underrated [Mit01]. Understanding [Kra04, Mas07, Tho10]. Unfair [Tab03]. Unfortunately [Col04]. Uniqueness [Poc01]. United} [Ger15, Mat07b, Mil15, Reb11, Tab03]. Universal [Pal19]. Universe [Kir13, Ste04a, Mar10a]. University [Atk04, Dud01, Dir10, SCG^{+}12].

Unknowingness [Pip17]. Unleashing [NM05]. Unpublished [Coh03]. Unshackled [Ker16]. Upward [Fis13]. Urbanization [LHM05]. Use [Odo07]. Useful [Dar18, SGTG10]. Uses [Blo12, Pye02]. Using [ISO^{+}07]. USSR [Gui02].

V [Cro02, Kra04, Les04, McK04, Pol04b, Pol04a, WH14]. Vaccination [Plo18]. Vaccine [Hil00, Pal19]. Values [Arr14, Fla05, Wol14a]. Vann [Hac01]. Variation [SPF03]. Various [Ven01]. Varnedoe [Kar05]. Vatican [Vir02]. Venomous [ISO^{+}07]. Verdi [Gos12]. Verification [Sil16]. Vermeule [Lan02]. Vertigo [Pip17]. Vest [Bir15]. Viagra [Mar17]. Vichy [Ful06]. Victor [Bar16, Got08]. View [Arr14, Kai07, Ulr20, Van07]. Vincent [Lan16]. Virgil [Put01]. Virginia [Dun07]. virorum [Vir02]. Virulence [Tau06]. Virus [Kil04, Pal19, Tau06]. Viruses [Vil04]. Vision [Zew06]. Visits [Cas11]. Vitalii [DT03]. Vive [Eas00]. Vladimir [Mis00, Yak12]. Vlastos [DF04]. Vogt [Bri06]. Voice [Böw03, Coo03, Tre19]. Voidness [Fel04]. Volume [Ano00n, Ano01n, Ano02m, Ano03m, Ano04m, Ano05m, Ano06m, Ano07m, Ano08m, Ano13i, Ano13o, Ano14i, Ano14o, Ano15i, Ano15o, Ano17a, Ano17c, Ano18a, Ano18c, Ano19i, Ano19k]. Vote [GBG15, Tat15]. Voting [Pil18].

W [Cas14, Eas18, McC16, Mei19, Mos05, Rim08, VV11, Wha03]. W. [Eng01]. Wade [Fon07]. Wagley [Kot00]. Wald [Dow02b]. Walkabout [Bea04]. Walt [Bob07]. Walter [Kot00, Aya12, Cun16, Jam04, Orc19, SG18]. Walton [Fle17]. Wampum [Bru18]. War [Gly16, Ken15a, O’S14, Win15, Hag02]. Ward [Oki12, Mar15b]. Warming [Ken15b]. Warren [Mad15, Nel11]. Was [Mar09, Men17]. Washington [Cas11, Hag02]. Wassily [Ber00]. Water [Eag00a]. Way [Hei06, War15v]. Wayne [Bev08]. Weapons [Gar01, Gar08b]. Weber [Hun09]. Wechsler [Hen02]. Weevils [Pre13]. Weigley [Lev07]. Weil [Bor01]. Weinberg [Zuc08]. Weiner [Pye01]. Weisskopf [Got08]. Welcome [San06]. Welfare [Smi01]. Well [Mar18]. Wells [Wal02]. Were [Col04, Nas19]. Wesley [VW12]. West [Lig17, HD08, Mar01a, Ulr20]. Western [Ham08a]. Westheimer [Fis09]. Wetherill [All09]. Whales [Gin12]. What/When [Win09b]. Wheeler [Dys10]. Whig [App19]. Whipple [Gin10]. White [Bak10, Hin08]. Whitfield [SCL^{+}10]. Who [Lea15b, Wol14b]. Wife [Fin15]. Wild [FEE11]. Will [Fur15, Lea15b, Ste04b]. Willard [Kat00, Føl12]. Willey [Sab04]. William [App19, Bra06, Bri00, DH07, Fla05, Fut12, Gar03, Gra02b, HCC06, Hel05, Ker00, Mad15, Mar13, Mat10, Nic05, Rho09, RK09, Sei08, Slo12, Wri07]. Williams [Jay14, Sha05]. Willson [Mar12]. Wilson [Dal05, MS02, Ken15a, Lea15b]. Window [Gel01]. Wine [Sto15]. Within [Res16]. Witkop [Spa12]. Wolf [Jor11]. Wolfgang [Sil04, Gar13b]. Woman [Ulr20]. Women [Gly16, Mal17]. Woodrow [Ken15a, Lea15b]. Woodward [Hac01]. Woolf [Wil05a]. Word [Wee04]. Words [VF10]. Work [Mar10b]. Works [Win19]. World [Dys01, Fas15, Her05, Kai07, Mar01b, Mat07b, NM05, Sch05, Tho10, Tho17, Wol00, Hag09, Ken15a, Win15]. Worms [FL13]. Worth [Gut14, SR00]. Wright [Deg02, Mul10]. Writing [Bri07, Sta06, Sto06]. Wu [Lid01, Sei02]. Wyeth [Sto11].

XenoLungs [Rot12]. Xenotransplantation [Wei04].

Year [Cer14, Van18, Von11]. Years [EN10, Gar08b, Gre04a, LHM05, Plo18, Sim09]. Yoder [Mys05]. Yorkshire [Kor00]. Young [Cas11, Nas19].

Zamecnik [Agr11]. Zartman [Bri06]. Zero [Blu11, Hei06]. Zero-Sum [Hei06]. Zewail [Tho18].

References

Abelson:2003:CPH

[Abe03] Philip Hauge Abelson. Caryl Parker Haskins, 12 August 1908–8 October 2001. Proceedings of the American Philosophical Society, 147(4):387–392, December 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

        stable/1558301.

Ackerman:2004:RKJ

[Ack04] James S. Ackerman. Richard Krautheimer, 6 July 1897–1 November 1994. Proceedings of the American Philosophical Society, 148(2):229– 234, June 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

        9243 (electronic). URL https://www.jstor.org/stable/1558288.

Adams:2000:LHJ

[Ada00] Arlin M. Adams. A. Leon HigginBotham, Jr., 25 February 1928– 14 December 1998. Proceedings of the American Philosophical Society, 144(2):219–223, June 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

        stable/1515635.

Adams:2003:MHM

[Ada03] Arlin M. Adams. Maurice Heckscher, 24 May 1907–31 October 2001. Proceedings of the American Philosophical Society, 147(4):393–397, December 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

        9243 (electronic). URL https://www.jstor.org/stable/1558302.

Adams:2005:CIL

[Ada05] Robert McC. Adams. Contexts of Iraqi looting. Proceedings of the American Philosophical Society, 149(1):56–62, March 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/4598908.

Adzick:2013:TTF

[Adz13] N. Scott Adzick. The tribulations and trials of fetal surgery for spina bifida. Proceedings of the American Philosophical Society, 157(4):409–437, December 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640225.

Augier:2003:HSJ

[AF03] Mie Augier and Edward Feigenbaum. Herbert A. Simon, 15 June 1916–9 February 2001. Proceedings of the American Philosophical Society, 147(2):193–198, June 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558264.

Agrawal:2011:PCZ

[Agr11] Sudhir Agrawal. Paul C. Zamecnik: 22 November 1912–27 October 2009. Proceedings of the American Philosophical Society, 155(2):241– 246, June 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23056841.

Alberts:2014:PMD

[Alb14] Bruce Alberts. Paul Mead Doty: 1 June 1920–5 December 2011. Proceedings of the American Philosophical Society, 158 (4):423–428, December 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Doty.pdf;   https://www.jstor.org/stable/24640186.

Allan:2000:DSE

[All00] D. G. C. Allan. “Dear and serviceable to each other”: Benjamin Franklin and the Royal Society of Arts. Proceedings of the American Philosophical Society, 144(3):245–266, September 2000. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/1515588. See correction to caption of fig-

ure 2 [Ano00i].

Allen:2007:FFI

[All07] Anita L. Allen. Face to face with “it”: And other neglected contexts of health privacy. Proceedings of the American Philosophical Society, 151(3):300–308, September 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599073.

Allegre:2009:GWA

[All09] Claude Allègre. George Wetherill 12 August 1925–19 July 2006. Proceedings of the American Philosophical Society, 153(1):127–130, March 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/40541637.

Aronson:2010:GLN

[ALSS10] Elliot Aronson, R. Duncan Luce, Claude Steele, and Patrick Suppes. Gardner Lindzey, 27 November 1920–4 February 2008. Proceedings of the American Philosophical Society, 154(1):103–107, March 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/20721531.

Alter:2009:TGR

[Alt09] Robert Alter. Thomas Gustav Rosenmeyer 3 April 1920–6 February 2007. Proceedings of the American Philosophical Society, 153(2):233– 237, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541664.

Altman:2015:PJA

[Alt15] Joel B. Altman. Paul Joel Alpers: 16 October 1932–19 May 2013. Proceedings of the American Philosophical Society, 159(1):113–117, March 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Alpers.pdf;   https://
www.jstor.org/stable/24640175.

Anchukaitis:2017:TRR

[Anc17] Kevin J. Anchukaitis. Tree rings reveal climate change past, present, and future. Proceedings of the American Philosophical Society, 161(3):244–263, September 2017. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://pdfs.

semanticscholar.org/a012/745a389e09d961ba4192fe68ae4a9362aed2.
pdf;   https://www.amphilsoc.org/sites/default/files/2017-
12/attachments/Anchukaitis.pdf.

Andler:2000:JLN

[And00] Martin Andler. Jean Leray, 7 November 1906–10 November 1998. Proceedings of the American Philosophical Society, 144(4):469–478, December 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1515623.

Anderson:2002:ORS

[And02] Don L. Anderson. Occam’s Razor: Simplicity, complexity, and global geodynamics. Proceedings of the American Philosophical Society, 146(1):56–76, March 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558156.

Anderson:2003:SOL

[And03] Philip W. Anderson. Stig Olov Lundquist, 9 August 1925–6 April 2000. Proceedings of the American Philosophical Society, 147(3): 287–291, September 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558217.

Anderson:2019:GCP

[And19] Margo Anderson. The ghosts of census past and their relevance for 2020. Proceedings of the American Philosophical Society, 163(3): ??, September 2019. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2020-08/attachments/Anderson.pdf.

Angeletti:2019:RLM

[Ang19] Ruth H. Angeletti. Rita Levi-Montalcini. Proceedings of the American Philosophical Society, 163(4):??, December 2019. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2021-
01/attachments/Levi-Montalcini.pdf.

Anonymous:2000:Aa

[Ano00a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 144(1):??, March 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2000:Ab

[Ano00b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 144(2):??, June 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2000:Ac

[Ano00c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 144(3):??, September 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2000:Ad

[Ano00d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 144(4):??, December 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2000:BMa

[Ano00e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 144(1):??, March 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515613.

Anonymous:2000:BMb

[Ano00f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 144(2):??, June 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515638.

Anonymous:2000:BMc

[Ano00g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 144(3):??, September 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515600.

Anonymous:2000:BMd

[Ano00h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 144(4):??, December 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515627.

Anonymous:2000:CDS

[Ano00i] Anonymous. Correction: “Dear and Serviceable to Each Other”: Benjamin Franklin and the Royal Society of Arts. Proceedings of the American Philosophical Society, 144(4):504, December 2000. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1515626. See [All00].

Anonymous:2000:FMa

[Ano00j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 144(1):83, March 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515601.

Anonymous:2000:FMb

[Ano00k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 144(2):199, June 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515628.

Anonymous:2000:FMc

[Ano00l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 144(3):289, September 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515586.

Anonymous:2000:FMd

[Ano00m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 144(4):451–496, December 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/1515615.

Anonymous:2000:VI

[Ano00n] Anonymous. Volume information. Proceedings of the American Philosophical Society, 144(4):497–503, December 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1515614.

Anonymous:2001:Aa

[Ano01a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 145(1):??, March 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2001:Ab

[Ano01b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 145(2):??, June 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2001:Ac

[Ano01c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 145(3):??, September 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2001:Ad

[Ano01d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 145(4):??, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2001:BMa

[Ano01e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 145(1):??, March 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558331.

Anonymous:2001:BMb

[Ano01f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 145(2):??, June 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558280.

Anonymous:2001:BMc

[Ano01g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 145(3):??, September 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558120.

Anonymous:2001:BMd

[Ano01h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 145(4):596–598, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/1558197.

Anonymous:2001:FMa

[Ano01i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 145(1):81, March 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558322.

Anonymous:2001:FMb

[Ano01j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 145(2):187, June 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558266.

Anonymous:2001:FMc

[Ano01k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 145(3):335, September 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558106.

Anonymous:2001:FMd

[Ano01l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 145(4):i–ii, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558178.

Anonymous:2001:PM

[Ano01m] Anonymous. Program of the meeting. Proceedings of the American Philosophical Society, 145(4):385–388, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1558179.

Anonymous:2001:VI

[Ano01n] Anonymous. Volume information. Proceedings of the American Philosophical Society, 145(4):599–605, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1558177.

Anonymous:2002:Aa

[Ano02a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 146(1):??, March 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2002:Ab

[Ano02b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 146(2):??, June 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2002:Ac

[Ano02c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 146(3):??, September 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2002:Ad

[Ano02d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 146(4):??, December 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2002:BMa

[Ano02e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 146(1):??, March 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558164.

Anonymous:2002:BMb

[Ano02f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 146(2):??, June 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558208.

Anonymous:2002:BMc

[Ano02g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 146(3):??, September 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558176.

Anonymous:2002:BMd

[Ano02h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 146(4):440–442, December 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/1558321.

Anonymous:2002:FMa

[Ano02i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 146(1):i–99, March 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558152.

Anonymous:2002:FMb

[Ano02j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 146(2):193, June 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558198.

Anonymous:2002:FMc

[Ano02k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 146(3):283, September 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558165.

Anonymous:2002:FMd

[Ano02l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 146(4):i–385, December 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558308.

Anonymous:2002:VI

[Ano02m] Anonymous. Volume information. Proceedings of the American Philosophical Society, 146(4):443–449, December 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1558307.

Anonymous:2003:Aa

[Ano03a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 147(1):??, March 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2003:Ab

[Ano03b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 147(2):??, June 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2003:Ac

[Ano03c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 147(3):??, September 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2003:Ad

[Ano03d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 147(4):??, December 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2003:BMa

[Ano03e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 147(1):??, March 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558133.

Anonymous:2003:BMb

[Ano03f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 147(2):??, June 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558265.

Anonymous:2003:BMc

[Ano03g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 147(3):??, September 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558223.

Anonymous:2003:BMd

[Ano03h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 147(4):417–419, December 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/1558306.

Anonymous:2003:FMa

[Ano03i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 147(1):i–65, March 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558121.

Anonymous:2003:FMb

[Ano03j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 147(2):i–151, June 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558252.

Anonymous:2003:FMc

[Ano03k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 147(3):i–273, September 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/1558209.

Anonymous:2003:FMd

[Ano03l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 147(4):i–377, December 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558294.

Anonymous:2003:VI

[Ano03m] Anonymous. Volume information. Proceedings of the American Philosophical Society, 147(4):420–426, December 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1558293.

Anonymous:2004:Aa

[Ano04a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 148(1):??, March 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2004:Ab

[Ano04b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 148(2):??, June 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2004:Ac

[Ano04c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 148(3):??, September 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2004:Ad

[Ano04d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 148(4):??, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2004:BMa

[Ano04e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 148(1):??, March 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558251.

Anonymous:2004:BMb

[Ano04f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 148(2):??, June 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558292.

Anonymous:2004:BMc

[Ano04g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 148(3):??, September 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558238.

Anonymous:2004:BMd

[Ano04h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 148(4):555–557, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/1558151.

Anonymous:2004:FMa

[Ano04i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 148(1):i–121, March 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558239.

Anonymous:2004:FMb

[Ano04j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 148(2):213, June 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558281.

Anonymous:2004:FMc

[Ano04k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 148(3):i–359, September 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/1558224.

Anonymous:2004:FMd

[Ano04l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 148(4):i–491, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558135.

Anonymous:2004:VI

[Ano04m] Anonymous. Volume information. Proceedings of the American Philosophical Society, 148(4):558–565, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1558134.

Anonymous:2005:Aa

[Ano05a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 149(1):??, March 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2005:Ab

[Ano05b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 149(2):??, June 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2005:Ac

[Ano05c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 149(3):??, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2005:Ad

[Ano05d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 149(4):??, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2005:BMa

[Ano05e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 149(1):??, March 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598919.

Anonymous:2005:BMb

[Ano05f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 149(2):??, June 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598934.

Anonymous:2005:BMc

[Ano05g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 149(3):??, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598951.

Anonymous:2005:BMd

[Ano05h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 149(4):643–645, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/4598969.

Anonymous:2005:FMa

[Ano05i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 149(1):i–81, March 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598904.

Anonymous:2005:FMb

[Ano05j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 149(2):i–227, June 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598920.

Anonymous:2005:FMc

[Ano05k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 149(3):i–387, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/4598935.

Anonymous:2005:FMd

[Ano05l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 149(4):i–591, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598953.

Anonymous:2005:VI

[Ano05m] Anonymous. Volume information. Proceedings of the American Philosophical Society, 149(4):646–654, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/4598952.

Anonymous:2006:Aa

[Ano06a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 150(1):??, March 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2006:Ab

[Ano06b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 150(2):??, June 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2006:Ac

[Ano06c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 150(3):??, September 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2006:Ad

[Ano06d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 150(4):??, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2006:BMa

[Ano06e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 150(1):??, March 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598986.

Anonymous:2006:BMb

[Ano06f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 150(2):??, June 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599002.

Anonymous:2006:BMc

[Ano06g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 150(3):??, September 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599019.

Anonymous:2006:BMd

[Ano06h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 150(4):703–705, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/4599038.

Anonymous:2006:FMa

[Ano06i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 150(1):i–149, March 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598970.

Anonymous:2006:FMb

[Ano06j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 150(2):i–333, June 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598987.

Anonymous:2006:FMc

[Ano06k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 150(3):i–455, September 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/4599003.

Anonymous:2006:FMd

[Ano06l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 150(4):i–647, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599021.

Anonymous:2006:VI

[Ano06m] Anonymous. Volume information. Proceedings of the American Philosophical Society, 150(4):706–714, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/4599020.

Anonymous:2007:Aa

[Ano07a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 151(1):??, March 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2007:Ab

[Ano07b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 151(2):??, June 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2007:Ac

[Ano07c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 151(3):??, September 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2007:Ad

[Ano07d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 151(4):??, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2007:BMa

[Ano07e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 151(1):??, March 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599053.

Anonymous:2007:BMb

[Ano07f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 151(2):??, June 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599069.

Anonymous:2007:BMc

[Ano07g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 151(3):??, September 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599084.

Anonymous:2007:BMd

[Ano07h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 151(4):??, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478465.

Anonymous:2007:FMa

[Ano07i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 151(1):i–79, March 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599039.

Anonymous:2007:FMb

[Ano07j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 151(2):i–231, June 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599054.

Anonymous:2007:FMc

[Ano07k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 151(3):i–329, September 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/4599070.

Anonymous:2007:FMd

[Ano07l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 151(4):??, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478449.

Anonymous:2007:VI

[Ano07m] Anonymous. Volume information. Proceedings of the American Philosophical Society, 151(4):??, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/25478448.

Anonymous:2008:Aa

[Ano08a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 152(1):??, March 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2008:Ab

[Ano08b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 152(2):??, June 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2008:Ac

[Ano08c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 152(3):??, September 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2008:Ad

[Ano08d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 152(4):??, December 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2008:BMa

[Ano08e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 152(1):??, March 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478481.

Anonymous:2008:BMb

[Ano08f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 152(2):??, June 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478498.

Anonymous:2008:BMc

[Ano08g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 152(3):??, September 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541598.

Anonymous:2008:BMd

[Ano08h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 152(4):??, December 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541613.

Anonymous:2008:FMa

[Ano08i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 152(1):??, March 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478467.

Anonymous:2008:FMb

[Ano08j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 152(2):??, June 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478483.

Anonymous:2008:FMc

[Ano08k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 152(3):??, September 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541588.

Anonymous:2008:FMd

[Ano08l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 152(4):??, December 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541600.

Anonymous:2008:VI

[Ano08m] Anonymous. Volume information. Proceedings of the American Philosophical Society, 152(4):??, December 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/40541599.

Anonymous:2009:Aa

[Ano09a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 153(1):??, March 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2009:Ab

[Ano09b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 153(2):??, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2009:Ac

[Ano09c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 153(3):??, September 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2009:Ad

[Ano09d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 153(4):??, December 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2009:BMa

[Ano09e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 153(1):??, March 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541638.

Anonymous:2009:BMb

[Ano09f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 153(2):??, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541667.

Anonymous:2009:BMc

[Ano09g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 153(3):??, September 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541681.

Anonymous:2009:BMd

[Ano09h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 153(4):??, December 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/20721519.

Anonymous:2009:FMa

[Ano09i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 153(1):??, March 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541626.

Anonymous:2009:FMb

[Ano09j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 153(2):??, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541651.

Anonymous:2009:FMc

[Ano09k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 153(3):??, September 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541668.

Anonymous:2009:FMd

[Ano09l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 153(4):??, December 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/20721508.

Anonymous:2010:Aa

[Ano10a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 154(1):??, March 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2010:Ab

[Ano10b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 154(2):??, June 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2010:Ac

[Ano10c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 154(3):??, September 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2010:Ad

[Ano10d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 154(4):??, December 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2010:BMa

[Ano10e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 154(1):??, March 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/20721536.

Anonymous:2010:BMb

[Ano10f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 154(2):??, June 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/41000101.

Anonymous:2010:BMc

[Ano10g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 154(3):??, September 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/41000092.

Anonymous:2010:BMd

[Ano10h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 154(4):??, December 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056871.

Anonymous:2010:FMa

[Ano10i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 154(1):??, March 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/20721521.

Anonymous:2010:FMb

[Ano10j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 154(2):??, June 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/41000093.

Anonymous:2010:FMc

[Ano10k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 154(3):??, September 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/41000079.

Anonymous:2010:FMd

[Ano10l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 154(4):??, December 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056858.

Anonymous:2011:Aa

[Ano11a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 155(1):??, March 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2011:Ab

[Ano11b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 155(2):??, June 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2011:Ac

[Ano11c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 155(3):??, September 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2011:Ad

[Ano11d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 155(4):??, December 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2011:BMa

[Ano11e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 155(1):??, March 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056856.

Anonymous:2011:BMb

[Ano11f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 155(2):??, June 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056842.

Anonymous:2011:BMc

[Ano11g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 155(3):??, September 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23208696.

Anonymous:2011:BMd

[Ano11h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 155(4):??, December 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23208790.

Anonymous:2011:FMa

[Ano11i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 155(1):??, March 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056843.

Anonymous:2011:FMb

[Ano11j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 155(2):??, June 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056828.

Anonymous:2011:FMc

[Ano11k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 155(3):??, September 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23208682.

Anonymous:2011:FMd

[Ano11l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 155(4):??, December 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23208779.

Anonymous:2012:Aa

[Ano12a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 156(1):??, March 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2012:Ab

[Ano12b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 156(2):??, June 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2012:Ac

[Ano12c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 156(3):??, September 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2012:Ad

[Ano12d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 156(4):??, December 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2012:BMa

[Ano12e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 156(1):??, March 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558086.

Anonymous:2012:BMb

[Ano12f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 156(2):??, June 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558201.

Anonymous:2012:BMc

[Ano12g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 156(3):??, September 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558099.

Anonymous:2012:BMd

[Ano12h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 156(4):??, December 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558241.

Anonymous:2012:FMa

[Ano12i] Anonymous. Front matter. Proceedings of the American Philosophical Society, 156(1):??, March 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558074.

Anonymous:2012:FMb

[Ano12j] Anonymous. Front matter. Proceedings of the American Philosophical Society, 156(2):??, June 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558189.

Anonymous:2012:FMc

[Ano12k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 156(3):??, September 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558087.

Anonymous:2012:FMd

[Ano12l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 156(4):??, December 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558229.

Anonymous:2013:Aa

[Ano13a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 157(1):??, March 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2013:Ab

[Ano13b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 157(2):??, June 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2013:Ac

[Ano13c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 157(3):??, September 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2013:Ad

[Ano13d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 157(4):??, December 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2013:BMa

[Ano13e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 157(1):??, March 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558147.

Anonymous:2013:BMb

[Ano13f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 157(2):??, June 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640246.

Anonymous:2013:BMc

[Ano13g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 157(3):??, September 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640403.

Anonymous:2013:BMd

[Ano13h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 157(4):??, December 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640235.

Anonymous:2013:CV

[Ano13i] Anonymous. Contents of volume 157. Proceedings of the American Philosophical Society, 157(4):491–495, December 2013. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/24640233.

Anonymous:2013:DMR

[Ano13j] Anonymous. Deaths of members recorded through 31 December 2013. Proceedings of the American Philosophical Society, 157(4):489–490, December 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/24640232.

Anonymous:2013:FMa

[Ano13k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 157(1):??, March 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558136.

Anonymous:2013:FMb

[Ano13l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 157(2):??, June 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640237.

Anonymous:2013:FMc

[Ano13m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 157(3):??, September 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640394.

Anonymous:2013:FMd

[Ano13n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 157(4):??, December 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640222.

Anonymous:2013:IV

[Ano13o] Anonymous. Index to volume 157. Proceedings of the American Philosophical Society, 157(4):496–501, December 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/24640234.

Anonymous:2013:ME

[Ano13p] Anonymous. Members elected in 2013. Proceedings of the American Philosophical Society, 157(4):487–488, December 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/24640231.

Anonymous:2014:Aa

[Ano14a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 158(1):??, March 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2014:Ab

[Ano14b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 158(2):??, June 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2014:Ac

[Ano14c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 158(3):??, September 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2014:Ad

[Ano14d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 158(4):??, December 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2014:BMa

[Ano14e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 158(1):??, March 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640148.

Anonymous:2014:BMb

[Ano14f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 158(2):??, June 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640208.

Anonymous:2014:BMc

[Ano14g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 158(3):??, September 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640137.

Anonymous:2014:BMd

[Ano14h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 158(4):??, December 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640196.

Anonymous:2014:CV

[Ano14i] Anonymous. Contents of volume 158. Proceedings of the American Philosophical Society, 158(4):465–470, December 2014. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/24640194.

Anonymous:2014:DMR

[Ano14j] Anonymous. Deaths of members recorded through 31 December 2014. Proceedings of the American Philosophical Society, 158(4):463–464, December 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/24640193.

Anonymous:2014:FMa

[Ano14k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 158(1):??, March 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640138.

Anonymous:2014:FMb

[Ano14l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 158(2):??, June 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640197.

Anonymous:2014:FMc

[Ano14m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 158(3):??, September 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640124.

Anonymous:2014:FMd

[Ano14n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 158(4):??, December 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640178.

Anonymous:2014:IV

[Ano14o] Anonymous. Index to volume 158. Proceedings of the American Philosophical Society, 158(4):471–476, December 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/24640195.

Anonymous:2014:ME

[Ano14p] Anonymous. Members elected in 2014. Proceedings of the American Philosophical Society, 158(4):461–462, December 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/24640192.

Anonymous:2015:Aa

[Ano15a] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 159(1):??, March 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2015:Ab

[Ano15b] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 159(2):??, June 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2015:Ac

[Ano15c] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 159(3):??, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2015:Ad

[Ano15d] Anonymous. Accessibility. Proceedings of the American Philosophical Society, 159(4):??, December 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Anonymous:2015:BMa

[Ano15e] Anonymous. Back matter. Proceedings of the American Philosophical Society, 159(1):??, March 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640177.

Anonymous:2015:BMb

[Ano15f] Anonymous. Back matter. Proceedings of the American Philosophical Society, 159(2):??, June 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640221.

Anonymous:2015:BMc

[Ano15g] Anonymous. Back matter. Proceedings of the American Philosophical Society, 159(3):??, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/26159164.

Anonymous:2015:BMd

[Ano15h] Anonymous. Back matter. Proceedings of the American Philosophical Society, 159(4):??, December 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/26159205.

Anonymous:2015:CV

[Ano15i] Anonymous. Contents of volume 159. Proceedings of the American Philosophical Society, 159(4):487–492, December 2015. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/26159203.

Anonymous:2015:DMR

[Ano15j] Anonymous. Deaths of members recorded through 31 December 2015. Proceedings of the American Philosophical Society, 159(4):485–486, December 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/26159202.

Anonymous:2015:FMa

[Ano15k] Anonymous. Front matter. Proceedings of the American Philosophical Society, 159(1):??, March 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640165.

Anonymous:2015:FMb

[Ano15l] Anonymous. Front matter. Proceedings of the American Philosophical Society, 159(2):??, June 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640209.

Anonymous:2015:FMc

[Ano15m] Anonymous. Front matter. Proceedings of the American Philosophical Society, 159(3):??, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/26159133.

Anonymous:2015:FMd

[Ano15n] Anonymous. Front matter. Proceedings of the American Philosophical Society, 159(4):??, December 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/26159190.

Anonymous:2015:IV

[Ano15o] Anonymous. Index to volume 159. Proceedings of the American Philosophical Society, 159(4):493–498, December 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/26159204.

Anonymous:2015:ME

[Ano15p] Anonymous. Members elected in 2015. Proceedings of the American Philosophical Society, 159(4):483–484, December 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/26159201.

Anonymous:2017:CV

[Ano17a] Anonymous. Contents of volume 161. Proceedings of the American Philosophical Society, 161(4):367–371, ???? 2017. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2018-
03/attachments/Contents%20of%20Volume%20161.pdf.

Anonymous:2017:DMR

[Ano17b] Anonymous. Deaths of members recorded through 31 December 2017. Proceedings of the American Philosophical Society, 161(4): 365–366, ???? 2017. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2018-03/attachments/Deaths%20of%20Members%
20Recorded%20through%2031%20December%202017.pdf.

Anonymous:2017:IV

[Ano17c] Anonymous. Index to volume 161. Proceedings of the American Philosophical Society, 161(4):372–376, ???? 2017. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2018-
03/attachments/Index%20to%20Volume%20161.pdf.

Anonymous:2017:ME

[Ano17d] Anonymous. Members elected in 2017. Proceedings of the American Philosophical Society, 161(4):363, ???? 2017. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2018-
03/attachments/Members%20Elected%20in%202017.pdf.

Anonymous:2018:CV

[Ano18a] Anonymous. Contents of volume 162. Proceedings of the American Philosophical Society, 162(4):405–408, ???? 2018. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2019-
06/attachments/Contents%20of%20Volume%20162.pdf.

Anonymous:2018:DMR

[Ano18b] Anonymous. Deaths of members recorded through 31 December 2018. Proceedings of the American Philosophical Society, 162(4): 403–404, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2019-06/attachments/Deaths%20of%20Members%
202018.pdf.

Anonymous:2018:IV

[Ano18c] Anonymous. Index to volume 162. Proceedings of the American Philosophical Society, 162(4):409–413, ???? 2018. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2019-
06/attachments/Index%20to%20Volume%20162.pdf.

Anonymous:2018:ME

[Ano18d] Anonymous. Members elected in 2018. Proceedings of the American Philosophical Society, 162(4):401–402, ???? 2018. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2019-
06/attachments/Members%20Elected%20in%202018.pdf.

Anonymous:2019:AAOa

[Ano19a] Anonymous. APS Administrative Offices & APS Museum. Proceedings of the American Philosophical Society, 163(1):??, March 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL /visit-museum.

Anonymous:2019:AAOb

[Ano19b] Anonymous. APS Administrative Offices & APS Museum. Proceedings of the American Philosophical Society, 163(2):??, June 2019. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

/visit-museum.

Anonymous:2019:AAOc

[Ano19c] Anonymous. APS Administrative Offices & APS Museum. Proceedings of the American Philosophical Society, 163(3):??, September 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL /visit-museum.

Anonymous:2019:AAOd

[Ano19d] Anonymous. APS Administrative Offices & APS Museum. Proceedings of the American Philosophical Society, 163(4):??, December 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL /visit-museum.

Anonymous:2019:ALa

[Ano19e] Anonymous. APS Library. Proceedings of the American Philosophical Society, 163(1):??, March 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL /library.

Anonymous:2019:ALb

[Ano19f] Anonymous. APS Library. Proceedings of the American Philosophical Society, 163(2):??, June 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL /library.

Anonymous:2019:ALc

[Ano19g] Anonymous. APS Library. Proceedings of the American Philosophical Society, 163(3):??, September 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL /library.

Anonymous:2019:ALd

[Ano19h] Anonymous. APS Library. Proceedings of the American Philosophical Society, 163(4):??, December 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL /library.

Anonymous:2019:CV

[Ano19i] Anonymous. Contents of volume 163. Proceedings of the American Philosophical Society, 163(4):??, December 2019. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2021-
01/attachments/Contents%20of%20Volume%20163.pdf.

Anonymous:2019:DMR

[Ano19j] Anonymous. Deaths of members recorded through 31 December 2019. Proceedings of the American Philosophical Society, 163(4): ??, December 2019. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2021-01/attachments/Deaths%20of%20Members%
20Recorded%20through%2031%20December%202019.pdf.

Anonymous:2019:IV

[Ano19k] Anonymous. Index to volume 163. Proceedings of the American Philosophical Society, 163(4):??, December 2019. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2021-
01/attachments/Index%20to%20Volume%20163.pdf.

Anonymous:2019:ME

[Ano19l] Anonymous. Members elected in 2019. Proceedings of the American Philosophical Society, 163(4):??, December 2019. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2021-
01/attachments/Members%20Elected%20in%202019.pdf.

Anonymous:2019:RTa

[Ano19m] Anonymous. Return to top. Proceedings of the American Philosophical Society, 163(1):??, March 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL #block-sitebranding.

Anonymous:2019:RTb

[Ano19n] Anonymous. Return to top. Proceedings of the American Philosophical Society, 163(2):??, June 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL #block-sitebranding.

Anonymous:2019:RTc

[Ano19o] Anonymous. Return to top. Proceedings of the American Philosophical Society, 163(3):??, September 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL #block-

sitebranding.

Anonymous:2019:RTd

[Ano19p] Anonymous. Return to top. Proceedings of the American Philosophical Society, 163(4):??, December 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL #block-

sitebranding.

Anonymous:2020:AAO

[Ano20a] Anonymous. APS Administrative Offices & APS Museum. Proceedings of the American Philosophical Society, 164(1):??, March 2020. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL /visit-museum.

Anonymous:2020:AL

[Ano20b] Anonymous. APS Library. Proceedings of the American Philosophical Society, 164(1):??, March 2020. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL /library.

Anonymous:2020:RT

[Ano20c] Anonymous. Return to top. Proceedings of the American Philosophical Society, 164(1):??, March 2020. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL #block-sitebranding.

Atkinson:2006:ODS

[AP06] Richard C. Atkinson and Patricia A. Pelfrey. Opportunity in a democratic society: Race and economic status in higher education. Proceedings of the American Philosophical Society, 150(2):318–332, June 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/4598992.

Appelhans:2019:CWA

[App19] Jeffery R. Appelhans. A Catholic Whig in the age of reason: Science, sociability, and the everyday life of William H. Keating. Proceedings of the American Philosophical Society, 163(1):??, March 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2019-11/attachments/Appelhans.pdf.

Al-Qadi:2003:PCH

[AQ03] Wadad Kadi Al-Qadi. The primordial covenant and human history in the Qur’ān. Proceedings of the American Philosophical Society, 147(4):332–338, December 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558296.

Armstrong:2010:ISE

[Arm10] Neil A. Armstrong. Introduction: Space: The evolving frontier. Proceedings of the American Philosophical Society, 154(4):399–401, December 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23056859.

Arrow:2011:GDJ

[Arr11] Kenneth Arrow. Gérard Debreu: 4 July 1921–31 December 2004. Proceedings of the American Philosophical Society, 155(3):319–325, September 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23208689.

Arrow:2014:CVD

[Arr14] Kenneth J. Arrow. Conflict of values: A decision view. Proceedings of the American Philosophical Society, 158(1):25–30, March 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Arrow.pdf;   https://www.
jstor.org/stable/24640142.

Ascheri:2011:DMS

[Asc11] Mario Ascheri. Domenico Maffei: 26 September 1925–4 July 2009. Proceedings of the American Philosophical Society, 155(3):351–355, September 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23208692.

Ashforth:2009:CTM

[Ash09] Adam Ashforth. Charles Tilly: 27 May 1929–29 April 2008. Proceedings of the American Philosophical Society, 153(3):371–380, September 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/40541680.

Atiyah:2006:BFE

[Ati06] Michael Atiyah. Benjamin Franklin and the Edinburgh Enlightenment. Proceedings of the American Philosophical Society, 150 (4):591–606, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599027.

Atkinson:2004:AFR

[Atk04] Richard C. Atkinson. Academic freedom and the research university. Proceedings of the American Philosophical Society, 148(2):195–204, June 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558284.

Atkinson:2008:DSS

[Atk08] Richard C. Atkinson. David Stephen Saxon, 8 February 1920–8 December 2005. Proceedings of the American Philosophical Society, 152(2):261–265, June 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478496.

Awodey:2007:SML

[Awo07] Steve Awodey. Saunders Mac Lane, 4 August 1909–14 April 2005. Proceedings of the American Philosophical Society, 151(3):351–356, September 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4599078.

Ayala:2012:WMF

[Aya12] Francisco J. Ayala. Walter Monroe Fitch: 21 May 1929–10 March 2011. Proceedings of the American Philosophical Society, 156(4): 435–442, December 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558236.

Badia:2017:MRM

[Bad17] Lola Badia. Martí De Riquer 3 May 1914–17 September 2013. Proceedings of the American Philosophical Society, 161(2):157–161, June 2017. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/De%20Riquer.pdf.

Bagnall:2002:ALD

[Bag02] Roger S. Bagnall. Alexandria: Library of dreams. Proceedings of the American Philosophical Society, 146(4):348–362, December 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/1558311.

Bailyn:2010:EFR

[Bai10] Bernard Bailyn. Eugene F. Rice Jr.: 20 August 1924–4 August 2008. Proceedings of the American Philosophical Society, 154(4):493–497, December 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23056868.

Bailyn:2013:OHS

[Bai13] Bernard Bailyn. Oscar Handlin: 29 September 1915–20 September 2011. Proceedings of the American Philosophical Society, 157(2):243– 247, June 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/24640243.

Baker:2010:CFB

[Bak10] Lee D. Baker. The cult of Franz Boas and his “conspiracy” to destroy the white race. Proceedings of the American Philosophical Society, 154(1):8–18, March 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/20721523.

Baldwin:2012:TNG

[Bal12] Ian T. Baldwin. Training a new generation of biologists: The genomeenabled field biologists. Proceedings of the American Philosophical Society, 156(2):205–214, June 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558192.

Bartoshuk:2000:CPM

[Bar00] Linda M. Bartoshuk. Carl Pfaffmann, 27 May 1913–16 April 1994. Proceedings of the American Philosophical Society, 144(3):323–327, September 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1515597.

Barker:2003:ISH

[Bar03a] Clyde F. Barker. Introduction to the Symposium in Honor of Jonathan E. Rhoads. Proceedings of the American Philosophical Society, 147(3):199–203, September 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/1558210.

Barker:2003:ETR

[Bar03b] Jane Barker. Elizabeth (Tibby) S. Russell, 1 May 1913–28 May 2001. Proceedings of the American Philosophical Society, 147(3):303–309, September 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558220.

Barker:2003:FBA

[Bar03c] Nicolas Barker. Frederick B. Adams Jr., 28 March 1910–7 January 2001. Proceedings of the American Philosophical Society, 147(2):153– 156, June 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558257.

Barker:2005:JRM

[Bar05] Clyde F. Barker. Jonathan E. Rhoads, 9 May 1907–3 January 2002. Proceedings of the American Philosophical Society, 149(3):431–440, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598948.

Barker:2009:JLT

[Bar09] Clyde F. Barker. The Jayne Lecture: Thomas Eakins and his medical clinics. Proceedings of the American Philosophical Society, 153(1):1– 47, March 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541627.

Barker:2010:JCT

[Bar10a] Clyde F. Barker. James C. Thompson, 16 August 1928–9 May 2008. Proceedings of the American Philosophical Society, 154(1):117–123, March 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/20721534.

Barondes:2010:GSM

[Bar10b] Samuel Barondes. Gunther S. Stent, 28 March 1924–12 June 2008. Proceedings of the American Philosophical Society, 154(1):113–116, March 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/20721533.

Barker:2016:VMO

[Bar16] Clyde F. Barker. Victor A. McKusick: 21 October 1921–22 July 2008. Proceedings of the American Philosophical Society, 160(1):103–107, March 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/VA_McKusick.pdf.

Barker:2018:IAS

[Bar18a] Clyde F. Barker. Introduction to APS session in celebration of Keith Stewart Thomson. Proceedings of the American Philosophical Society, 162(2):155–157, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.org/

sites/default/files/2018-11/attachments/Barker.pdf.

Barker:2018:TSM

[Bar18b] Clyde F. Barker. Thomas E. Starzl 11 March 1926–4 March 2017. Proceedings of the American Philosophical Society, 162(2):213–220, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2018-11/attachments/Starzl.pdf.

Barry:2019:PND

[Bar19] John M. Barry. The past is never dead. It’s not even past: The 1918 influenza pandemic and its lessons. Proceedings of the American Philosophical Society, 163(4):304–311, December 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2021-01/attachments/Barry_0.pdf.

Bass:2006:NTA

[Bas06] George F. Bass. New techniques of archaeology and Greek shipwrecks of the Sixth and Fifth Centuries BC. Proceedings of the American Philosophical Society, 150(1):1–14, March 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/4598971.

Bassler:2010:CCC

[Bas10] Bonnie Bassler. Cell to cell communication. Proceedings of the American Philosophical Society, 154(3):307–314, September 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/41000083.

Bass:2013:CHG

[Bas13] George F. Bass. Crawford H. Greenewalt Jr.: 3 June 1937–4 May 2012. Proceedings of the American Philosophical Society, 157 (3):355–360, September 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640400.

Baumol:2005:DPM

[Bau05] William J. Baumol. Daniel Patrick Moynihan, 16 March 1927– 26 March 2003. Proceedings of the American Philosophical Society, 149(2):267–272, June 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598932.

Babcock:2019:LLS

[BD19] Todd M. Babcock and Edwin F. S. Danson. Locating liberty: The 1769 State House Observatory. Proceedings of the American Philosophical Society, 163(4):??, December 2019. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2021-
01/attachments/Babcock%20and%20Danson.pdf.

Baker:2004:JRP

[BDN04] William O. Baker, Edward E. David, and A. Michael Noll. John R. Pierce, 27 March 1910–2 April 2002. Proceedings of the American Philosophical Society, 148(1):145–149, March 2004. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/1558250.

Bear:2004:WGG

[Bea04] Greg Bear. When genes go walkabout. Proceedings of the American Philosophical Society, 148(3):324–331, September 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1558229.

Bedini:2003:BHS

[Bed03] Silvio A. Bedini. Brooke Hindle, 28 September 1918–3 June 2001. Proceedings of the American Philosophical Society, 147(3):281–285, September 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558216.

Bederson:2014:DFH

[Bed14] Benjamin Bederson. Donald Frederick Hornig: 17 March 1920–21 January 2013. Proceedings of the American Philosophical Society, 158(4):449–454, December 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Hornig.pdf;   https://www.jstor.org/stable/24640190.

Begley:2007:SBJ

[Beg07] Louis Begley. Saul Bellow, 10 June 1915–5 April 2005. Proceedings of the American Philosophical Society, 151(4):435–439, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/25478456.

Bejan:2019:TCF

[Bej19] Teresa M. Bejan. Two concepts of freedom (of speech). Proceedings of the American Philosophical Society, 163(2):??, June 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2020-03/attachments/Bejan.pdf.

Bendiner:2002:JBO

[Ben02] Robert Bendiner. John B. Oakes, 23 April 1913–5 April 2001. Proceedings of the American Philosophical Society, 146(3):307–310, September 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558173.

Benhabib:2004:JNS

[Ben04] Seyla Benhabib. Judith Nisse Shklar, 24 September 1928–16 September 1992. Proceedings of the American Philosophical Society, 148 (4):529–534, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558147.

Bergson:2000:WLA

[Ber00] Abram Bergson. Wassily Leontief, 5 August 1906–5 February 1999. Proceedings of the American Philosophical Society, 144(4):465–468, December 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1515622.

Berg:2006:MBB

[Ber06] Howard C. Berg. Marvels of bacterial behavior. Proceedings of the American Philosophical Society, 150(3):428–442, September 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/4599006.

Berg:2009:AKM

[Ber09] Paul Berg. Arthur Kornberg, 3 March 1918–26 October 2007. Proceedings of the American Philosophical Society, 153(4):467–475, December 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/20721515.

Berg:2013:RJG

[Ber13] Paul Berg. Robert Joy Glaser: 11 September 1918–7 June 2012. Proceedings of the American Philosophical Society, 157(4):465–469, December 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/24640228.

Berenbaum:2014:BCC

[Ber14] May R. Berenbaum. Bees in crisis: Colony collapse, honey laundering, and other problems bee-setting American apiculture. Proceedings of the American Philosophical Society, 158(3):229–247, September 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Berenbaum.pdf;   https:
//www.jstor.org/stable/24640128.

Bevington:2008:WCB

[Bev08] David Bevington. Wayne C. Booth, 22 February 1921–10 October 2005. Proceedings of the American Philosophical Society, 152(2):217– 220, June 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/25478488.

Baker:2013:JGB

[BF13] Neal K. Baker and Stephen D. Fantone. James Gilbert Baker: 11 November 1914–29 June 2005. Proceedings of the American Philosophical Society, 157(2):237–241, June 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/24640242.

Bodmer:2011:JLM

[BG11] Walter Bodmer and Ann Ganesan. Joshua Lederberg: 23 May 1925– 2 February 2008. Proceedings of the American Philosophical Society, 155(4):477–499, December 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23208787.

Bryson:2008:GFC

[BGS^{+}08] Arthur E. Bryson, Harvey Greenspan, Howard A. Stone, Tai T. Wu, and Frederick H. Abernathy. George Francis Carrier, 4 May 1918–8 March 2002. Proceedings of the American Philosophical Society, 152(2):229–233, June 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478490.

Bialkin:2004:ISJ

[Bia04] Kenneth J. Bialkin. Isaac Stern, 21 July 1920–22 September 2001. Proceedings of the American Philosophical Society, 148(4):535–538, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558148.

Bailyn:2008:FFD

[BIK^{+}08] Bernard Bailyn, Akira Iriye, William C. Kirby, Ernest R. May, and Charles S. Maier. Franklin L. Ford 26 December 1920–31 August 2003. Proceedings of the American Philosophical Society, 152(4):541–545, December 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541607.

Billington:2001:HIR

[Bil01] James H. Billington. Humanizing the information revolution. Proceedings of the American Philosophical Society, 145(4):579–586, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558195.

Birgeneau:2015:CMV

[Bir15] Robert J. Birgeneau. Charles M. Vest: 9 September 1941–12 December 2013. Proceedings of the American Philosophical Society, 159(1):119–127, March 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Vest.pdf;   https://www.jstor.org/stable/24640176.

Bishop:2003:GBO

[Bis03] Tom Bishop. Germaine Brée, 2 October 1907–22 September 2001. Proceedings of the American Philosophical Society, 147(3):275–279, September 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558215.

Bisson:2005:RBM

[Bis05] Thomas N. Bisson. Robert Brentano, 19 May 1926–21 November 2002. Proceedings of the American Philosophical Society, 149(2):243– 248, June 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598928.

Bearn:2007:MMJ

[BK07] Alexander G. Bearn and Richard M. Krause. Maclyn McCarty, 9 June 1911–2 January 2005. Proceedings of the American Philosophical Society, 151(2):247–253, June 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599063.

Burnham:2007:ETC

[BL07] Scott Burnham and Paul Lansky. Edward Toner Cone, 4 May 1917– 23 October 2004. Proceedings of the American Philosophical Society, 151(1):103–106, March 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599048.

Blight:2016:RMJ

[BL16] James G. Blight and Janet M. Lang. Robert S. McNamara 9 June 1916-6 July 2009. Proceedings of the American Philosophical Society, 160(4):425–435, December 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.org/

sites/default/files/2017-07/attachments/McNamara.pdf.

Blaauw:2002:HCV

[Bla02] Adriaan Blaauw. Hendrik Christoffel van de Hulst: 19 November 1918–31 July 2000. Proceedings of the American Philosophical Society, 146(4):419–423, December 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558318.

Blau:2010:PMB

[Bla10] Reva Blau. Peter Michael Blau: 7 February 1918–12 March 2002. Proceedings of the American Philosophical Society, 154(3):317–320, September 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/41000084.

Blasi:2016:ALM

[Bla16] Vincent Blasi. Anthony Lewis 27 March 1927–25 March 2013. Proceedings of the American Philosophical Society, 160(2):207–211, June 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Lewis.pdf.

Blinder:2019:LTE

[Bli19] Alan S. Blinder. The lamppost theory of economic policy. Proceedings of the American Philosophical Society, 163(3):??, September 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2020-08/attachments/Blinder.pdf.

Bloch:2012:GUH

[Blo12] R. Howard Bloch. Good uses of the humanities in bad times. Proceedings of the American Philosophical Society, 156(4):416–422, December 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23558234.

Blume:2011:MCZ

[Blu11] Anna Blume. Maya concepts of zero. Proceedings of the American Philosophical Society, 155(1):51–88, March 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/23056849.

Bobbitt:2007:WRO

[Bob07] Philip Chase Bobbitt. Walt Rostow, 7 October 1916–13 February 2003. Proceedings of the American Philosophical Society, 151 (4):479–486, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478464.

Bock:2007:EMJ

[Boc07] Walter J. Bock. Ernst Mayr, 5 July 1904–3 February 2005. Proceedings of the American Philosophical Society, 151(3):357–370, September 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4599079.

Boer:2007:EGM

[Boe07] Charles Boer. Eugenio Garin, 9 May 1909–29 December 2004. Proceedings of the American Philosophical Society, 151(1):107–111, March 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4599049.

Bogle:2014:NLH

[Bog14] John C. Bogle. The (non) lessons of history — and the (real) lessons of return sources and investment costs. Proceedings of the American Philosophical Society, 158(3):209–221, September 2014. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Bogle.pdf;   https://www.jstor.org/stable/24640126.

Bois:2016:CGP

[Boi16] Yve-Alain Bois. Can a genuine Picasso be a fake? Proceedings of the American Philosophical Society, 160(1):6–17, March 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Bois.pdf.

Bok:2006:ACM

[Bok06] Derek Bok. Archibald Cox, 17 May 1912–29 May 2004. Proceedings of the American Philosophical Society, 150(3):461–465, September 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/4599010.

Bombieri:2010:ABM

[Bom10] Enrico Bombieri. Armand Borel: 21 May 1923–11 August 2003. Proceedings of the American Philosophical Society, 154(3):321–328, September 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/41000085.

Boon:2010:ABG

[Boo10] James A. Boon. On alternating Boasians: Generational connections. Proceedings of the American Philosophical Society, 154(1):19– 30, March 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/20721524.

Borel:2001:AWM

[Bor01] Armand Borel. André Weil, 6 May 1906–6 August 1998. Proceedings of the American Philosophical Society, 145(1):107–114, March 2001.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/1558329.

Borucki:2017:KBD

[Bor17] William J. Borucki. Kepler: A brief discussion of the mission and exoplanet results. Proceedings of the American Philosophical Society, 161 (1):38–65, March 2017. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-07/attachments/Borucki.pdf.

Bowering:2003:QVG

[Böw03] Gerhard H. Böwering. The Qur’ān as the voice of God. Proceedings of the American Philosophical Society, 147(4):347–353, December 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/1558298.

Bowersock:2009:BSG

[Bow09] G. W. Bowersock. Beverly Sills Greenough 25 May 1929–2 July 2007. Proceedings of the American Philosophical Society, 153(1):89– 95, March 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541633.

Bowen:2010:RFG

[Bow10] William G. Bowen. Robert Francis Goheen, 15 August 1919–31 March 2008. Proceedings of the American Philosophical Society, 154(1):95– 102, March 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/20721530.

Bower:2011:NMA

[Bow11] Gordon H. Bower. Neal E. Miller: 3 August 1909–23 March 2002. Proceedings of the American Philosophical Society, 155(3):357–365, September 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23208693.

Bowersock:2018:ACM

[Bow18] Glen W. Bowersock. Alan Cameron 13 March 1938–31 July 2017. Proceedings of the American Philosophical Society, 162(3):305–312, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2019-03/attachments/Cameron.pdf.

Boyd:2007:NDN

[Boy07] Donald W. Boyd. Norman Dennis Newell, 27 January 1909–18 April 2005. Proceedings of the American Philosophical Society, 151 (3):385–390, September 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599082.

Ben-Rafael:2012:SES

[BR12] Eliezer Ben-Rafael. Shmuel N. Eisenstadt: 10 September 1923–2 September 2010. Proceedings of the American Philosophical Society, 156(2):223–227, June 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558194.

Brady:2006:WJB

[Bra06] Thomas A. Brady, Jr. William J. Bouwsma, 22 November 1923– 2 March 2004. Proceedings of the American Philosophical Society, 150(3):457–460, September 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599009.

Bradley:2008:HMI

[Bra08] William G. Bradley. History of medical imaging. Proceedings of the American Philosophical Society, 152(3):349–361, September 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/40541591.

Branscomb:2010:HBA

[Bra10] Lewis M. Branscomb. Harvey Brooks: 5 August 1915–28 May 2004. Proceedings of the American Philosophical Society, 154(4):461–469, December 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23056864.

Breslow:2000:EN

[Bre00] Ronald Breslow. Extending nature. Proceedings of the American Philosophical Society, 144(3):267–274, September 2000. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/1515589.

Brest:2005:DSP

[Bre05] Paul Brest. In defense of strategic philanthropy. Proceedings of the American Philosophical Society, 149(2):132–140, June 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/4598922.

Brimmer:2000:WMM

[Bri00] Andrew F. Brimmer. William McChesney Martin, Jr., 17 December 1906–27 July 1998. Proceedings of the American Philosophical Society, 144(2):231–235, June 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515637.

Brieger:2004:OTO

[Bri04] Gert H. Brieger. Owsei Temkin, 6 October 1902–18 July 2002. Proceedings of the American Philosophical Society, 148(4):539–545, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558149.

Bricker:2006:EZV

[Bri06] Victoria R. Bricker. Evon Zartman Vogt Jr., 20 August 1918– 13 May 2004. Proceedings of the American Philosophical Society, 150(3):517–522, September 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599018.

Bricker:2007:LCA

[Bri07] Victoria R. Bricker. Literary continuities across the transformation from Maya hieroglyphic to alphabetic writing. Proceedings of the American Philosophical Society, 151(1):27–42, March 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/4599042.

Bright:2008:MBE

[Bri08] William Bright. Murray Barnson Emeneau, 28 February 1904– 29 August 2005. Proceedings of the American Philosophical Society, 152(1):141–143, March 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478476.

Brillinger:2012:DHB

[Bri12] David R. Brillinger. David H. Blackwell: 24 April 1919–8 July 2010. Proceedings of the American Philosophical Society, 156(2):217–222, June 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23558193.

Browne:2001:DCS

[Bro01] Janet Browne. Darwin in caricature: A study in the popularisation and dissemination of evolution. Proceedings of the American Philosophical Society, 145(4):496–509, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1558189.

Broecker:2002:DCR

[Bro02a] W. S. Broecker. Dust: Climate’s Rosetta Stone. Proceedings of the American Philosophical Society, 146(1):77–80, March 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1558157.

Brown:2002:JRM

[Bro02b] Jonathan Brown. John Rupert Martin, 27 September 1916–26 July 2000. Proceedings of the American Philosophical Society, 146 (3):303–306, September 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558172.

Brombert:2005:GMO

[Bro05] Victor Brombert. Georges May, 7 October 1920–28 February 2003. Proceedings of the American Philosophical Society, 149(2):261–265, June 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4598931.

Brombert:2009:RFS

[Bro09] Victor Brombert. Robert Fagles: 11 September 1933–26 March 2008. Proceedings of the American Philosophical Society, 153(3):323–326, September 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541674.

Brown:2010:HCJ

[Bro10] Peter Brown. Henry Chadwick, 23 June 1920–17 June 2008. Proceedings of the American Philosophical Society, 154(1):79–86, March 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/20721528.

Brookfield:2017:BCC

[Bro17] John Brookfield. Bryan Campbell Clarke 24 June 1932–27 February 2014. Proceedings of the American Philosophical Society, 161(1): 85–95, March 2017. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-07/attachments/Clarke.pdf.

Bruchac:2018:BCC

[Bru18] Margaret M. Bruchac. Broken chains of custody: Possessing, dispossessing, and repossessing lost wampum belts. Proceedings of the American Philosophical Society, 162(1):59–108, March 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2018-08/attachments/Bruchac.pdf.

Billington:2006:SIB

[BSW06] James H. Billington, Arthur Schlesinger, Jr., and Robert Wokler. Sir Isaiah Berlin, 6 June 1909–5 November 1997. Proceedings of the American Philosophical Society, 150(4):663–672, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/4599033.

Bucher:2013:EDK

[Buc13] Doris J. Bucher. Edwin D. Kilbourne: 10 July 1920–21 February 2011. Proceedings of the American Philosophical Society, 157(1):121–125, March 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23558145.

Burkhardt:2001:DCM

[Bur01] Frederick H. Burkhardt. Darwin and the Copley Medal. Proceedings of the American Philosophical Society, 145(4):510–518, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/1558190.

Burbidge:2003:SFH

[Bur03] E. Margaret Burbidge. Sir Fred Hoyle, 24 June 1915–20 August 2001. Proceedings of the American Philosophical Society, 147(4):405–412, December 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558304.

Burnett:2008:DEP

[Bur08] Charles Burnett. David Edwin Pingree, 2 January 1933–11 November 2005. Proceedings of the American Philosophical Society, 152(2):255– 259, June 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/25478495.

Becklin:2016:GNS

[BWW16] Eric Becklin and Gareth Wynn-Williams. Gerry Neugebauer 3 September 1932–26 September 2014. Proceedings of the American Philosophical Society, 160(4):443–447, December 2016. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2017-
07/attachments/Neugebauer.pdf.

Cairns:2005:GHA

[Cai05] John Cairns, Jr. Garrett Hardin, 21 April 1915–14 September 2003. Proceedings of the American Philosophical Society, 149(3):413–419, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598945.

Canavan:2001:SBM

[Can01] Gregory H. Canavan. Space-based missile defense and stability. Proceedings of the American Philosophical Society, 145(3):270–282, September 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558109.

Carter:2006:CSD

[Car06] Joseph Coleman Carter. The contributions of the sciences to the discovery of ancient rural civilization. Proceedings of the American Philosophical Society, 150(1):15–61, March 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/4598972.

Casper:2004:GGM

[Cas04] Gerhard Casper. Gerald Gunther, 26 May 1927–30 July 2002. Proceedings of the American Philosophical Society, 148(4):493–497, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558143.

Casper:2011:YMU

[Cas11] Gerhard Casper. A young man from “Ultima Thule” visits Jefferson: Alexander von Humboldt in Philadelphia and Washington. Proceedings of the American Philosophical Society, 155(3):247–262, September 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23208683.

Casper:2014:RWL

[Cas14] Gerhard Casper. Richard W. Lyman: 18 October 1923–27 May 2012. Proceedings of the American Philosophical Society, 158 (4):435–439, December 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Lyman.pdf;   https://www.jstor.org/stable/24640188.

Cropper:2012:AHS

[CD12] Elizabeth Cropper and Charles Dempsey. Anne D’Harnoncourt: 7 September 1943–1 June 2008. Proceedings of the American Philosophical Society, 156(3):333–338, September 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/23558094.

Cerami:2014:FYO

[Cer14] Anthony Cerami. A forty-year odyssey in the sea of translational medicine. Proceedings of the American Philosophical Society, 158 (4):400–410, December 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Cerami_0.pdf;   https://www.jstor.org/stable/24640184.

Chafe:2000:FGL

[Cha00] Wallace Chafe. Floyd Glenn Lounsbury, 25 April 1914–14 May 1998. Proceedings of the American Philosophical Society, 144(2):225–229, June 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1515636.

Chandler:2001:TCC

[Cha01] Alfred D. Chandler. Thomas Childs Cochran, 29 April 1902–2 May 1999. Proceedings of the American Philosophical Society, 145(2):193– 196, June 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558272.

Challinor:2003:DRS

[Cha03] David Challinor. S. Dillon Ripley, 20 September 1913–12 March 2001. Proceedings of the American Philosophical Society, 147(3):297–302, September 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558219.

Clair:2013:TTE

[CHH13] Matthew Clair, Peter Blair Henry, and Sandile Hlatshwayo. Two tales of entrepreneurship: Barbados, Jamaica, and the 1973 oil price shock. Proceedings of the American Philosophical Society, 157(1):32– 57, March 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23558139.

Cho:2000:RLS

[Cho00] Alfred Y. Cho. Robert A. Laudise, 2 September 1930–20 August 1998. Proceedings of the American Philosophical Society, 144(3):309–312, September 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1515594.

Choppin:2012:AGB

[Cho12] Purnell W. Choppin. Alexander Gordon Bearn: 29 March 1923– 15 May 2009. Proceedings of the American Philosophical Society, 156(1):67–73, March 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558079.

Churchland:2011:BS

[Chu11] Patricia Smith Churchland. The brain and its self. Proceedings of the American Philosophical Society, 155(1):41–50, March 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/23056848.

Clark:2000:BBR

[Cla00] George W. Clark. Bruno Benedetto Rossi, 13 April 1905–21 November 1993. Proceedings of the American Philosophical Society, 144 (3):329–341, September 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515598.

Clardy:2007:DNC

[Cla07] Jon Clardy. Discovery of new compounds in nature. Proceedings of the American Philosophical Society, 151(2):201–210, June 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/4599058.

Clingman:2015:NGN

[Cli15] Stephen Clingman. Nadine Gordimer: 20 November 1923–13 July 2014. Proceedings of the American Philosophical Society, 159(4):467– 473, December 2015. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Gordimer.pdf;
 https://www.jstor.org/stable/26159199.

Coffin:2004:ERF

[Cof04] John M. Coffin. Evolution of retroviruses: Fossils in our DNA. Proceedings of the American Philosophical Society, 148(3):264–280, September 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558226.

Cohen:2003:CTJ

[Coh03] Seymour S. Cohen. The correspondence of Thomas Jefferson and Thomas Cooper: A previously unpublished manuscript of Dumas Malone. Proceedings of the American Philosophical Society, 147 (1):39–64, March 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558127.

Cohen:2004:ECA

[Coh04] Seymour S. Cohen. Erwin Chargaff, 11 August 1905–20 June 2002. Proceedings of the American Philosophical Society, 148(2):221–228, June 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558287.

Cohen:2009:DRM

[Coh09] Joel E. Cohen. The demography of the resident membership of the American Philosophical Society. Proceedings of the American Philosophical Society, 153(2):200–213, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/40541659.

Cohen:2010:PCC

[Coh10] Joel E. Cohen. Population and climate change. Proceedings of the American Philosophical Society, 154(2):158–182, June 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/41000096.

Cohen:2011:MAO

[Coh11] Joel E. Cohen. Mixing apples and oranges: What poetry and applied mathematics have in common. Proceedings of the American Philosophical Society, 155(2):189–202, June 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/23056834.

Coleman:2004:TUW

[Col04] William Thaddeus Coleman, Jr. Truths that unfortunately were not, and still are not, sufficiently self-evident. Proceedings of the American Philosophical Society, 148(4):434–454, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1558139.

Connick:2000:KPJ

[Con00a] Robert E. Connick. Kenneth Pitzer, 6 January 1914–26 December 1997. Proceedings of the American Philosophical Society, 144 (4):479–483, December 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515624.

Constable:2000:CTD

[Con00b] Giles Constable. Charles Till Davis, 14 April 1929–10 April 1998. Proceedings of the American Philosophical Society, 144(3):297–300, September 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1515592.

Constable:2003:SSR

[Con03] Giles Constable. Sir Steven Runciman, 7 July 1903–1 November 2000. Proceedings of the American Philosophical Society, 147(1):95– 101, March 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558132.

Connor:2010:JHA

[Con10] W. Robert Connor. John H. D’Arms: 27 November 1934–22 January 2002. Proceedings of the American Philosophical Society, 154 (3):329–336, September 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/41000086.

Cook:2003:VHI

[Coo03] Michael Cook. The voice of honest indignation is the voice of god. Proceedings of the American Philosophical Society, 147(4):325–331, December 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558295.

Cooper:2006:ASL

[Coo06] John Milton Cooper, Jr. Arthur Stanley Link, 8 August 1920– 26 March 1998. Proceedings of the American Philosophical Society, 150(4):679–683, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599035.

Cornel:2017:SOS

[Cor17] Tabea Cornel. Something old, something new, something pseudo, something true: Pejorative and deferential references to phrenology since 1840. Proceedings of the American Philosophical Society, 161 (4):299–332, December 2017. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.org/

sites/default/files/2018-03/attachments/Cornel.pdf.

Costa:2002:SMW

[Cos02] James T. Costa. Scale models? What insect societies teach us about ourselves. Proceedings of the American Philosophical Society, 146(2):170–180, June 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558200.

Cotton:2000:DHR

[Cot00] F. Albert Cotton. Derek H. R. Barton, 8 September 1918–16 March 1998. Proceedings of the American Philosophical Society, 144(3):291–296, September 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515591.

Cox:2019:DMB

[Cox19] John D. Cox. David Martin Bevington. Proceedings of the American Philosophical Society, 163(3):??, September 2019. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2020-
08/attachments/Bevington.pdf.

Clegg:2010:DKJ

[CR10] Dennis O. Clegg and Teisha Rowland. Daniel E. Koshland Jr.: 30 March 1920–23 July 2007. Proceedings of the American Philosophical Society, 154(4):477–483, December 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/23056866.

Craig:2001:GMS

[Cra01] Gordon A. Craig. George L. Mosse, 20 September 1918–22 January 1999. Proceedings of the American Philosophical Society, 145(2):221– 225, June 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558277.

Crighton:2000:SJL

[Cri00] David G. Crighton. Sir James Lighthill, 23 January 1924–17 July 1998. Proceedings of the American Philosophical Society, 144(3): 313–317, September 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515595.

Crow:2002:JVN

[Cro02] James F. Crow. James V. Neel, 22 March 1915–1 February 2000. Proceedings of the American Philosophical Society, 146(1):123–127, March 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558161.

Cropper:2009:CHS

[Cro09] Elizabeth Cropper. Craig Hugh Smyth, 28 July 1915–22 December 2006. Proceedings of the American Philosophical Society, 153 (4):495–500, December 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/20721518.

Croce:2018:AGK

[Cro18] Carlo Croce. Alfred G. Knudson 9 August 1922–10 July 2016. Proceedings of the American Philosophical Society, 162(4):395–400, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2019-06/attachments/Knudson.pdf.

Culler:2017:MHA

[CSG17] Jonathan Culler, Paul Sawyer, and Roger Gilbert. Meyer Howard Abrams: 23 July 1912–21 April 2015. Proceedings of the American Philosophical Society, 161(2):183–188, June 2017. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Abrams.pdf.

Chu:2004:ASM

[CT04] Steven Chu and Charles H. Townes. Arthur Schawlow, 5 May 1921–28 April 1999. Proceedings of the American Philosophical Society, 148(2):243–254, June 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558290.

Culler:2014:EKS

[Cul14] Jonathan Culler. Eve Kosofsky Sedgwick: 2 May 1950–12 April 2009. Proceedings of the American Philosophical Society, 158(3):281– 286, September 2014. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Sedgwick.pdf;
 https://www.jstor.org/stable/24640130.

Culler:2017:BRO

[Cul17] Jonathan Dwight Culler. Benedict Richard O’Gorman Anderson 26 August 1936–13 December 2015. Proceedings of the American Philosophical Society, 161(1):107–110, March 2017. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2017-
07/attachments/Anderson.pdf.

Cunningham:2016:WLC

[Cun16] Jeff Cunningham. Walter Leland Cronkite, Jr. 4 November 1916–17 July 2009. Proceedings of the American Philosophical Society, 160 (1):75–95, March 2016. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Cronkite.pdf.

deAlmeida:2003:CCF

[dA03] Darcy Fontoura de Almeida. Carlos Chagas Filho, 12 September 1910–16 February 2000. Proceedings of the American Philosophical Society, 147(1):77–82, March 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558130.

Dalgarno:2005:SRW

[Dal05] Alexander Dalgarno. Sir Robert Wilson, 16 April 1927–2 September 2002. Proceedings of the American Philosophical Society, 149(1):119– 122, March 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598918.

Danna:2010:DNO

[Dan10] Kathleen J. Danna. Daniel Nathans: 30 October 1928–16 November 1999. Proceedings of the American Philosophical Society, 154 (3):337–354, September 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/41000087.

Darnton:2001:LSD

[Dar01] Robert Darnton. Lawrence Stone, 4 December 1919–16 June 1999. Proceedings of the American Philosophical Society, 145(3):379–383, September 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558119.

Darnell:2015:FI

[Dar15] Regna Darnell. A framing introduction. Proceedings of the American Philosophical Society, 159(3):247–250, September 2015. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Darnell.pdf;   https://www.jstor.org/stable/
26159135.

Darnell:2018:FBL

[Dar18] Regna Darnell. Franz Boas’s legacy of “useful knowledge”: The APS Archives and the future of Americanist anthropology. Proceedings of the American Philosophical Society, 162(1):1–14, March 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2018-08/attachments/Darnell.pdf.

Darnell:2019:CFC

[Dar19] Regna Darnell. Conversation framed — The Center for Native American and Indigenous Research: a Jeffersonian legacy. Proceedings of the American Philosophical Society, 163(3):??, September 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2020-08/attachments/Darnell.pdf.

Davies:2012:HMH

[Dav12] Anna Morpurgo Davies. Henry Max Hoenigswald: 17 April 1915– 16 June 2003. Proceedings of the American Philosophical Society, 156(1):81–88, March 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558081.

Dunn:2006:ECC

[DDV06] Mary Maples Dunn, Richard S. Dunn, and John C. Van Horne. Edward C. Carter II, 10 January 1928–1 October 2002. Proceedings of the American Philosophical Society, 150(1):173–179, March 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/4598980.

DeMontebello:2003:RFP

[De 03] Philippe De Montebello. The Richard A. F. Penrose Lecture: The changing landscape of museums. Proceedings of the American Philosophical Society, 147(3):259–272, September 2003. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/1558214.

DeSoto:2013:TIB

[De 13] Hernando De Soto. Thoughts on the importance of boundaries. Proceedings of the American Philosophical Society, 157(1):22–31, March 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/23558138.

deCallatay:2017:GRJ

[de 17] François de Callataÿ. Georges Le Rider 27 January 1928–3 July 2014. Proceedings of the American Philosophical Society, 161(2):163–167, June 2017. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Le%20Rider.pdf.

Degler:2002:GWA

[Deg02] Carl N. Degler. Gordon Wright, 24 April 1912–11 January 2000. Proceedings of the American Philosophical Society, 146(3):317–322, September 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558175.

Delbanco:2012:WHO

[Del12a] Andrew H. Delbanco. What’s happening in our colleges: Thoughts on the new meritocracy. Proceedings of the American Philosophical Society, 156(3):303–308, September 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/23558091.

DeLuca:2012:HLA

[DeL12b] Hector F. DeLuca. Henry A. Lardy: 19 August 1917–4 August 2010. Proceedings of the American Philosophical Society, 156(3):353–356, September 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23558097.

Desai:2005:ABA

[Des05] Padma Desai. Abram Bergson, 21 April 1914–23 April 2003. Proceedings of the American Philosophical Society, 149(3):389–394, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4598942.

DesMarais:2010:EME

[Des10] David J. Des Marais. Exploring Mars for evidence of habitable environments and life. Proceedings of the American Philosophical Society, 154(4):402–421, December 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056860.

Davidson:2004:GVJ

[DF04] Donald Davidson and John Ferrari. Gregory Vlastos, 27 July 1907– 12 October 1991. Proceedings of the American Philosophical Society, 148(2):255–259, June 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558291.

Davis:2007:WAN

[DH07] Russ Davis and William Happer. William Aaron Nierenberg, 13 February 1919–10 September 2000. Proceedings of the American Philosophical Society, 151(2):261–266, June 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/4599065.

Dickey:2012:MLL

[DH12] Eleanor Dickey and Richard Hamilton. Mabel Louise Lang: 12 November 1917–21 July 2010. Proceedings of the American Philosophical Society, 156(2):245–250, June 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/23558197.

Dirks:2010:FBA

[Dir10] Nicholas B. Dirks. Franz Boas and the American University: A personal account. Proceedings of the American Philosophical Society, 154(1):31–39, March 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/20721525.

Doolittle:2004:MDK

[Doo04] Russell F. Doolittle. Martin David Kamen, 27 August 1913–31 August 2002. Proceedings of the American Philosophical Society, 148(4):499–503, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558144.

Doty:2005:JEN

[Dot05] Paul Doty. John T. Edsall, 3 November 1902–12 June 2002. Proceedings of the American Philosophical Society, 149(1):89–92, March 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/4598912.

Dowling:2002:FNG

[Dow02a] John E. Dowling. Fishing for novel genes. Proceedings of the American Philosophical Society, 146(4):337–347, December 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1558310.

Dowling:2002:GWN

[Dow02b] John E. Dowling. George Wald: 18 November 1906–12 April 1997. Proceedings of the American Philosophical Society, 146(4):431–439, December 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558320.

Doyle:2019:MIM

[Doy19] Michael W. Doyle. The model international mobility convention: Beyond migrants and refugees. Proceedings of the American Philosophical Society, 163(3):??, September 2019. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2020-
08/attachments/Doyle.pdf.

Davis:2007:GPM

[DP07] Philip J. Davis and Neil Patterson. Gerard Piel, 1 March 1915–5 September 2004. Proceedings of the American Philosophical Society, 151(2):271–277, June 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599067.

Drell:2007:HBJ

[Dre07] Sidney D. Drell. Hans Bethe, 2 July 1906–6 March 2005. Proceedings of the American Philosophical Society, 151(3):337–344, September 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4599076.

Drickamer:2002:HSG

[Dri02] Harry G. Drickamer. Herbert Sanders Gutowsky, 8 November 1919– 13 January 2000. Proceedings of the American Philosophical Society, 146(2):205–210, June 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558203.

Drickamer:2003:VGJ

[DT03] Harry G. Drickamer and John Meurig Thomas. Vitalii I. Gol’danskii, 18 June 1923–14 January 2001. Proceedings of the American Philosophical Society, 147(2):157–159, June 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1558258.

Dubernard:2011:HFA

[Dub11] Jean-Michel Dubernard. Hand and face allografts: Myth, dream, and reality. Proceedings of the American Philosophical Society, 155(1):13– 22, March 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23056845.

Duderstadt:2001:FUD

[Dud01] James J. Duderstadt. The future of the university in the digital age. Proceedings of the American Philosophical Society, 145(1):54–72, March 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558325.

Dunn:2007:DCB

[Dun07] Richard S. Dunn. The demographic contrast between slave life in Jamaica and Virginia, 1760–1865. Proceedings of the American Philosophical Society, 151(1):43–60, March 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/4599043.

Dunn:2008:IGE

[Dun08] Richard S. Dunn. Introduction to the GAFOH essays. Proceedings of the American Philosophical Society, 152(4):425–426, December 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/40541601.

Dunn:2011:RMJ

[Dun11] Richard S. Dunn. Robert L. Mcneil Jr.: 13 July 1915–20 May 2010. Proceedings of the American Philosophical Society, 155(2):225–230, June 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23056838.

Dutton:2016:BCJ

[Dut16] P. Leslie Dutton. Britton Chance 24 July 1913–16 November 2010. Proceedings of the American Philosophical Society, 160(3):309–312, September 2016. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-07/attachments/Chance.pdf.

Dyson:2001:WFT

[Dys01] Tim Dyson. World food trends: A neo-Malthusian prospect? Proceedings of the American Philosophical Society, 145(4):438–455, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558183.

Dyson:2006:MRF

[Dys06] Freeman Dyson. Marshall N. Rosenbluth, 5 February 1927–28 September 2003. Proceedings of the American Philosophical Society, 150(2):403–407, June 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599001.

Dyson:2010:JAW

[Dys10] Freeman Dyson. John Archibald Wheeler, 9 July 1911–13 April 2008. Proceedings of the American Philosophical Society, 154(1):125–129, March 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/20721535.

Eagleson:2000:RWC

[Eag00a] Peter S. Eagleson. The role of water in climate. Proceedings of the American Philosophical Society, 144(1):33–38, March 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1515603.

Eagleson:2000:GRC

[Eag00b] William B. Eagleson, Jr. George Roberts Clark, 12 January 1910– 15 June 1998. Proceedings of the American Philosophical Society, 144(2):201–205, June 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515632.

Eagleson:2008:DPE

[Eag08] William B. Eagleson, Jr. David P. Eastburn, 9 January 1921– 11 October 2005. Proceedings of the American Philosophical Society, 152(1):135–140, March 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478475.

Eastburn:2000:VD

[Eas00] David Eastburn. Vive le dilettante! Proceedings of the American Philosophical Society, 144(2):192–197, June 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1515631.

Easterling:2018:BMW

[Eas18] Pat Easterling. Bernard M. W. Knox 24 November 1914. 22 July 2010. Proceedings of the American Philosophical Society, 162(1): 109–114, March 2018. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2018-08/attachments/Knox.pdf.

Eagleson:2002:RSR

[EC02] William B. Eagleson and William T. Coleman. R. Stewart Rauch, Jr.: 15 July 1914–16 November 2001. Proceedings of the American Philosophical Society, 146(4):403–410, December 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1558315.

Ernst:2012:LKJ

[EDJG12] W. G. Ernst, Paul M. Davis, David D. Jackson, and Freeman Gilbert. Leon Knopoff: 1 July 1925–20 January 2011. Proceedings of the American Philosophical Society, 156(3):347–352, September 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/23558096.

Ekers:2011:IDA

[EK11] R. D. Ekers and K. I. Kellermann. Introduction: Discoveries in astronomy. Proceedings of the American Philosophical Society, 155(2):129– 133, June 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23056829.

Easton:2010:IPT

[EN10] Roger L. Easton, Jr. and William Noel. Infinite possibilities: Ten years of study of the Archimedes Palimpsest. Proceedings of the American Philosophical Society, 154(1):50–76, March 2010. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/20721527.

Engell:2001:WJB

[Eng01] James Engell. W. Jackson Bate, 23 May 1918–26 July 1999. Proceedings of the American Philosophical Society, 145(3):337–343, September 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558111.

Erk:2009:HBG

[Erk09] Frank C. Erk. H. Bentley Glass: 17 January 1906–16 January 2005. Proceedings of the American Philosophical Society, 153(3):327–339, September 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541675.

Ernst:2000:CCM

[Ern00] W. G. Ernst. Continental collision, mountain building, and diamonds in the rough. Proceedings of the American Philosophical Society, 144(1):1–32, March 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515602.

Ernst:2005:KBK

[Ern05] W. G. Ernst. Konrad B. Krauskopf, 30 November 1910–4 May 2003. Proceedings of the American Philosophical Society, 149(3):421–425, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598946.

Evans:2010:ICC

[Eva10] James Warren Evans. The immediate challenge of climate change: Reducing impacts on the poorest communities in the poorest countries. Proceedings of the American Philosophical Society, 154(2):183–191, June 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/41000097.

Evenson:2005:BMG

[Eve05] Robert E. Evenson. Besting Malthus: The Green Revolution. Proceedings of the American Philosophical Society, 149(4):469–486, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4598956.

Fackler:2009:FAC

[Fac09] John P. Fackler. Frank Albert Cotton 7 April 1930–20 February 2007. Proceedings of the American Philosophical Society, 153(1):83– 88, March 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541632.

Fass:2015:HAR

[Fas15] Paula S. Fass. How Americans raise their children: Generational relations from the Revolution to the global world. Proceedings of the American Philosophical Society, 159(1):85–94, March 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Fass.pdf;   https://www.
jstor.org/stable/24640172.

Fauci:2007:MRH

[Fau07] Anthony S. Fauci. Maurice R. Hilleman, 30 August 1919–11 April 2005. Proceedings of the American Philosophical Society, 151(4): 451–455, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478459.

Fauci:2014:CEK

[Fau14] Anthony S. Fauci. C. Everett Koop: 14 October 1916–25 February 2013. Proceedings of the American Philosophical Society, 158(4):455– 460, December 2014. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Koop_2.pdf;
https://www.jstor.org/stable/24640191.

Falthammar:2006:HAM

[FD06] Carl-Gunne Fälthammar and Alexander J. Dessler. Hannes Alfvén, 30 May 1908–2 April 1995. Proceedings of the American Philosophical Society, 150(4):649–662, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/4599032.

Frelinger:2017:RDO

[FDBH17] Jeffrey Frelinger, Tommy C. Douglas, Elizabeth P. Blankenhorn, and Lee Hood. Ray David Owen 30 October 1915–21 September 2014.

Proceedings of the American Philosophical Society, 161(2):169–176, June 2017. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Owen.pdf.

Frater:2011:JPW

[FEE11] Bob Frater, Philip Edwards, and Ron Ekers. John Paul Wild: 17 May 1923–10 May 2008. Proceedings of the American Philosophical Society, 155(3):377–381, September 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/23208695.

Feldman:2004:VRS

[Fel04] Noah Feldman. The voidness of repugnant statutes: Another look at the meaning of Marbury. Proceedings of the American Philosophical Society, 148(1):27–37, March 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558242.

Finger:2008:BFR

[FH08] Stanley Finger and Ian S. Hagemann. Benjamin Franklin’s risk factors for gout and stones: From genes and diet to possible lead poisoning. Proceedings of the American Philosophical Society, 152(2):189–206, June 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/25478486.

Finkleman:2001:CSS

[Fin01] David Finkleman. The contribution of space systems and strategic defense to nuclear stability. Proceedings of the American Philosophical Society, 145(3):260–269, September 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/1558108.

Fine:2005:EMR

[Fin05] Ruth Fine. Expanding the mainstream: Romare Bearden revisited. Proceedings of the American Philosophical Society, 149(1):40–55, March 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4598907.

Finger:2010:RCS

[Fin10] Harold B. Finger. Robert C. Seamans Jr.: 30 October 1918– 28 June 2008. Proceedings of the American Philosophical Society, 154(4):503–506, December 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056870.

Finch:2012:EHL

[Fin12] Caleb E. Finch. Evolution of the human lifespan, past, present, and future: Phases in the evolution of human life expectancy in relation to the inflammatory load. Proceedings of the American Philosophical Society, 156(1):9–44, March 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558076.

Finger:2015:BFW

[Fin15] Stanley Finger. Benjamin Franklin’s wife’s apoplexy and Mid-Eighteenth Century medicine. Proceedings of the American Philosophical Society, 159(2):169–197, June 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Finger.pdf;   https://www.jstor.org/stable/24640214.

Fischer:2008:CPK

[Fis08] Stanley Fischer. Charles P. Kindleberger, 12 October 1910–7 July 2003. Proceedings of the American Philosophical Society, 152(1):145– 149, March 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/25478477.

Fisher:2009:FHW

[Fis09] Harvey F. Fisher. Frank H. Westheimer 15 January 1912–14 April 2007. Proceedings of the American Philosophical Society, 153(2):243– 246, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541666.

Fisher:2012:CKM

[Fis12] Franklin M. Fisher. Carl Kaysen: 5 March 1920–8 February 2010. Proceedings of the American Philosophical Society, 156(3):339–346, September 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23558095.

Fiske:2013:DSU

[Fis13] Susan T. Fiske. Divided by status: Upward envy and downward scorn. Proceedings of the American Philosophical Society, 157(3):261–268, September 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/24640395.

Fishkin:2019:DWP

[Fis19] James Fishkin. Democracy when the people are thinking: Deliberation and democratic renewal. Proceedings of the American Philosophical Society, 163(2):??, June 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.

org/sites/default/files/2020-03/attachments/Fishkin.pdf.

Fitch:2002:SBT

[Fit02] Val L. Fitch. Sam Bard Treiman, 27 May 1925–30 November 1999. Proceedings of the American Philosophical Society, 146(1):133–138, March 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558163.

Fumaroli:2013:JWR

[FL13] Marc Fumaroli and Catherine Lafarge. Jacqueline Worms De Romilly: 26 March 1913–18 December 2010. Proceedings of the American Philosophical Society, 157(1):109–114, March 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/23558143.

Flathman:2005:BLV

[Fla05] Richard E. Flathman. The bases, limits, and values of pluralism: An engagement with William James. Proceedings of the American Philosophical Society, 149(2):159–198, June 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/4598924.

Fleming:2015:SCA

[Fle15] John V. Fleming. Sanative contagion among Cambridge Platonists. Proceedings of the American Philosophical Society, 159(4):409– 414, December 2015. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Fleming.pdf;
https://www.jstor.org/stable/26159193.

Fleming:2017:AWL

[Fle17] John V. Fleming. Arthur Walton Litz 31 October 1929–4 June 2014. Proceedings of the American Philosophical Society, 161(4):349– 352, December 2017. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2018-03/attachments/Litz.pdf.

Follesdal:2012:WVO

[Føl12] Dagfinn Føllesdal. Willard Van Orman Quine: 25 June 1908–25 December 2000. Proceedings of the American Philosophical Society, 156(1):99–104, March 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558084.

Fong:2007:FWM

[Fon07] Wen C. Fong. Frederick Wade Mote, 2 June 1922–10 February 2005. Proceedings of the American Philosophical Society, 151(2):255–260, June 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4599064.

Forney:2003:ISJ

[For03] Robert C. Forney. Irving S. Shapiro, 15 July 1916–13 September 2001. Proceedings of the American Philosophical Society, 147(4):413–416, December 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558305.

Fox:2009:NHM

[Fox09] Marye Anne Fox. Norman Hackerman: 2 March 1912–16 June 2007. Proceedings of the American Philosophical Society, 153(3):351–357, September 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541677.

Fee:2006:DBM

[FP06] Elizabeth Fee and Manon Parry. David Bodian, 15 May 1910–18 September 1992. Proceedings of the American Philosophical Society, 150(1):167–172, March 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598979.

Franklin:2004:SDO

[Fra04] Simon Franklin. Sir Dimitri Obolensky, 1 April 1918–23 December 2001. Proceedings of the American Philosophical Society, 148(1):139– 144, March 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558249.

Fre:2016:TRJ

[Frè16] Pietro Frè. Tullio Regge 11 July 1931–23 October 2014. Proceedings of the American Philosophical Society, 160(3):275–282, September 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2017-07/attachments/Regge.pdf.

Frantz-Szabo:2008:DOE

[FS08] Gabriella Frantz-Szabó. Dietz Otto Edzard 28 August 1930–2 June 2004. Proceedings of the American Philosophical Society, 152 (3):391–399, September 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541594.

Fuchs:2001:PPD

[Fuc01] Ira H. Fuchs. Prospects and possibilities of the digital age. Proceedings of the American Philosophical Society, 145(1):45–53, March 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/1558324.

Fulcher:2006:FIF

[Ful06] Jane F. Fulcher. French identity in flux: Vichy’s collaboration, and Antigone’s operatic triumph. Proceedings of the American Philosophical Society, 150(2):261–295, June 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/4598990.

Furstenberg:2015:WMD

[Fur15] Frank F. Furstenberg. Will marriage disappear? Proceedings of the American Philosophical Society, 159(3):241–246, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL http://live-amphil.pantheonsite.io/
sites/default/files/2017-10/attachments/Furstenberg.pdf;
 https://www.jstor.org/stable/26159134.

Futter:2012:TWG

[Fut12] Ellen V. Futter. Tribute to William T. Golden: A man for all seasons, a museum for all people and all time. Proceedings of the American Philosophical Society, 156(1):1–8, March 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/23558075.

Gomory:2006:HHG

[GA06] Ralph E. Gomory and Arlin M. Adams. Herman Heine Goldstine, 13 September 1913–16 June 2004. Proceedings of the American Philosophical Society, 150(2):367–377, June 2006. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/4598998.

Gallager:2003:CES

[Gal03] Robert G. Gallager. Claude Elwood Shannon, 30 April 1916–24 February 2001. Proceedings of the American Philosophical Society, 147(2):187–191, June 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558263.

Garwin:2001:SWS

[Gar01] Richard L. Garwin. Space weapons or space arms control? Proceedings of the American Philosophical Society, 145(3):243–259, September 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558107.

Gardner:2003:WRH

[Gar03] David Pierpont Gardner. William Redington Hewlett, 20 May 1913– 12 January 2001. Proceedings of the American Philosophical Society, 147(2):161–166, June 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558259.

Gardner:2008:JFS

[Gar08a] Howard Gardner. James O. Freedman 21 September 1935–21 March 2006. Proceedings of the American Philosophical Society, 152(3): 401–408, September 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541595.

Garwin:2008:LNW

[Gar08b] Richard L. Garwin. Living with nuclear weapons: Sixty years and counting. Proceedings of the American Philosophical Society, 152(1):69–82, March 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478469.

Gardner:2012:TRS

[Gar12] Howard Gardner. Theodore Ryland Sizer: 23 June 1932–21 October 2009. Proceedings of the American Philosophical Society, 156(2):261– 265, June 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23558199.

Garrow:2013:NDK

[Gar13a] David J. Garrow. Nicholas deb. Katzenbach: 17 January 1922– 8 May 2012. Proceedings of the American Philosophical Society, 157(4):471–477, December 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640229.

Garwin:2013:WKH

[Gar13b] Richard L. Garwin. Wolfgang Kurt Hermann Panofsky: 24 April 1919–24 September 2007. Proceedings of the American Philosophical Society, 157(4):479–485, December 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/24640230.

Garwin:2016:IDS

[Gar16] Richard L. Garwin. The Iran deal in six minutes. Proceedings of the American Philosophical Society, 160(2):201–204, June 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Garwin.pdf.

Gardner:2017:JSB

[Gar17] Howard Gardner. Jerome Seymour Bruner 1 October 1915–5 June 2016. Proceedings of the American Philosophical Society, 161(4):353– 361, December 2017. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2018-03/attachments/Bruner.pdf.

Garrett:2018:NPI

[Gar18] Andrew Garrett. New perspectives on Indo-European phylogeny and chronology. Proceedings of the American Philosophical Society, 162 (1):27–40, March 2018. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2018-08/attachments/Garrett.pdf.

Gerken:2015:RVC

[GBG15] Heather K. Gerken, Robert F. Bauer, and Benjamin L. Ginsberg. The right to vote: A conversation with the co-chairs of the Presidential Commission on Election Administration. Proceedings of the American Philosophical Society, 159(1):17–31, March 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Right%20to%20Vote.pdf;
 https://www.jstor.org/stable/24640168.

Gray:2014:HTN

[GC14] Harry B. Gray and James P. Collman. Henry Taube: 30 November 1915–16 November 2005. Proceedings of the American Philosophical Society, 158(2):195–198, June 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Taube.pdf;   https://www.jstor.org/stable/24640207.

Gelletly:2001:RIB

[Gel01] William Gelletly. Radioactive ion beams: A new window on atomic nuclei. Proceedings of the American Philosophical Society, 145(4):519–554, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558191.

Georgi:2007:ABP

[Geo07] Howard Georgi. Abraham “Bram” Pais, 19 May 1918–28 July 2000. Proceedings of the American Philosophical Society, 151(3):391–394, September 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4599083.

Gerken:2015:RPC

[Ger15] Heather K. Gerken. The real problem with “Citizens United”: Campaign finance, dark money, and shadow parties. Proceedings of the American Philosophical Society, 159(1):5–16, March 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Gerken.pdf;   https://
www.jstor.org/stable/24640167.

Goldstine:2001:ERP

[GG01] Herman H. Goldstine and Ralph E. Gomory. Emanuel Ruben Piore, 19 July 1908–9 May 2000. Proceedings of the American Philosophical Society, 145(3):359–363, September 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/1558115.

Glazer:2013:ESJ

[GH13] Alexander N. Glazer and Robert L. Hill. Emil L. Smith: 5 July 1911–31 May 2009. Proceedings of the American Philosophical Society, 157(1):127–141, March 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558146.

Goodman:2019:ODD

[GHX19] Leo A. Goodman, Robert M. Hauser, and Yu Xie. Otis Dudley Duncan. Proceedings of the American Philosophical Society, 163(2):??, June 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.amphilsoc.org/sites/default/
files/2020-03/attachments/Duncan.pdf.

Gibbard:2014:NC

[Gib14] Allan Gibbard. Normative consistency. Proceedings of the American Philosophical Society, 158(1):10–17, March 2014. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Gibbard.pdf;   https://www.jstor.org/stable/
24640140.

Giegengack:2015:CCM

[Gie15] Robert Giegengack. The Carrington coronal mass ejection of 1859. Proceedings of the American Philosophical Society, 159(4):421–433, December 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL http://live-amphil.pantheonsite.io/
sites/default/files/2017-10/attachments/Giegengack.pdf;
https://www.jstor.org/stable/26159195.

Gilmore:2008:TRM

[Gil08] David D. Gilmore. “Tarasca”: Ritual monster of Spain. Proceedings of the American Philosophical Society, 152(3):362–382, September 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/40541592.

Gillman:2014:ABR

[Gil14] Derek Gillman. Albert Barnes and the rejection of history. Proceedings of the American Philosophical Society, 158(2):107–119, June 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Gillman.pdf;   https://
www.jstor.org/stable/24640198.

Gingerich:2005:BCM

[Gin05] Owen Gingerich. I. Bernard Cohen, 1 March 1914–20 June 2003. Proceedings of the American Philosophical Society, 149(3):395–398, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598943.

Ginsburg:2009:APF

[Gin09] Jane C. Ginsburg. The author’s place in the future of copyright. Proceedings of the American Philosophical Society, 153(2):147–159, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/40541655.

Gingerich:2010:FLW

[Gin10] Owen Gingerich. Fred Lawrence Whipple: 5 November 1906–30 August 2004. Proceedings of the American Philosophical Society, 154(3):385–389, September 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/41000090.

Gingerich:2011:GIT

[Gin11] Owen Gingerich. Galileo, the impact of the telescope, and the birth of modern astronomy. Proceedings of the American Philosophical Society, 155(2):134–141, June 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056830.

Gingerich:2012:EWL

[Gin12] Philip D. Gingerich. Evolution of whales from land to sea. Proceedings of the American Philosophical Society, 156(3):309–323, September 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23558092.

Gourevitch:2017:SHN

[GK17] Peter A. Gourevitch and Robert O. Keohane. Stanley Hoffmann 27 November 1928–13 September 2015. Proceedings of the American Philosophical Society, 161(1):97–105, March 2017. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2017-07/attachments/Hoffmann.pdf.

Gleason:2002:AUA

[Gle02] Abbott Gleason. Adam Ulam: 8 April 1922–28 March 2000. Proceedings of the American Philosophical Society, 146(4):415–418, December 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558317.

Glymph:2016:IDB

[Gly16] Thavolia Glymph. “Invisible disabilities”: Black women in war and in freedom. Proceedings of the American Philosophical Society, 160 (3):237–246, September 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.org/

sites/default/files/2017-07/attachments/Glymph.pdf.

Goheen:2005:FTJ

[Goh05] Robert F. Goheen. Frank E. Taplin, 22 June 1915–11 May 2003. Proceedings of the American Philosophical Society, 149(3):447–450, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598950.

Goldsby:2012:TLC

[Gol12] Richard A. Goldsby. Theodore Lamont Cross: 12 February 1924– 28 February 2010. Proceedings of the American Philosophical Society, 156(1):75–80, March 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558080.

Gomory:2014:CG

[Gom14] Ralph Gomory. Corporate goals. Proceedings of the American Philosophical Society, 158(4):392–399, December 2014. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Gomory.pdf;   https://www.jstor.org/stable/24640183.

Gordon:2008:FSD

[Gor08] Terri J. Gordon. Film in the second degree: “Cabaret” and the dark side of laughter. Proceedings of the American Philosophical Society, 152(4):440–465, December 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541603.

Gossman:2001:MNH

[Gos01] Lionel Gossman. Michelet and natural history: The alibi of nature. Proceedings of the American Philosophical Society, 145(3):283–333, September 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558110.

Gossett:2012:GVI

[Gos12] Philip Gossett. Giuseppe Verdi and the Italian Risorgimento. Proceedings of the American Philosophical Society, 156(3):271–282, September 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23558088.

Gossett:2016:SPS

[Gos16] Suzanne Gossett. Shakespeare from page to stage. Proceedings of the American Philosophical Society, 160(2):183–200, June 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Gossett.pdf.

Gottfried:2008:VFW

[Got08] Kurt Gottfried. Victor F. Weisskopf, 19 September 1908–22 April 2002. Proceedings of the American Philosophical Society, 152(1):163– 166, March 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/25478480.

Gould:2008:JBD

[Gou08] Andrew P. Gould. John Bahcall, 30 December 1934–17 August 2005. Proceedings of the American Philosophical Society, 152(1):103–107, March 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/25478471.

Gould:2011:RIS

[Gou11] John Wells Gould. Robert Imbrie Smith: 23 February 1931–1 November 2008. Proceedings of the American Philosophical Society, 155(1):109–114, March 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056853.

Gursky:2002:HFJ

[GP02] Herbert Gursky and Frank Press. Herbert Friedman, 21 June 1916– 9 September 2000. Proceedings of the American Philosophical Society, 146(2):195–203, June 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558202.

Greenstein:2006:PHO

[GR06] Fred I. Greenstein and Austin Ranney. Pendleton Herring, 27 October 1903–17 August 2004. Proceedings of the American Philosophical Society, 150(3):487–492, September 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/4599014.

Graham:2002:ERW

[Gra02a] Patricia Albjerg Graham. Educational reform: Why now? Proceedings of the American Philosophical Society, 146(3):256–263, September 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558167.

Grant:2002:WDH

[Gra02b] Peter R. Grant. William D. Hamilton: 1 August 1936–7 March 2000. Proceedings of the American Philosophical Society, 146(4):387–394, December 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558313.

Gray:2002:EHL

[Gra02c] Hanna Holborn Gray. Edward Hirsch Levi, 26 June 1911–7 March 2000. Proceedings of the American Philosophical Society, 146(3): 297–302, September 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558171.

Grant:2008:MCJ

[Gra08] Edward Grant. Marshall Clagett, 23 January 1916–21 October 2005. Proceedings of the American Philosophical Society, 152(1):121–127, March 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/25478473.

Grabar:2009:WBB

[Gra09] Oleg Grabar. When is a bird a bird? Proceedings of the American Philosophical Society, 153(3):247–253, September 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/40541669.

Grabar:2011:IIS

[Gra11] Oleg Grabar. Ihor Ivanovich Ševčenko: 10 February 1922–26 December 2009. Proceedings of the American Philosophical Society, 155(4):513–517, December 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23208789.

Graham:2013:PNC

[Gra13] Loren Graham. The power of names: In culture and in mathematics. Proceedings of the American Philosophical Society, 157(2):229–234, June 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/24640241.

Gray:2014:OUG

[Gra14] Hanna Holborn Gray. Overgrowth and undergrowth in the groves of academe. Proceedings of the American Philosophical Society, 158(2): 128–135, June 2014. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Gray_0.pdf;
https://www.jstor.org/stable/24640200.

Gray:2015:CSA

[Gra15] Harry B. Gray. The 21st Century solar army. Proceedings of the American Philosophical Society, 159(4):361–366, December 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Gray.pdf;   https://www.
jstor.org/stable/26159191.

Graham:2016:ER

[Gra16] Loren Graham. Epigenetics and Russia. Proceedings of the American Philosophical Society, 160(3):266–271, September 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2017-07/attachments/Graham.pdf.

Grant:2019:JB

[Gra19] Peter R. Grant. John T. Bonner. Proceedings of the American Philosophical Society, 163(4):??, December 2019. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2021-
01/attachments/Bonner.pdf.

Greenhouse:2004:BWF

[Gre04a] Linda Greenhouse. “Because we are final” judicial review two hundred years after Marbury. Proceedings of the American Philosophical Society, 148(1):38–52, March 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558243.

Greenhouse:2004:BP

[Gre04b] Linda Greenhouse. The Blackmun papers. Proceedings of the American Philosophical Society, 148(3):332–357, September 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1558230.

Greenhouse:2009:IFI

[Gre09] Linda Greenhouse. Introduction: The future of information. Proceedings of the American Philosophical Society, 153(2):131–133, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/40541652.

Greenhouse:2010:ICS

[Gre10a] Carol J. Greenhouse. Introduction: Cultural subjects and objects: The legacy of Franz Boas and its futures in anthropology, academe, and human rights. Proceedings of the American Philosophical Society, 154(1):1–7, March 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/20721522.

Greenhouse:2010:HDJ

[Gre10b] Linda Greenhouse. How do judges know what they know? Proceedings of the American Philosophical Society, 154(3):287–293, September 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/41000081.

Greenhouse:2019:DBF

[Gre19] Linda Greenhouse. Dinner with Ben Franklin: The origins of the American Philosophical Society. Proceedings of the American Philosophical Society, 163(1):1–9, March 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

amphilsoc.org/sites/default/files/2019-11/attachments/Greenhouse.
pdf.

Grosz:2005:BMM

[Gro05] Barbara J. Grosz. Beyond mice and menus. Proceedings of the American Philosophical Society, 149(4):529–543, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/4598959.

Gordon:2013:MMO

[GTCD13] Ann B. Gordon, Samuel O. Thier, John E. Craig, and Karen Davis. Margaret E. Mahoney: 24 October 1924–22 December 2011. Proceedings of the American Philosophical Society, 157(2):249–253, June 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/24640244.

Guillemin:2002:AEU

[Gui02] Jeanne Guillemin. The 1979 anthrax epidemic in the USSR: Applied science and political controversy. Proceedings of the American Philosophical Society, 146(1):18–36, March 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1558154.

Gutas:2005:FRA

[Gut05] Dimitri Gutas. Franz Rosenthal, 31 August 1914–8 April 2003. Proceedings of the American Philosophical Society, 149(3):441–446, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598949.

Gutmann:2014:FWH

[Gut14] Amy Gutmann. The fundamental worth of higher education. Proceedings of the American Philosophical Society, 158(2):136–143, June 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Gutmann.pdf;   https://
www.jstor.org/stable/24640201.

Gutmann:2015:FSH

[Gut15] Amy Gutmann. Francis Sheldon Hackney: 5 December 1933–12 September 2013. Proceedings of the American Philosophical Society, 159(3):335–341, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Hackney.pdf;   https://www.jstor.org/stable/26159160.

Hackney:2001:CVW

[Hac01] Sheldon Hackney. C. Vann Woodward, 13 November 1908–17 December 1999. Proceedings of the American Philosophical Society, 145(2):233–240, June 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558279.

Hackney:2009:MMN

[Hac09] Sheldon Hackney. Martin Meyerson: 14 November 1922–2 June 2007. Proceedings of the American Philosophical Society, 153(3):359–363, September 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541678.

Haggard:2002:NAL

[Hag02] Robert F. Haggard. The Nicola Affair: Lewis Nicola, George Washington, and American military discontent during the Revolutionary War. Proceedings of the American Philosophical Society, 146(2):139– 169, June 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558199.

Haggard:2009:PFD

[Hag09] Robert F. Haggard. The politics of friendship: Du Pont, Jefferson, Madison, and the physiocratic dream for the New World. Proceedings of the American Philosophical Society, 153(4):419–440, December 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/20721511.

Halley:2018:JOR

[Hal18] Matthew R. Halley. Jefferson’s ornithology reconsidered. Proceedings of the American Philosophical Society, 162(3):231–258, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2019-03/attachments/Halley.pdf.

Hamburg:2007:SBJ

[Ham07] David A. Hamburg. Sune Bergstrom, 10 January 1916–15 August 2004. Proceedings of the American Philosophical Society, 151(1):91– 96, March 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4599046.

Hamann:2008:HMH

[Ham08a] Byron Ellsworth Hamann. How Maya hieroglyphs got their name: Egypt, Mexico, and China in Western grammatology since the Fifteenth Century. Proceedings of the American Philosophical Society, 152(1):1–68, March 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478468.

Hamburg:2008:AGM

[Ham08b] David A. Hamburg. Alexander L. George 31 May 1920–16 August 2006. Proceedings of the American Philosophical Society, 152 (4):547–551, December 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541608.

Harper:2010:JMS

[Har10] David Harper. John Maynard Smith, 6 January 1920–19 April 2004. Proceedings of the American Philosophical Society, 154(1):109–112, March 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/20721532.

Harris:2015:NRM

[Har15] Verne Harris. Nelson Rolihlahla Mandela: 18 July 1918–5 December 2013. Proceedings of the American Philosophical Society, 159(3):349– 353, September 2015. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Mandela.pdf;
https://www.jstor.org/stable/26159162.

Hauser:2020:WMA

[Hau20] Robert M. Hauser. What is the matter with America’s schools? Proceedings of the American Philosophical Society, 164(1):??, March 2020. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2021-12/attachments/Hauser.pdf.

Hauser:2006:WHS

[HCC06] Robert M. Hauser, Charles Camic, and Colleagues at the University of Wisconsin-Madison. William Hamilton Sewell, 27 November 1909–24 June 2001. Proceedings of the American Philosophical Society, 150(1):201–204, March 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598985.

Hauptman:2008:CDM

[HD08] Laurence M. Hauptman and Heriberto Dixon. Cadet David Moniac: A Creek Indian’s schooling at West Point, 1817–1822. Proceedings of the American Philosophical Society, 152(3):322–348, September 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/40541590.

Heaney:2004:JLT

[Hea04] Seamus Heaney. The Jayne Lecture: Title deeds: Translating a classic. Proceedings of the American Philosophical Society, 148 (4):411–426, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558136.

Heilbron:2005:FLH

[Hei05] J. L. Heilbron. Frederic Lawrence Holmes, 6 February 1932–27 March 2003. Proceedings of the American Philosophical Society, 149(2):255– 260, June 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598930.

Heilbron:2006:PMF

[Hei06] J. L. Heilbron. Plus and minus: Franklin’s zero-sum way of thinking. Proceedings of the American Philosophical Society, 150(4):607–617, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4599028.

Heilbron:2018:TFJ

[Hei18] John L. Heilbron. Tore Frängsmyr 8 July 1938–28 August 2017. Proceedings of the American Philosophical Society, 162(3):313–319, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2019-03/attachments/Frangsmyr.pdf.

Helinski:2005:WDM

[Hel05] Donald R. Helinski. William David McElroy, 22 January 1917– 17 February 1999. Proceedings of the American Philosophical Society, 149(1):103–106, March 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598915.

Henkin:2002:HWD

[Hen02] Louis Henkin. Herbert Wechsler, 4 December 1909–26 April 2000. Proceedings of the American Philosophical Society, 146(3):311–316, September 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558174.

Henkin:2004:JMG

[Hen04] Louis Henkin. John Marshall globalized. Proceedings of the American Philosophical Society, 148(1):53–63, March 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1558244.

Herdt:2005:TIP

[Her05] Robert W. Herdt. Treaties, intellectual property, market power, and food in the developing world. Proceedings of the American Philosophical Society, 149(4):487–514, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/4598957.

Hertz:2014:PFS

[Her14] Paul Hertz. The present and future of space science at NASA. Proceedings of the American Philosophical Society, 158(4):329–353, December 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Hertz.pdf;   https://www.
jstor.org/stable/24640179.

Herschbach:2017:JRK

[Her17] Dudley Herschbach. Jeremy Randall Knowles 2 April 1935–3 April 2008. Proceedings of the American Philosophical Society, 161(4):335– 342, December 2017. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2018-03/attachments/Knowles.pdf.

Hersch:2013:NAA

[HGG13] Matthew H. Hersch, Loren Graham, and Patricia Albjerg Graham. Neil A. Armstrong: 5 August 1930–25 August 2012. Proceedings of the American Philosophical Society, 157(3):347–353, September 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/24640399.

Hijiya:2000:GJR

[Hij00] James A. Hijiya. The “Gita” of J. Robert Oppenheimer. Proceedings of the American Philosophical Society, 144(2):123–167, June 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL http://www.amphilsoc.org/sites/default/files/
Hijiya.pdf;   https://www.jstor.org/stable/1515629.

Hilleman:2000:FJT

[Hil00] Maurice R. Hilleman. The frustrating journey toward an AIDS vaccine. Proceedings of the American Philosophical Society, 144 (4):349–360, December 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515616.

Hindle:2001:CBA

[Hin01] Brooke Hindle. Carl Bridenbaugh, 10 August 1903–6 January 1992. Proceedings of the American Philosophical Society, 145(3):345–347, September 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558112.

Hinshaw:2008:GFW

[Hin08] Robert Hinshaw. Gilbert F. White 26 November 1911–5 October 2006. Proceedings of the American Philosophical Society, 152(3):419–423, September 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541597.

Hiraizumi:2000:MLT

[Hir00] Wataru Hiraizumi. Mass longevity transforms our society. Proceedings of the American Philosophical Society, 144(4):361–383, December 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1515617.

Hindle:2014:KHO

[HKT14] Win Hindle, Peter Koch, and Dan Tymann. Kenneth H. Olsen: 20 February 1926–6 February 2011. Proceedings of the American Philosophical Society, 158(3):293–303, September 2014. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Olsen.pdf;   https://www.jstor.org/stable/24640132.

Hastings:2006:LBA

[HM06] J. Woodland Hastings and Laurens Mets. Lawrence Bogorad, 29 August 1921–28 December 2003. Proceedings of the American Philosophical Society, 150(2):341–345, June 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/4598994.

Haug-Moritz:2008:HRE

[HM08] Gabriele Haug-Moritz. The Holy Roman Empire, the Schmalkald League, and the idea of confessional nation-building. Proceedings of the American Philosophical Society, 152(4):427–439, December 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/40541602.

Henry:2014:FOI

[HO14] Martin Henry and Franz Overbeck. Franz Overbeck’s inaugural lecture in Basel (1870). Proceedings of the American Philosophical Society, 158(3):248–277, September 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Henry.pdf;   https://www.jstor.org/stable/24640129.

Hodgson:2012:JKG

[Hod12] Godfrey Hodgson. John Kenneth Galbraith: 15 October 1908– 29 April 2006. Proceedings of the American Philosophical Society, 156(4):443–450, December 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558237.

Hoffman:2013:CWB

[Hof13] David C. Hoffman. “The creation we behold”: Thomas Paine’s “The Age of Reason” and the tradition of physico-theology. Proceedings of the American Philosophical Society, 157(3):281–303, September 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/24640397.

Hogg:2014:DSH

[Hog14] David E. Hogg. David Sutphin Heeschen: 12 March 1926–13 April 2012. Proceedings of the American Philosophical Society, 158(4):429– 434, December 2014. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Heeschen.pdf;
 https://www.jstor.org/stable/24640187.

Holton:2004:RKM

[Hol04] Gerald Holton. Robert K. Merton, 4 July 1910–23 February 2003. Proceedings of the American Philosophical Society, 148(4):505–517, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558145.

Holldobler:2010:MLD

[Höl10] Bert Hölldobler. Martin Lindauer: 19 December 1918–13 November 2008. Proceedings of the American Philosophical Society, 154 (4):485–491, December 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056867.

Holt:2016:LDR

[Hol16] Thomas C. Holt. Legacies “disremembered:” re-reading moments of emancipation. Proceedings of the American Philosophical Society, 160 (3):247–253, September 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.org/

sites/default/files/2017-07/attachments/Holt.pdf.

Hope:2014:EHG

[Hop14] Charles Hope. Ernst Hans Gombrich: 30 March 1909–3 November 2001. Proceedings of the American Philosophical Society, 158(2): 161–165, June 2014. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Gombrich.pdf;
 https://www.jstor.org/stable/24640204.

Horowitz:2014:DRS

[Hor14] Daniel Horowitz. David Riesman: 22 September 1909–10 May 2002. Proceedings of the American Philosophical Society, 158(2):173–193, June 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Riesman.pdf;   https://
www.jstor.org/stable/24640206.

Horowitz:2015:RDR

[Hor15] Alexandra Horowitz. Reading dogs reading us. Proceedings of the American Philosophical Society, 159(2):141–155, June 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic).  URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Horowitz.pdf;   https:/
/www.jstor.org/stable/24640211.

Happer:2006:RHD

[HP06] W. Happer and P. J. E. Peebles. Robert Henry Dicke, 6 May 1916–4 March 1997. Proceedings of the American Philosophical Society, 150(1):181–186, March 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598981.

Huang:2006:CPR

[Hua06] Nian-Sheng Huang. From the “fur cap” to Poor Richard: The Chinese connection. Proceedings of the American Philosophical Society, 150(2):205–240, June 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598988.

Humphrey:2007:AF

[Hum07] Caroline Humphrey. Alternative freedoms. Proceedings of the American Philosophical Society, 151(1):1–10, March 2007. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/4599040.

Hunger:2007:HHD

[Hun07] Hermann Hunger. Herbert Hunger, 9 December 1914–9 July 2000. Proceedings of the American Philosophical Society, 151(4):457–460, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/25478460.

Hunt:2009:EWA

[Hun09] Lynn Hunt. Eugen Weber 24 April 1925–17 May 2007. Proceedings of the American Philosophical Society, 153(2):239–242, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/40541665.

Huxley:2001:SAH

[Hux01] Andrew Huxley. Sir Alan Hodgkin, 5 February 1914–20 December 1998. Proceedings of the American Philosophical Society, 145(1):97– 105, March 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558328.

Isaacson:2006:GM

[Isa06] Walter Isaacson. Great minds. Proceedings of the American Philosophical Society, 150(4):568–574, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/4599025.

Imperial:2007:UCR

[ISO^{+}07] Julita S. Imperial, Natalie Silverton, Baldomero M. Olivera, Pradip K. Bandyopadhyay, Annett Sporning, Michael Ferber, and Heinrich Terlau. Using chemistry to reconstruct evolution: On the origins of fish-hunting in venomous cone snails. Proceedings of the American Philosophical Society, 151(2):185–200, June 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/4599057.

Jablonski:2012:HSP

[Jab12] Nina G. Jablonski. Human skin pigmentation as an example of adaptive evolution. Proceedings of the American Philosophical Society, 156(1):45–57, March 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558077.

Jacob:2003:LLR

[Jac03] Maurice Jacob. Louis Leprince-Ringuet, 27 March 1901–23 December 2000. Proceedings of the American Philosophical Society, 147(2):167– 173, June 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558260.

Jackson:2010:BEC

[Jac10] Jason Baird Jackson. Boasian ethnography and contemporary intellectual property debates. Proceedings of the American Philosophical Society, 154(1):40–49, March 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/20721526.

Jacobs:2017:MIY

[Jac17] Irwin Mark Jacobs. The magic inside your cell phone and its broad societal impact. Proceedings of the American Philosophical Society, 161 (2):138–154, June 2017. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Jacobs.pdf.

Jacob:2020:EOT

[Jac20] Margaret C. Jacob. Enlightened outrage: The three impostors. Proceedings of the American Philosophical Society, 164(2):??, ???? 2020. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2022-06/attachments/Jacob.pdf.

Jamieson:2004:WHA

[Jam04] Kathleen Hall Jamieson. Walter H. Annenberg, 13 March 1908– 1 October 2002. Proceedings of the American Philosophical Society, 148(2):215–220, June 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558286.

Jamieson:2011:LAF

[Jam11] Kathleen Hall Jamieson. Leonore Annenberg: 20 February 1918– 12 March 2009. Proceedings of the American Philosophical Society, 155(1):91–96, March 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056850.

Jamieson:2015:IDF

[Jam15] Kathleen Hall Jamieson. Implications of the demise of “fact” in political discourse. Proceedings of the American Philosophical Society, 159 (1):66–84, March 2015. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Jamieson.pdf;
 https://www.jstor.org/stable/24640171.

Jamieson:2019:HRH

[Jam19] Kathleen Hall Jamieson. How Russian hackers and trolls exploited U.S. media in 2016. Proceedings of the American Philosophical Society, 163(2):??, June 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.org/

sites/default/files/2020-03/attachments/Jamieson.pdf.

Jaynes:2014:RMW

[Jay14] Gerald D. Jaynes. Robin M. Williams, Jr.: 11 October 1914–3 June 2006. Proceedings of the American Philosophical Society, 158(1): 101–105, March 2014. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Williams.pdf;
 https://www.jstor.org/stable/24640147.

Jeffrey:2000:CGH

[JB00] Richard Jeffrey and Paul Benacerraf. Carl Gustav Hempel, 8 January 1905–9 November 1997. Proceedings of the American Philosophical Society, 144(1):95–98, March 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515608.

Jeffery:2016:KLF

[Jef16] Peter Jeffery. Kenneth Levy 26 February 1927–15 August 2013. Proceedings of the American Philosophical Society, 160(2):227–231, June 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Levy.pdf.

Jacob:2018:JOA

[JH18] Margaret Jacob and Lynn Hunt. Joyce Oldham Appleby 9 April 1929–23 December 2016. Proceedings of the American Philosophical Society, 162(2):207–211, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.

org/sites/default/files/2018-11/attachments/Appleby.pdf.

Johnson:2006:DBT

[Joh06] Richard A. Johnson. David Bicknell Truman, 1 June 1913–28 August 2003. Proceedings of the American Philosophical Society, 150(4):697–702, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599037.

Johns:2016:LAS

[Joh16] Adrian Johns. Lay assaying and the scientific citizen. Proceedings of the American Philosophical Society, 160(1):18–25, March 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Johns.pdf.

Jones:2008:HBA

[Jon08] Christopher P. Jones. Herbert Bloch 18 August 1911–6 September 2006. Proceedings of the American Philosophical Society, 152 (4):533–540, December 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541606.

Jones:2016:SF

[Jon16] Christopher P. Jones. A syntax of forgery. Proceedings of the American Philosophical Society, 160(1):26–36, March 2016. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Jones.pdf.

Jones:2018:LOP

[Jon18] Alexander Jones. “Like opening a pyramid and finding an atomic bomb”: Derek de Solla Price and the Antikythera mechanism. Proceedings of the American Philosophical Society, 162(3):259–294, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2019-03/attachments/Jones.pdf.

Jorgenson:2006:RLM

[Jor06] Dale W. Jorgenson. Ralph Landau, 19 May 1916–6 April 2004. Proceedings of the American Philosophical Society, 150(1):191–195, March 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4598983.

Jordan:2011:CRP

[Jor11] William Chester Jordan. Count Robert’s “pet” wolf. Proceedings of the American Philosophical Society, 155(4):404–417, December 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/23208782.

Joseph:2018:IEE

[Jos18] Brian D. Joseph. The Indo-Europeanization of Europe: An introduction to the issues. Proceedings of the American Philosophical Society, 162(1):17–26, March 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.org/

sites/default/files/2018-08/attachments/Joseph.pdf.

Kaiser:2007:AWT

[Kai07] Karl Kaiser. America in the world today: A European view. Proceedings of the American Philosophical Society, 151(4):419–424, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/25478454.

Karmel:2005:JKV

[Kar05] Pepe Karmel. J. Kirk T. Varnedoe, 18 January 1946–14 August 2003. Proceedings of the American Philosophical Society, 149(4):631–638, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598967.

Katz:2000:JWH

[Kat00] Stanley N. Katz. James Willard Hurst, 6 October 1910–18 June 1997. Proceedings of the American Philosophical Society, 144(4):459–463, December 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1515621.

Katz:2005:FCR

[Kat05a] Samuel L. Katz. Frederick Chapman Robbins, 25 August 1916–4 August 2003. Proceedings of the American Philosophical Society, 149(4):615–618, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598964.

Katz:2005:WDI

[Kat05b] Stanley N. Katz. What does it mean to say that philanthropy is “effective”? The philanthropists’ new clothes. Proceedings of the American Philosophical Society, 149(2):123–131, June 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/4598921.

Keasling:2012:EBD

[Kea12] Jay D. Keasling. Engineering biology for drugs and fuels. Proceedings of the American Philosophical Society, 156(3):283–294, September 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23558089.

Keller:2000:PBT

[Kel00] Christian B. Keller. Philanthropy betrayed: Thomas Jefferson, the Louisiana Purchase, and the origins of Federal Indian removal policy. Proceedings of the American Philosophical Society, 144(1):39–66, March 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1515604.

Kelly:2001:BOJ

[Kel01] Donald H. Kelly. Brian O’Brien, 2 January 1898–1 July 1992. Proceedings of the American Philosophical Society, 145(3):355–357, September 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558114.

Kennedy:2003:KBJ

[Ken03] Eugene P. Kennedy. Konrad Bloch, 12 January 1912–15 October 2000. Proceedings of the American Philosophical Society, 147(1):67– 72, March 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558128.

Kennedy:2015:NIS

[Ken15a] Ross Kennedy. “A net of intrigue and selfish rivalry”: Woodrow Wilson and power politics during World War I. Proceedings of the American Philosophical Society, 159(2):156–161, June 2015. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/

attachments/Kennedy.pdf;   https://www.jstor.org/stable/
24640212.

Kennel:2015:HGW

[Ken15b] Charles F. Kennel. “Hiatus” in global warming: Paradox and complexity in climate science. Proceedings of the American Philosophical Society, 159(4):367–408, December 2015. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Kennel.pdf;   https://www.jstor.org/stable/26159192.

Kennel:2016:EFJ

[Ken16] Charles F. Kennel. Edward A. Frieman 19 January 1926–11 April 2013. Proceedings of the American Philosophical Society, 160(3):283– 293, September 2016. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-07/attachments/Frieman.pdf.

Kerman:2000:MPC

[Ker00] Joseph Kerman. Music and politics: The case of William Byrd (1540–1623). Proceedings of the American Philosophical Society, 144(3):275–287, September 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515590.

Kerber:2009:WDR

[Ker09] Linda K. Kerber. Why diamonds really are a girl’s best friend. Proceedings of the American Philosophical Society, 153(1):56–66, March 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/40541629.

Kerber:2016:PUW

[Ker16] Linda K. Kerber. The past unshackled: When revolutions go backward. Proceedings of the American Philosophical Society, 160(3):233– 236, September 2016. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-07/attachments/Kerber.pdf.

Kevles:2008:PPP

[Kev08] Daniel J. Kevles. Protections, privileges, and patents: Intellectual property in American horticulture, the Late Nineteenth Century to 1930. Proceedings of the American Philosophical Society, 152(2):207– 213, June 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/25478487.

Keynes:2003:JLB

[Key03] Richard D. Keynes. The Jayne Lecture. From bryozoans to tsunami: Charles Darwin’s findings on the Beagle. Proceedings of the American Philosophical Society, 147(2):103–127, June 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1558253.

Kyte:2020:RD

[KFW20] Jack Kyte, Robert Fahey, and Joseph Watson. Russell Doolittle. Proceedings of the American Philosophical Society, 164(2):??, ???? 2020. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2022-06/attachments/Doolittle.pdf.

Kitagawa:2014:BCJ

[KG14] Teizo Kitagawa and Thomas Goplen. Bryce L. Crawford, Jr.: 27 November 1914–16 September 2011. Proceedings of the American Philosophical Society, 158(3):317–321, September 2014. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Crawford.pdf;   https://www.jstor.org/stable/
24640135.

Kennedy:2008:CPO

[KHEP08] Donald Kennedy, Albert Hastorf, David Epel, and David Perkins. Colin S. Pittendrigh, 13 October 1918–19 March 1996. Proceedings of the American Philosophical Society, 152(1):157–161, March 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/25478479.

Kihlstrom:2003:ERH

[Kih03] John F. Kihlstrom. Ernest Ropiequet Hilgard, 25 July 1904–22 October 2001. Proceedings of the American Philosophical Society, 147(4):399–404, December 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558303.

Kilbourne:2004:VPD

[Kil04] Edwin D. Kilbourne. Virus paleontology, disease, and evolution: An introduction. Proceedings of the American Philosophical Society, 148(3):261–263, September 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558225.

Kinney:2001:RAS

[Kin01] Dale Kinney. Roman architectural spolia. Proceedings of the American Philosophical Society, 145(2):138–161, June 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1558268.

Kirby:2003:HSA

[Kir03] Alec Kirby. Harold Stassen, 13 April 1907–4 March 2001. Proceedings of the American Philosophical Society, 147(3):317–324, September 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558222.

Kirshner:2013:AUN

[Kir13] Robert P. Kirshner. The accelerating universe: A Nobel surprise. Proceedings of the American Philosophical Society, 157(4):438–456, December 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/24640226.

Kitsikopoulos:2000:TCM

[Kit00] Harry Kitsikopoulos. Technological change in Medieval England: A critique of the neo-Malthusian argument. Proceedings of the American Philosophical Society, 144(4):397–449, December 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1515619.

Kitsikopoulos:2013:HNC

[Kit13] Harry Kitsikopoulos. From hero to Newcomen: The critical scientific and technological developments that led to the invention of the steam engine. Proceedings of the American Philosophical Society, 157(3):304–344, September 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640398.

Klein:2003:TCO

[KK03] George Klein and Eva Klein. Torbjörn Caspersson, 15 October 1910– 7 December 1997. Proceedings of the American Philosophical Society, 147(1):73–75, March 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558129.

Klein:2011:PSM

[Kle11] Lawrence R. Klein. Paul A. Samuelson: 15 May 1915–13 December 2009. Proceedings of the American Philosophical Society, 155(4):501–511, December 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23208788.

Kleppner:2016:NFR

[Kle16] Daniel Kleppner. Norman Foster Ramsey 27 August 1915–4 November 2011. Proceedings of the American Philosophical Society, 160 (3):313–334, September 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.org/

sites/default/files/2017-07/attachments/Ramsey.pdf.

Knight:2008:JBM

[Kni08] Douglas A. Knight. James Barr 20 March 1924–14 October 2006. Proceedings of the American Philosophical Society, 152(3):385–390, September 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541593.

Knoll:2017:DMR

[Kno17] Andrew H. Knoll. David Malcolm Raup 24 April 1933–9 July 2015. Proceedings of the American Philosophical Society, 161(2):177–182, June 2017. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Raup.pdf.

Kolenda:2001:MNS

[Kol01] Pauline Kolenda. Mysore Narasimhachar Srinivas, 16 November 1916–30 November 1999. Proceedings of the American Philosophical Society, 145(3):373–378, September 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1558118.

Koptur:2006:HGB

[Kop06] Suzanne Koptur. H. G. Baker, 23 February 1920–2 July 2001. Proceedings of the American Philosophical Society, 150(2):335–339, June 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/4598993.

Kornberg:2000:FSD

[Kor00] Hans Kornberg. Frederick Sydney Dainton (Baron Dainton of Hallam Moors in South Yorkshire), 11 November 1914–5 December 1997. Proceedings of the American Philosophical Society, 144(2):207–212, June 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1515633.

Koshland:2005:GSA

[Kos05] Daniel E. Koshland, Jr. Glenn Seaborg, 19 April 1912–25 February 1999. Proceedings of the American Philosophical Society, 149 (4):619–623, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598965.

Kottak:2000:CWW

[Kot00] Conrad P. Kottak. Charles Walter Wagley, 9 November 1913–25 November 1991. Proceedings of the American Philosophical Society, 144(1):119–122, March 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515612.

Kovach:2009:FN

[Kov09] Bill Kovach. The future of news. Proceedings of the American Philosophical Society, 153(2):134–139, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/40541653.

Kramer:2004:UMV

[Kra04] Larry Kramer. Understanding Marbury v. Madison. Proceedings of the American Philosophical Society, 148(1):14–26, March 2004. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1558241.

Kroto:2001:GHD

[Kro01] H. W. Kroto. Gerhard Herzberg, 25 December 1904–3 March 1999. Proceedings of the American Philosophical Society, 145(2):197–201, June 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558273.

Kennel:2012:CCC

[KRV12] Charles F. Kennel, V. Ramanathan, and David G. Victor. Coping with climate change in the next half-century. Proceedings of the American Philosophical Society, 156(4):398–415, December 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/23558233.

Kuklick:2014:KL

[Kuk14] Bruce Kuklick. Killing Lumumba. Proceedings of the American Philosophical Society, 158(2):144–152, June 2014. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Kuklick.pdf;   https://www.jstor.org/stable/
24640202.

Lakin:2011:SML

[Lak11] Jason M. Lakin. Seymour Martin Lipset: 18 March 1922–31 December 2006. Proceedings of the American Philosophical Society, 155(3):335–350, September 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23208691.

Lamoreaux:2002:HBG

[Lam02] Steve K. Lamoreaux. Hendrik Brugt Gerhard Casimir, 15 July 1909– 4 May 2000. Proceedings of the American Philosophical Society, 146(3):285–290, September 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558169.

Lang:2002:EDT

[Lan02] Mabel L. Lang. Emily Dickinson Townsend Vermeule: 11 August 1928–6 February 2001. Proceedings of the American Philosophical Society, 146(4):425–429, December 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/1558319.

Landes:2005:SJH

[Lan05] David S. Landes. Sir John Habakkuk, 13 May 1915–3 November 2002. Proceedings of the American Philosophical Society, 149(1):99– 101, March 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598914.

Lantieri:2011:FTL

[Lan11] Laurent A. Lantieri. Face transplant: Learning from the past, facing the future. Proceedings of the American Philosophical Society, 155(1):23–28, March 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056846.

Langhauser:2016:VMO

[Lan16] Derek Langhauser. Vincent L. McKusick 21 October 1921–3 December 2014. Proceedings of the American Philosophical Society, 160(1):109–115, March 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
VL_McKusick.pdf.

Lowrey:2013:RIS

[LB13] Ying Lowrey and William J. Baumol. Rapid invention, slow industrialization, and the absent innovative entrepreneur in Medieval China. Proceedings of the American Philosophical Society, 157(1):1– 21, March 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23558137.

LeBlanc:2006:DBM

[Le 06] Charles Le Blanc. Derk Bodde, 9 March 1909–3 November 2003. Proceedings of the American Philosophical Society, 150(1):161–165, March 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4598978.

Leavell:2003:MMH

[Lea03] Linda Leavell. Marianne Moore, her family, and their language. Proceedings of the American Philosophical Society, 147(2):140–149, June 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/1558256.

Leal:2012:AOS

[Lea12] Walter S. Leal. An acute olfactory system is essential for reproduction — the raison d’être for adult insects. Proceedings of the American Philosophical Society, 156(3):295–302, September 2012. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/23558090.

Lean:2015:ESS

[Lea15a] Judith L. Lean. Everyday solar storms … and their everyday impacts. Proceedings of the American Philosophical Society, 159 (4):434–452, December 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Lean.pdf;   https://www.jstor.org/stable/26159196.

Leavell:2015:FMW

[Lea15b] Perry Leavell. Finding a man who will take advice: Woodrow Wilson and Edward House. Proceedings of the American Philosophical Society, 159(1):57–65, March 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Leavell.pdf;   https://www.jstor.org/stable/24640170.

Lepore:2016:AA

[Lep16] Jill Lepore. Asymmetry in the archives. Proceedings of the American Philosophical Society, 160(2):178–182, June 2016. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Lepore.pdf.

Lester:2004:BVB

[Les04] Anthony Lester. Brown v. Board of Education overseas. Proceedings of the American Philosophical Society, 148(4):455–463, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/1558140.

Lessig:2007:CP

[Les07] Lawrence Lessig. The code of privacy. Proceedings of the American Philosophical Society, 151(3):283–290, September 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/4599071.

Levitt:2007:RFW

[Lev07] Martin L. Levitt. Russell F. Weigley, 2 July 1930–3 March 2004. Proceedings of the American Philosophical Society, 151(2):279–282, June 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4599068.

Lewis:2011:JHF

[Lew11] David Levering Lewis. John Hope Franklin: 2 January 1915–25 March 2009. Proceedings of the American Philosophical Society, 155(1):97– 100, March 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23056851.

Leopold:2005:GEU

[LHM05] Luna B. Leopold, Reed Huppman, and Andrew Miller. Geomorphic effects of urbanization in forty-one years of observation. Proceedings of the American Philosophical Society, 149(3):349–371, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/4598940.

Lidofsky:2001:CSW

[Lid01] Leon Lidofsky. Chien-Shiung Wu, 29 May 1912–16 February 1997. Proceedings of the American Philosophical Society, 145(1):115–126, March 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558330.

Liebman:2007:AFJ

[Lie07] Lance Liebman. Allan Farnsworth, 30 June 1928–31 January 2005. Proceedings of the American Philosophical Society, 151(4):445–449, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/25478458.

Lightfoot:2017:MLW

[Lig17] Jane L. Lightfoot. Martin Litchfield West 23 September 1937–13 July 2015. Proceedings of the American Philosophical Society, 161(3):285– 292, September 2017. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-12/attachments/West.pdf.

Linderberg:2003:POL

[Lin03] Jan Linderberg. Per-Olov Löwdin, 28 October 1916–6 October 2000. Proceedings of the American Philosophical Society, 147(2):175–180, June 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://www.amphilsoc.org/sites/default/
files/210.pdf;   https://www.jstor.org/stable/1558261.

Link:2008:PFM

[Lin08] Dorothea Link. Performing the fandango in Mozart’s “Le nozze di Figaro”. Proceedings of the American Philosophical Society, 152(2):167– 178, June 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/25478484.

Lipshitz:2006:EBL

[Lip06] Howard D. Lipshitz. Edward B. Lewis, 20 May 1918–21 July 2004. Proceedings of the American Philosophical Society, 150(2):379–395, June 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4598999.

Little:2010:RWP

[Lit10] Mark Little. Redefining what’s possible. Proceedings of the American Philosophical Society, 154(2):192–200, June 2010. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/41000098.

Lloyd-Jones:2001:AGR

[LJ01] Hugh Lloyd-Jones. Ancient Greek religion. Proceedings of the American Philosophical Society, 145(4):456–464, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1558184.

Leonard:2008:THB

[LL08] Janet L. Leonard and G. David Lange. Theodore Holmes Bullock, 16 May 1915–19 December 2005. Proceedings of the American Philosophical Society, 152(2):221–228, June 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/25478489.

Lo:2013:OBR

[Lo13] Andrew W. Lo. The origin of bounded rationality and intelligence. Proceedings of the American Philosophical Society, 157(3):269–280, September 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/24640396.

Longair:2011:DPA

[Lon11] Malcolm Longair. The discovery of pulsars and the aftermath. Proceedings of the American Philosophical Society, 155(2):147–157, June 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/23056832.

Lovei:2019:TJ

[Löv19] Gábor L. Lövei. Tibor Jermy. Proceedings of the American Philosophical Society, 163(1):??, March 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.

org/sites/default/files/2019-11/attachments/Lovei.pdf.

Lunine:2009:STS

[Lun09] Jonathan I. Lunine. Saturn’s Titan: A strict test for life’s cosmic ubiquity. Proceedings of the American Philosophical Society, 153(4):403–418, December 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/20721510.

Madonna:2015:WWS

[Mad15] G. Terry Madonna. William Warren Scranton: 19 July 1917–28 July 2013. Proceedings of the American Philosophical Society, 159 (2):227–231, June 2015. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Scranton.pdf;
 https://www.jstor.org/stable/24640219.

Malkiel:2009:CEM

[Mal09] Burton G. Malkiel. The Chinese economic miracle: Can it last? Proceedings of the American Philosophical Society, 153(2):193–199, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/40541658.

Malkiel:2017:KDW

[Mal17] Nancy Weiss Malkiel. “Keep the damned women out”: The struggle for coeducation in the Ivy League, the Seven Sisters, Oxford, and Cambridge. Proceedings of the American Philosophical Society, 161 (1):31–37, March 2017. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-07/attachments/Malkiel.pdf.

Mancall:2020:LHE

[Man20] Peter C. Mancall. Landscape and history in the early modern Atlantic basin. Proceedings of the American Philosophical Society, 164 (2):??, ???? 2020. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2022-06/attachments/Mancall.pdf.

Marchand:2001:GOD

[Mar01a] Suzanne Marchand. German orientalism and the decline of the West. Proceedings of the American Philosophical Society, 145(4):465–473, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558185.

Marcum:2001:DCG

[Mar01b] Deanna B. Marcum. Defining “common good” in the digital world. Proceedings of the American Philosophical Society, 145(1):73–80, March 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558326.

Mary:2003:JBB

[Mar03] Viscountess Eccles Mary. Julian Boyd and the battlefield of Hastings. Proceedings of the American Philosophical Society, 147(2):134–139, June 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558255.

Marty:2008:JJP

[Mar08] Martin E. Marty. Jaroslav Jan Pelikan 17 December 1923–13 May 2006. Proceedings of the American Philosophical Society, 152 (4):565–569, December 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541611.

Marcus:2009:JMW

[Mar09] Maeva Marcus. John Marshall was not the First Chief Justice. Proceedings of the American Philosophical Society, 153(1):48–55, March 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/40541628.

Marcy:2010:OES

[Mar10a] Geoffrey W. Marcy. Other Earths and the search for life in the Universe. Proceedings of the American Philosophical Society, 154 (4):422–438, December 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056861.

Marks:2010:PSW

[Mar10b] Arthur S. Marks. Palladianism on the Schuylkill: The work of Frederick Graff at Fairmount. Proceedings of the American Philosophical Society, 154(2):201–257, June 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/41000099.

Marks:2012:PPP

[Mar12] Arthur S. Marks. Private and public in “The Peale Family”: Charles Willson Peale as pater and painter. Proceedings of the American Philosophical Society, 156(2):109–187, June 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/23558190.

Marks:2013:JPP

[Mar13] Arthur S. Marks. Joining the past to the present: William Rush’s emblematic statuary at Fairmount. Proceedings of the American Philosophical Society, 157(2):176–228, June 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/24640240.

March:2015:MCN

[Mar15a] James G. March. Michel Crozier: 6 November 1922–23 May 2013. Proceedings of the American Philosophical Society, 159(2):201–205, June 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Crozier.pdf;   https://
www.jstor.org/stable/24640215.

Marshall:2015:WHG

[Mar15b] Mac Marshall. Ward H. Goodenough: 30 May 1919–9 June 2013. Proceedings of the American Philosophical Society, 159(2):207–214, June 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL http://live-amphil.pantheonsite.io/
sites/default/files/2017-10/attachments/Goodenough.pdf;
https://www.jstor.org/stable/24640216.

Marletta:2017:SDN

[Mar17] Michael A. Marletta. Serendipity in discovery: From nitric oxide to Viagra. Proceedings of the American Philosophical Society, 161 (3):189–201, September 2017. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.org/

sites/default/files/2017-12/attachments/Marletta.pdf.

Markmann:2018:LTS

[Mar18] James F. Markmann. The legacy of Tom Starzl is alive and well in transplantation today. Proceedings of the American Philosophical Society, 162(2):139–154, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.

org/sites/default/files/2018-11/attachments/Markmann.pdf.

Mastroianni:2000:EHM

[Mas00] Luigi Mastroianni. Emily Hartshorne Mudd, 6 September 1898–2 May 1998. Proceedings of the American Philosophical Society, 144(1):99– 104, March 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1515609.

Massey:2007:UAI

[Mas07] Douglas S. Massey. Understanding America’s immigration “crisis”. Proceedings of the American Philosophical Society, 151(3):309–327, September 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4599074.

Massey:2016:MUB

[Mas16] Douglas S. Massey. The Mexico–US border in the American imagination. Proceedings of the American Philosophical Society, 160(2): 160–177, June 2016. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Massey.pdf.

Matlock:2001:DSL

[Mat01] Jack F. Matlock. Dmitri Sergeyevich Likhachev, 28 November 1906– 30 September 1999. Proceedings of the American Philosophical Society, 145(3):349–353, September 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558113.

Matlock:2007:GFK

[Mat07a] Jack F. Matlock, Jr. George F. Kennan, 16 February 1904–17 March 2005. Proceedings of the American Philosophical Society, 151(2):233– 242, June 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4599061.

Matlock:2007:IAE

[Mat07b] Jack F. Matlock, Jr. Introduction: American empire? The role of the United States in the world today. Proceedings of the American Philosophical Society, 151(4):403, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/25478451.

Mattson:2009:ASJ

[Mat09] Kevin Mattson. Arthur Schlesinger Jr. 15 October 1917–28 February 2007. Proceedings of the American Philosophical Society, 153(1):115– 126, March 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541636.

Matlock:2010:WEO

[Mat10] Jack F. Matlock, Jr. William Eldridge Odom: 23 June 1932– 30 May 2008. Proceedings of the American Philosophical Society, 154(3):355–358, September 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/41000088.

Maxey:2007:LHV

[Max07] David W. Maxey. Lewis H. Van Dusen Jr., 18 December 1910–16 November 2004. Proceedings of the American Philosophical Society, 151(1):131–138, March 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599052.

Mayr:2001:PFD

[May01] Ernst Mayr. The philosophical foundations of Darwinism. Proceedings of the American Philosophical Society, 145(4):488–495, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558188.

Mayr:2002:GLS

[May02] Ernst Mayr. G. Ledyard Stebbins, 6 January 1906–19 January 2000. Proceedings of the American Philosophical Society, 146(1):129–131, March 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558162.

May:2009:RNJ

[May09] Ernest R. May. Richard E. Neustadt 26 June 1919–31 October 2003. Proceedings of the American Philosophical Society, 153(1):103–114, March 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/40541635.

Mayhew:2017:RDD

[May17] David Mayhew. Robert A. Dahl 17 December 1915–5 February 2014. Proceedings of the American Philosophical Society, 161(4):343– 347, December 2017. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2018-03/attachments/Dahl.pdf.

Meinwald:2007:ISD

[MB07] Jerrold Meinwald and John Brauman. Introduction to symposium on discovery and invention in contemporary chemistry. Proceedings of the American Philosophical Society, 151(2):184, June 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/4599056.

McAuliffe:2003:PPQ

[McA03] Jane Dammen McAuliffe. The persistent power of the Qur’ān. Proceedings of the American Philosophical Society, 147(4):339–346, December 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558297.

McCredie:2000:AFS

[McC00] James R. McCredie. Alison Frantz, 27 September 1903–1 February 1995. Proceedings of the American Philosophical Society, 144(2):213– 217, June 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1515634.

McCredie:2002:HAT

[McC02] James R. McCredie. Homer Armstrong Thompson: 7 September 1906–7 May 2000. Proceedings of the American Philosophical Society, 146(4):411–414, December 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558316.

Mcclellan:2005:ALS

[Mcc05] James E. Mcclellan III. Accident, luck, and serendipity in historical research. Proceedings of the American Philosophical Society, 149(1): 1–21, March 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598905.

McCredie:2014:EBH

[McC14] James R. McCredie. Evelyn Byrd Harrison: 5 June 1920–3 November 2012. Proceedings of the American Philosophical Society, 158(1): 63–67, March 2014. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Harrison.pdf;
 https://www.jstor.org/stable/24640144.

McCormick:2016:JWB

[McC16] Michael McCormick. John W. Baldwin 13 July 1929–8 February 2015. Proceedings of the American Philosophical Society, 160(4):449– 455, December 2016. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-07/attachments/Baldwin.pdf.

McElheny:2002:EHL

[McE02] Victor K. McElheny. Edwin Herbert Land, 7 May 1909–1 March 1991. Proceedings of the American Philosophical Society, 146(1):111–122, March 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558160.

McKusick:2000:MHJ

[McK00] Victor A. McKusick. A. McGehee Harvey, 30 July 1911–8 May 1998. Proceedings of the American Philosophical Society, 144(1):85– 94, March 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1515607.

McKenzie:2001:GDS

[McK01] Ian F. C. McKenzie. George D. Snell, 19 December 1903–6 June 1996. Proceedings of the American Philosophical Society, 145(3):369–372, September 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558117.

McKusick:2004:IRS

[McK04] Vincent L. McKusick. Introductory remarks: Symposium on the fiftieth anniversary of Brown v. Board of Education. Proceedings of the American Philosophical Society, 148(4):427, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1558137.

Machen:2013:RDK

[ME13] Terry Machen and David Erlij. Richard Darwin Keynes: 14 August 1919–12 June 2010. Proceedings of the American Philosophical Society, 157(1):115–120, March 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558144.

Megivern:2003:CPC

[Meg03] James J. Megivern. Capital punishment: The curious history of its privileged place in Christendom. Proceedings of the American Philosophical Society, 147(1):3–12, March 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/1558123.

Meikle:2019:AWC

[Mei19] Jeffrey L. Meikle. Alfred W. Crosby. Proceedings of the American Philosophical Society, 163(1):??, March 2019. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2019-
11/attachments/Meikle.pdf.

Mendelson:2017:WOW

[Men17] Edward Mendelson. Why Odysseus was right about persons. Proceedings of the American Philosophical Society, 161(2):132–137, June 2017. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Mendelson.pdf.

Mazur:2010:AGN

[MGM10] Barry Mazur, Benedict Gross, and David Mumford. Andrew Gleason: 4 November 1921–17 October 2008. Proceedings of the American Philosophical Society, 154(4):471–476, December 2010. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/23056865.

Middlekauff:2006:MTH

[Mid06] Robert L. Middlekauff. Mark Twain’s humor: With examples. Proceedings of the American Philosophical Society, 150(3):449–454, September 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4599008.

Middlekauff:2011:JTA

[Mid11] Robert Middlekauff. James Thorpe: 17 August 1915–4 January 2009. Proceedings of the American Philosophical Society, 155(2):235–239, June 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23056840.

Middlekauff:2015:ESM

[Mid15] Robert Middlekauff. Edmund Sears Morgan: 17 January 1916–8 July 2013. Proceedings of the American Philosophical Society, 159(2): 221–225, June 2015. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Morgan.pdf;
https://www.jstor.org/stable/24640218.

Miller:2006:JIM

[Mil06] Will Miller. Joseph Irwin Miller, 26 May 1909–16 August 2004. Proceedings of the American Philosophical Society, 150(3):493–497, September 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4599015.

Miller:2015:AIC

[Mil15] Robert J. Miller. American Indian constitutions and their influence on the United States Constitution. Proceedings of the American Philosophical Society, 159(1):32–56, March 2015.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic).  URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Miller.pdf;   https://
www.jstor.org/stable/24640169.

Miller:2016:PPC

[Mil16] David Philip Miller. Of patents, principles, and the construction of heroic invention: The case of Neilson’s hot blast in iron production. Proceedings of the American Philosophical Society, 160(4):361– 422, December 2016. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-07/attachments/Miller_1.pdf.

Mintz:2020:DBD

[Min20] Steven Mintz. David Brion Davis. Proceedings of the American Philosophical Society, 164(2):??, ???? 2020. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

amphilsoc.org/sites/default/files/2022-06/attachments/Davis.
pdf.

Mislow:2000:VPJ

[Mis00] Kurt Mislow. Vladimir Prelog, 23 July 1906–7 January 1998. Proceedings of the American Philosophical Society, 144(1):105–111, March 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/1515610.

Mitchell:2001:RRU

[Mit01] Thomas N. Mitchell. Roman republicanism: The underrated legacy. Proceedings of the American Philosophical Society, 145(2):127–137, June 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558267.

Mitchell:2007:EKD

[Mit07] John Mitchell. Ernst Kitzinger, 27 December 1912–22 January 2003. Proceedings of the American Philosophical Society, 151(3):345–350, September 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4599077.

Mittler:2008:PPA

[Mit08] Barbara Mittler. Popular propaganda? Art and culture in Revolutionary China. Proceedings of the American Philosophical Society, 152(4):466–489, December 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541604.

Miyashita:2011:FBL

[Miy11] Mizuki Miyashita. Five Blackfoot lullabies. Proceedings of the American Philosophical Society, 155(3):276–293, September 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/23208685.

McPherson:2011:BBJ

[MLT11] Mary Patterson McPherson, W. Thomas London, and Mark Thompson. Baruch S. Blumberg: 28 July 1925–5 April 2011. Proceedings of the American Philosophical Society, 155(3):301–311, September 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/23208687.

Meinwald:2011:TRO

[MN11] Jerrold Meinwald and Koji Nakanishi. Thomas Risley Odhiambo: 4 February 1931–20 May 2003. Proceedings of the American Philosophical Society, 155(2):231–234, June 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/23056839.

Mariano:2020:LRK

[MO20] Roberto S. Mariano and Suleyman Ozmucur. Lawrence R. Klein. Proceedings of the American Philosophical Society, 164(1):??, March 2020. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2021-12/attachments/Klein.pdf.

Moffett:2020:SIB

[Mof20] Mark W. Moffett. Societies, identity, and belonging. Proceedings of the American Philosophical Society, 164(2):??, ???? 2020. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2022-06/attachments/Moffett.pdf.

Monfasani:2001:POK

[Mon01] John Monfasani. Paul Oskar Kristeller, 22 May 1905–7 June 1999. Proceedings of the American Philosophical Society, 145(2):207–211, June 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558275.

Moncure:2008:EGJ

[Mon08] Sue Swyers Moncure. Edward Graham Jefferson, 15 July 1921– 9 February 2006. Proceedings of the American Philosophical Society, 152(2):243–248, June 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478493.

Morgan:2006:DJB

[Mor06] Edmund S. Morgan. Daniel J. Boorstin, 1 October 1914–28 February 2004. Proceedings of the American Philosophical Society, 150(2):347– 351, June 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598995.

Morrison:2010:IEP

[Mor10] David Morrison. Impacts and evolution: Protecting Earth from asteroids. Proceedings of the American Philosophical Society, 154 (4):439–450, December 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056862.

Morison:2013:SBL

[Mor13] Ian Morison. Sir Bernard Lovell: 31 August 1913–6 August 2012. Proceedings of the American Philosophical Society, 157(3):361–364, September 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/24640401.

Morange:2016:FJJ

[Mor16] Michel Morange. François Jacob 17 June 1920–19 April 2013. Proceedings of the American Philosophical Society, 160(2):213–217, June 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Jacob.pdf.

Mosteller:2005:JWT

[Mos05] Frederick Mosteller. John W. Tukey, 16 June 1915–26 July 2000. Proceedings of the American Philosophical Society, 149(4):625–630, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598966.

Moynihan:2001:CBA

[Moy01] Daniel Patrick Moynihan. Charles Blitzer, 10 August 1927–19 February 1999. Proceedings of the American Philosophical Society, 145(2):189–191, June 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558271.

McDaniel:2002:RRW

[MS02] Boyce McDaniel and Albert Silverman. Robert Rathbun Wilson, 4 March 1914–16 January 2000. Proceedings of the American Philosophical Society, 146(2):229–234, June 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/1558207.

Meserve:2007:PHA

[MS07] Richard A. Meserve and Maxine F. Singer. Philip Hauge Abelson, 27 April 1913–1 August 2004. Proceedings of the American Philosophical Society, 151(4):427–434, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/25478455.

Miller:2012:JCH

[MT12] Paul F. Miller, Jr. and Thomas R. Tritton. John C. Haas: 22 May 1918–2 April 2011. Proceedings of the American Philosophical Society, 156(4):457–462, December 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558239.

Mulford:2008:BFS

[Mul08] Carla Mulford. Benjamin Franklin’s savage eloquence: Hoaxes from the press at Passy, 1782. Proceedings of the American Philosophical Society, 152(4):490–530, December 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/40541605.

Mullen:2010:EWN

[Mul10] Richard Mullen. Esmond Wright: 5 November 1915–9 August 2003. Proceedings of the American Philosophical Society, 154(3):391–397, September 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/41000091.

Mumford:2008:GMF

[Mum08] David Mumford. George Mackey 1 February 1916–15 March 2006. Proceedings of the American Philosophical Society, 152(4):559–563, December 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541610.

Munk:2013:PS

[Mun13] Walter Munk. The perfect storm. Proceedings of the American Philosophical Society, 157(4):369–392, December 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/24640223.

Murtagh:2007:JBJ

[Mur07] William J. Murtagh. James Biddle, 8 July 1929–10 March 2005. Proceedings of the American Philosophical Society, 151(4):441–444, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/25478457.

Mysen:2005:HYJ

[Mys05] Bjorn Mysen. Hatten S. Yoder Jr., 20 March 1921–2 August 2003. Proceedings of the American Philosophical Society, 149(4):639–642, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598968.

Napier:2005:PEJ

[Nap05] Susan J. Napier. The problem of existence in Japanese animation. Proceedings of the American Philosophical Society, 149(1):72–79, March 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4598910.

Nash:2006:FS

[Nas06] Gary B. Nash. Franklin and slavery. Proceedings of the American Philosophical Society, 150(4):618–635, December 2006. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/4599029.

Nasar:2013:GPA

[Nas13] Sylvia Nasar. The grand pursuit of Alfred Marshall and Joseph Schumpeter: The firm, the entrepreneur, and economic growth. Proceedings of the American Philosophical Society, 157(1):58–66, March 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/23558140.

Nash:2019:WWW

[Nas19] Gary B. Nash. When we were young: The American Philosophical Society in the 18th Century. Proceedings of the American Philosophical Society, 163(1):??, March 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.org/

sites/default/files/2019-11/attachments/Nash.pdf.

Nathan:2007:SJK

[Nat07] David G. Nathan. Stanley J. Korsmeyer, 8 June 1950–31 March 2005. Proceedings of the American Philosophical Society, 151(2):243–246, June 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4599062.

Necipoglu:2012:OGN

[Nec12] Gülru Necipoğlu. Oleg Grabar: 3 November 1929–8 January 2011. Proceedings of the American Philosophical Society, 156(4):451–456, December 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23558238.

Nelson:2011:MWN

[Nel11] Phillip G. Nelson. Marshall Warren Nirenberg: 10 April 1927–15 January 2010. Proceedings of the American Philosophical Society, 155(3):367–375, September 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23208694.

Nichols:2005:RAB

[Nic05] Ashton Nichols. Roaring alligators and burning tygers: Poetry and science from William Bartram to Charles Darwin. Proceedings of the American Philosophical Society, 149(3):304–315, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/4598937.

Nierenberg:2000:ASN

[Nie00] William A. Nierenberg. Athelstan Spilhaus, 25 November 1911– 30 March 1998. Proceedings of the American Philosophical Society, 144(3):343–347, September 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515599.

Nolan–Jr:2020:JCB

[NJ20] Michael W. Nolan-Jr. John C. Bogle. Proceedings of the American Philosophical Society, 164(2):??, ???? 2020. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2022-
06/attachments/Bogle.pdf.

Naylor:2005:UGG

[NM05] Rosamond Naylor and Richard Manning. Unleashing the genius of the genome to feed the developing world. Proceedings of the American Philosophical Society, 149(4):515–528, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/4598958.

Norrby:2002:CNP

[Nor02] Erling Norrby. A century of Nobel Prizes. Proceedings of the American Philosophical Society, 146(4):323–336, December 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1558309.

Norcini:2005:PPF

[Nor05] Marilyn Norcini. The political process of factionalism and selfgovernance at Santa Clara Pueblo, New Mexico. Proceedings of the American Philosophical Society, 149(4):544–590, December 2005. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/4598960.

North:2012:MOJ

[Nor12] Helen F. North. Martin Ostwald: 15 January 1922–10 April 2010. Proceedings of the American Philosophical Society, 156(2):251–259, June 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23558198.

Nottebohm:2016:PMF

[Not16] Fernando Nottebohm. Peter Marler 24 February 1928–5 July 2014. Proceedings of the American Philosophical Society, 160(4):437–441, December 2016. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-07/attachments/Marler.pdf.

Nuland:2014:BMR

[Nul14] Sherwin B. Nuland. Books and men, redux. Proceedings of the American Philosophical Society, 158(1):1–9, March 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic).  URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Nuland.pdf;   https://
www.jstor.org/stable/24640139.

Odom:2007:AHH

[Odo07] William E. Odom. American hegemony: How to use it, how to lose it. Proceedings of the American Philosophical Society, 151(4):404–411, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/25478452.

Okimoto:2012:RWJ

[Oki12] Daniel I. Okimoto. Robert E. Ward: 29 January 1916–7 December 2009. Proceedings of the American Philosophical Society, 156(1):105– 108, March 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23558085.

Olopade:2015:JDR

[Olo15] Olufunmilayo Falusi Olopade. Janet Davison Rowley: 5 April 1925–17 December 2013. Proceedings of the American Philosophical Society, 159(3):355–359, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Rowley.pdf;   https://www.jstor.org/stable/26159163.

Olszewski:2019:DET

[Ols19] Edward Olszewski. Dionysus’s enigmatic Thyrsus. Proceedings of the American Philosophical Society, 163(2):??, June 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2020-03/attachments/Olszewski.pdf.

OMalley:2012:CTM

[O’M12] John W. O’Malley. The Council of Trent (1545–63) and Michelangelo’s “Last Judgment” (1541). Proceedings of the American Philosophical Society, 156(4):388–397, December 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/23558232.

Orcutt:2019:WHM

[Orc19] John Orcutt. Walter Heinrich Munk. Proceedings of the American Philosophical Society, 163(3):??, September 2019. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2020-
08/attachments/Munk.pdf.

Ortiz:2004:SRF

[Ort04] Sutti Ortiz. Sir Raymond Firth, 25 March 1901–22 February 2002. Proceedings of the American Philosophical Society, 148(1):129–133, March 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558247.

OShaughnessy:2014:GHS

[O’S14] Andrew Jackson O’Shaughnessy. “To gain the hearts and subdue the minds of America”: General Sir Henry Clinton and the conduct of the British War for America. Proceedings of the American Philosophical Society, 158(3):199–208, September 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL http://live-amphil.pantheonsite.io/
sites/default/files/2017-10/attachments/OShaughnessy.pdf;
 https://www.jstor.org/stable/24640125.

Osmond:2003:SRN

[Osm03] Barry Osmond. Sir Rutherford Ness (Bob) Robertson, 29 September 1913–5 March 2001. Proceedings of the American Philosophical Society, 147(2):181–185, June 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558262.

Ostriker:2006:LSJ

[Ost06] Jeremiah P. Ostriker. Lyman Spitzer Jr., 26 June 1914–31 March 1997. Proceedings of the American Philosophical Society, 150(3): 505–516, September 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599017.

Ostwald:2007:MHJ

[Ost07] Martin Ostwald. Michael Hamilton Jameson, 15 October 1924– 18 August 2004. Proceedings of the American Philosophical Society, 151(1):113–123, March 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599050.

Otvos:2000:MCA

[OTC00] John W. Otvos, Marilyn Taylor, and Robert E. Connick. Melvin Calvin, 8 April 1911–8 January 1997. Proceedings of the American Philosophical Society, 144(4):453–457, December 2000. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/1515620.

Ozdogan:2017:HCA

[Özd17] Mehmet Özdoğan. Halet Çambel 27 August 1916–12 January 2014. Proceedings of the American Philosophical Society, 161(3):277–283, September 2017. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-12/attachments/Cambel.pdf.

Ozyar:2008:MJM

[Özy08] Asli Özyar. Machteld Johanna Mellink 26 October 1917–24 February 2006. Proceedings of the American Philosophical Society, 152 (3):409–417, September 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541596.

Palmer:2003:BHJ

[Pal03] Vaughn Palmer. Bruce Hutchison, 5 June 1901–14 September 1992. Proceedings of the American Philosophical Society, 147(1):83–93, March 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558131.

Palese:2019:LRU

[Pal19] Peter Palese. The long road to a universal influenza virus vaccine. Proceedings of the American Philosophical Society, 163(4):??, December 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2021-01/attachments/Palese.pdf.

Papathomas:2005:BJF

[Pap05] Thomas V. Papathomas. Bela Julesz, 19 February 1928–31 December 2003. Proceedings of the American Philosophical Society, 149(4):599–603, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598962.

Parker:2006:JAS

[Par06] Eugene N. Parker. John Alexander Simpson, 3 November 1916–31 August 2000. Proceedings of the American Philosophical Society, 150(3):499–503, September 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599016.

Parsons:2012:SHL

[Par12] Peter Parsons. Sir Hugh Lloyd-Jones: 21 September 1922–5 October 2009. Proceedings of the American Philosophical Society, 156(1):95– 98, March 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23558083.

Parkinson:2017:SCC

[Par17] Claire L. Parkinson. Satellite contributions to climate change studies. Proceedings of the American Philosophical Society, 161(3):208– 225, September 2017. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-12/attachments/Parkinson.pdf.

Patrick:2009:GAJ

[Pat09] Hugh Patrick. Gardner Ackley 30 June 1915–12 February 1998. Proceedings of the American Philosophical Society, 153(1):69–73, March 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/40541630.

Peachey:2001:RHA

[Pea01] Lee D. Peachey. Richard Hume Adrian, Second Baron Adrian of Cambridge, 16 October 1927–4 April 1995. Proceedings of the American Philosophical Society, 145(1):83–96, March 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1558327.

Peachey:2006:KRP

[Pea06] Lee D. Peachey. Keith Roberts Porter, 11 June 1912–2 May 1997. Proceedings of the American Philosophical Society, 150(4):685–696, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4599036.

Peachey:2014:SAF

[Pea14] Lee D. Peachey. Sir Andrew Fielding Huxley: 22 November 1917–30 May 2012. Proceedings of the American Philosophical Society, 158 (1):69–83, March 2014. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Huxley.pdf;
https://www.jstor.org/stable/24640145.

Peck:2018:RNE

[Pec18] Robert McCracken Peck. The remarkable nature of Edward Lear. Proceedings of the American Philosophical Society, 162(2):158–190, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2018-11/attachments/Peck.pdf.

Pei:2004:JCB

[Pei04] I. M. Pei. J. Carter Brown, 8 October 1934–17 June 2002. Proceedings of the American Philosophical Society, 148(3):369–373, September 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558232.

Pelikan:2003:HAO

[Pel03] Jaroslav J. Pelikan. Heiko Augustinus Oberman, 15 October 1930– 22 April 2001. Proceedings of the American Philosophical Society, 147(3):293–296, September 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558218.

Pelikan:2006:RMF

[Pel06] Jaroslav J. Pelikan. Roland Mushat Frye, 3 July 1921–13 January 2005. Proceedings of the American Philosophical Society, 150(2):361– 365, June 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598997.

Peterson:2001:AMJ

[Pet01] Peter G. Peterson. Akio Morita, 26 January 1921–3 October 1999. Proceedings of the American Philosophical Society, 145(2):213–220, June 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558276.

Pettit:2009:EFS

[Pet09] Donald Pettit. Exploring the frontier: Science of opportunity on the international space station. Proceedings of the American Philosophical Society, 153(4):381–402, December 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/20721509.

Phillips:2005:LPL

[Phi05] David Phillips. The Lord Porter of Luddenham, 6 December 1920– 31 August 2002. Proceedings of the American Philosophical Society, 149(1):107–111, March 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598916.

Phipps:2007:JKN

[Phi07] Charles Phipps. Jack S. Kilby, 8 November 1923–20 June 2005. Proceedings of the American Philosophical Society, 151(4):461–466, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/25478461.

Pierce:2002:NHO

[Pie02] William S. Pierce. New hearts for old: The mechanical option. Proceedings of the American Philosophical Society, 146(2):181–192, June 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/1558201.

Piel:2003:DRB

[Pie03a] Eleanor Jackson Piel. The death row brothers. Proceedings of the American Philosophical Society, 147(1):30–38, March 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1558126.

Piel:2003:IRS

[Pie03b] Gerard Piel. Introductory remarks: Symposium on capital punishment and the administration of justice. Proceedings of the American Philosophical Society, 147(1):1–2, March 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1558122.

Pildes:2018:PFV

[Pil18] Richard H. Pildes. The past and future of voting rights. Proceedings of the American Philosophical Society, 162(3):221–230, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2019-03/attachments/Pildes.pdf.

Pingree:2001:DHH

[Pin01] David Pingree. Daniel Henry Holmes Ingalls, 4 May 1916–17 July 1999. Proceedings of the American Philosophical Society, 145(2):203– 206, June 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558274.

Pines:2009:JBM

[Pin09] David Pines. John Bardeen: 23 May 1908–30 January 1991. Proceedings of the American Philosophical Society, 153(3):287–321, September 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/40541673.

Pinch:2016:TPH

[Pin16] Trevor Pinch. Thomas Parke Hughes 13 September 1923–3 February 2014. Proceedings of the American Philosophical Society, 160(3):295– 301, September 2016. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-07/attachments/Hughes.pdf.

Pippin:2017:PHV

[Pip17] Robert Pippin. The philosophical Hitchcock: Vertigo and the anxieties of unknowingness. Proceedings of the American Philosophical Society, 161(4):293–298, December 2017. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

amphilsoc.org/sites/default/files/2018-03/attachments/Pippin.
pdf.

Plotkin:2009:RAA

[Plo09] Stanley A. Plotkin. Robert Austrian 12 April 1916–17 January 2007. Proceedings of the American Philosophical Society, 153(2):217–220, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/40541660.

Plotkin:2018:THT

[Plo18] Stanley A. Plotkin. Two hundred twenty years of vaccination. Proceedings of the American Philosophical Society, 162(3):295–302, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2019-03/attachments/Plotkin.pdf.

Putnam:2013:HFN

[PO13] Michael C. J. Putnam and John O’Malley. Helen F. North: 31 January 1921–21 January 2012. Proceedings of the American Philosophical Society, 157(2):255–260, June 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/24640245.

Pocock:2001:UA

[Poc01] J. G. A. Pocock. The uniqueness of Aotearoa. Proceedings of the American Philosophical Society, 145(4):482–487, December 2001. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1558187.

Pollak:2004:MVM

[Pol04a] Louis H. Pollak. Marbury v. Madison: What did John Marshall decide and why? Proceedings of the American Philosophical Society, 148(1): 1–13, March 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558240.

Pollak:2004:MRS

[Pol04b] Louis H. Pollak. Moderator’s remarks: Symposium on the fiftieth anniversary of Brown v. Board of Education. Proceedings of the American Philosophical Society, 148(4):428–433, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1558138.

Pomeranz:2008:CDL

[Pom08] Kenneth Pomeranz. Chinese development in long-run perspective. Proceedings of the American Philosophical Society, 152(1):83–100, March 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/25478470.

Portes:2011:AIG

[Por11] Alejandro Portes. America and its immigrants: A game of mirrors. Proceedings of the American Philosophical Society, 155(4):418–432, December 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23208783.

Post:2012:SBJ

[Pos12] Robert C. Post. Silvio A. Bedini: 17 January 1917–14 November 2007. Proceedings of the American Philosophical Society, 156(4):425–433, December 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23558235.

Powell:2014:APS

[Pow14] Timothy B. Powell. The American Philosophical Society: Protocols for the treatment of indigenous materials. Proceedings of the American Philosophical Society, 158(4):411–420, December 2014. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Protocol.pdf;   https://www.jstor.org/stable/
24640185.

Powell:2015:SCS

[Pow15] Timothy B. Powell. Steering a course set by Thomas Jefferson: New developments in the Native American Collections at the American Philosophical Society. Proceedings of the American Philosophical Society, 159(3):282–295, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Powell.pdf;   https://www.jstor.org/stable/26159137.

Poggio:2014:DAG

[PP14] Tomaso Poggio and Allegra Poggio. Donald Arthur Glaser: 21 September 1926–28 February 2013. Proceedings of the American Philosophical Society, 158(3):311–315, September 2014. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Glaser.pdf;   https://www.jstor.org/stable/24640134.

Preston:2000:AAO

[Pre00] Richard S. Preston. The accuracy of the astronomical observations of Lewis and Clark. Proceedings of the American Philosophical Society, 144(2):168–191, June 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515630.

Preston:2005:AJC

[Pre05] Samuel H. Preston. Ansley Johnson Coale, 14 November 1917–5 November 2002. Proceedings of the American Philosophical Society, 149(2):249–253, June 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598929.

Prena:2013:TSW

[Pre13] Jens Prena. Thomas Say’s weevils, with special reference to his “curculionites” pamphlet and the extant type material. Proceedings of the American Philosophical Society, 157(4):393–408, December 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/24640224.

Price:2007:DJM

[Pri07] Raymond Price. Digby Johns McLaren, 11 December 1919–8 December 2004. Proceedings of the American Philosophical Society, 151(3):371–377, September 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599080.

Prown:2007:PMJ

[Pro07] Jules David Prown. Paul Mellon, 11 June 1907–1 February 1999. Proceedings of the American Philosophical Society, 151(3):379–384, September 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4599081.

Probstein:2012:HGS

[Pro12] Ronald F. Probstein. H. Guyford Stever: 24 October 1916–9 April 2010. Proceedings of the American Philosophical Society, 156(2):267– 270, June 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23558200.

Peters:2004:SLP

[PRP^{+}04] Alan Peters, Jack Rosenbluth, George D. Pappas, Lawrence Kruger, and Enrico Mugnaini. Sanford Louis Palay, 23 September 1918– 5 August 2002. Proceedings of the American Philosophical Society, 148(4):519–528, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558146.

Peters:2013:SCC

[PT13] Keith Peters and E. M. Tansey. Sir Christopher Charles Booth: 22 June 1924–13 July 2012. Proceedings of the American Philosophical Society, 157(4):459–463, December 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/24640227.

Putnam:2001:AAV

[Put01] Michael C. J. Putnam. The ambiguity of art in Virgil’s “Aeneid”. Proceedings of the American Philosophical Society, 145(2):162–183, June 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558269.

Putnam:2005:JRF

[Put05] Hilary Putnam. John Rawls, 21 February 1921–24 November 2002. Proceedings of the American Philosophical Society, 149(1):113–117, March 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4598917.

Putnam:2009:RRO

[Put09] Hilary Putnam. Richard Rorty 4 October 1931–8 June 2007. Proceedings of the American Philosophical Society, 153(2):229–232, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/40541663.

Perry:2001:NPG

[PW01] Alfred M. Perry and Alvin M. Weinberg. Nuclear power and the greenhouse effect. Proceedings of the American Philosophical Society, 145(2):184–186, June 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558270.

Pye:2001:MWM

[Pye01] Lucian W. Pye. Myron Weiner, 11 March 1931–3 June 1999. Proceedings of the American Philosophical Society, 145(2):227–231, June 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/1558278.

Pyenson:2002:UCH

[Pye02] Lewis Pyenson. Uses of cultural history: Karl Lamprecht in Argentina. Proceedings of the American Philosophical Society, 146 (3):235–255, September 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558166.

Pye:2005:GAJ

[Pye05] Lucian W. Pye. Gabriel A. Almond, 12 January 1911–25 December 2002. Proceedings of the American Philosophical Society, 149(2):229– 232, June 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598926.

Quadir:2013:ASE

[Qua13] Iqbal Z. Quadir. Adam Smith, economic development, and the global spread of cell phones. Proceedings of the American Philosophical Society, 157(1):67–91, March 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558141.

Ragai:2013:SDF

[Rag13] Jehane Ragai. The scientific detection of forgery in paintings. Proceedings of the American Philosophical Society, 157(2):164–175, June 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/24640239.

Raible:2009:KBN

[Rai09] Wolfgang Raible. Kurt Baldinger 17 November 1919–17 January 2007. Proceedings of the American Philosophical Society, 153(2):221–224, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/40541661.

Rakove:2016:CDF

[Rak16] Jack Rakove. Carl N. Degler 6 February 1921–27 December 2014. Proceedings of the American Philosophical Society, 160(3):303–308, September 2016. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-07/attachments/Degler.pdf.

Rakove:2019:DTA

[Rak19] Jack N. Rakove. Democracy today: Ancient lessons, modern challenges — introduction. Proceedings of the American Philosophical Society, 163(2):??, June 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.org/

sites/default/files/2020-03/attachments/Rakove.pdf.

Ramsey:2004:PKJ

[Ram04] Norman F. Ramsey. Polykarp Kusch, 26 January 1911–20 March 1993. Proceedings of the American Philosophical Society, 148(1):135– 137, March 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558248.

Ramsey:2008:DRS

[Ram08] John T. Ramsey. David Roy Shackleton Bailey, 10 December 1917– 28 November 2005. Proceedings of the American Philosophical Society, 152(2):267–278, June 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478497.

Rudenstine:2012:CRO

[RB12] Neil L. Rudenstine and Nicolas Barker. Charles A. Ryskamp: 21 October 1928–26 March 2010. Proceedings of the American Philosophical Society, 156(3):357–362, September 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/23558098.

Radcliffe:2000:SJP

[RC00] Anthony Radcliffe and Keith Christiansen. Sir John Pope-Hennessy, 13 December 1913–31 October 1994. Proceedings of the American Philosophical Society, 144(4):485–493, December 2000. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/1515625.

Resnik:2007:JLR

[RC07] Judith Resnik and Dennis E. Curtis. The Jayne Lecture: Representing justice: From Renaissance iconography to Twenty-First-Century courthouses. Proceedings of the American Philosophical Society, 151(2):139–183, June 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599055.

Rebert:2011:CSU

[Reb11] Paula Rebert. A civilian surveyor on the United States–Mexico boundary: The case of Arthur Schott. Proceedings of the American Philosophical Society, 155(4):433–462, December 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/23208784.

Redford:2008:MHE

[Red08] Bruce Redford. Mary Hyde Eccles, 12 July 1912–26 August 2003. Proceedings of the American Philosophical Society, 152(2):239–241, June 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/25478492.

Rees:2001:DSN

[Ree01] Martin Rees. Dennis Sciama, 18 November 1926–18 December 1999. Proceedings of the American Philosophical Society, 145(3):365–368, September 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558116.

Rees:2019:STB

[Ree19] Martin Rees. Some thoughts on 2050 and beyond. Proceedings of the American Philosophical Society, 163(4):??, December 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2021-01/attachments/Rees.pdf.

Reiner:2002:HGG

[Rei02] Erica Reiner. Hans Gustav Güterbock, 27 May 1908–29 March 2000. Proceedings of the American Philosophical Society, 146(3):291–296, September 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558170.

Reid:2007:IAI

[Rei07] Gerald F. Reid. Illegal alien? The immigration case of Mohawk ironworker Paul K. Diabo. Proceedings of the American Philosophical Society, 151(1):61–78, March 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599044.

Reid:2015:MCB

[Rei15] Lady Joan Reid. Mary, Countess of Bessborough: 9 March 1915– 13 April 2013. Proceedings of the American Philosophical Society, 159(1):107–111, March 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Bessborough.pdf;   https://www.jstor.org/stable/24640174.

Reich:2018:ADS

[Rei18] David Reich. Ancient DNA suggests steppe migrations spread Indo-European languages. Proceedings of the American Philosophical Society, 162(1):41–57, March 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.org/

sites/default/files/2018-08/attachments/Reich.pdf.

Remer:2006:AAB

[Rem06] Rosalind Remer. The art of the anniversary: Ben Franklin at three hundred. Proceedings of the American Philosophical Society, 150(4):636–642, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599030.

Resnik:2016:WJM

[Res16] Judith Resnik. “Within its jurisdiction”: Moving boundaries, people, and the law of migration. Proceedings of the American Philosophical Society, 160(2):117–159, June 2016. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Resnik.pdf.

Rhodes:2009:WGO

[Rho09] Frank H. T. Rhodes. William T. Golden, 25 October 1909–7 October 2007. Proceedings of the American Philosophical Society, 153(4):457–466, December 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/20721514.

Rice:2008:MS

[Ric08] John A. Rice. Mozart on the stage. Proceedings of the American Philosophical Society, 152(2):179–188, June 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/25478485.

Rice:2012:HCT

[Ric12] Charles M. Rice. Hepatitis c: Toward a cure and eradication. Proceedings of the American Philosophical Society, 156(3):324–330, September 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23558093.

Richardson:2015:JTJ

[Ric15] R. C. Richardson. Joan Thirsk: 19 June 1922–3 October 2013. Proceedings of the American Philosophical Society, 159(3):343–348, September 2015. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Thirsk.pdf;
https://www.jstor.org/stable/26159161.

Ridgway:2004:PPB

[Rid04] Brunilde Sismondo Ridgway. Phyllis Pray Bober. 2 December 1920– 30 May 2002. Proceedings of the American Philosophical Society, 148(3):361–367, September 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558231.

Ridgway:2005:SGS

[Rid05] Brunilde Sismondo Ridgway. The study of Greek sculpture in the Twenty-First Century. Proceedings of the American Philosophical Society, 149(1):63–71, March 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598909.

Ridgway:2011:ERK

[Rid11] Brunilde Sismondo Ridgway. Elfriede Regina Knauer: 3 July 1926– 7 June 2010. Proceedings of the American Philosophical Society, 155(3):327–334, September 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23208690.

Rigden:2000:IIR

[Rig00] John S. Rigden. Isidor Isaac Rabi, 29 July 1898–11 January 1988. Proceedings of the American Philosophical Society, 144(1):113–118, March 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1515611.

Rimel:2001:CSP

[Rim01] Rebecca W. Rimel. Charity and strategy: Philanthropy’s evolving role. Proceedings of the American Philosophical Society, 145 (4):587–595, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558196.

Rimel:2008:TWL

[Rim08] Rebecca W. Rimel. Thomas W. Langfitt, 20 April 1927–7 August 2005. Proceedings of the American Philosophical Society, 152(1):151– 155, March 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/25478478.

Ristau:2005:DRG

[Ris05] Carolyn A. Ristau. Donald Redfield Griffin, 3 August 1915–7 November 2003. Proceedings of the American Philosophical Society, 149 (3):399–411, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598944.

Richard:2009:WPJ

[RK09] John P. Richard and Jack F. Kirsch. William Platt Jencks 15 August 1927–3 January 2007. Proceedings of the American Philosophical Society, 153(1):97–101, March 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541634.

Roehl:2008:CMC

[Roe08] Richard W. Roehl. Carlo M. Cipolla, 15 August 1922–5 September 2000. Proceedings of the American Philosophical Society, 152(2):235– 238, June 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/25478491.

Romano:2006:RSC

[Rom06] David Gilman Romano. Roman surveyors in Corinth. Proceedings of the American Philosophical Society, 150(1):62–85, March 2006. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/4598973.

Ronda:2006:AWK

[Ron06] James P. Ronda. “To acquire what knolege you can”: Thomas Jefferson as exploration patron and planner. Proceedings of the American Philosophical Society, 150(3):409–413, September 2006. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/4599004.

Roos:2015:DCG

[Roo15] Raymond P. Roos. D. Carleton Gajdusek: 9 September 1923–12 December 2008. Proceedings of the American Philosophical Society, 159 (1):97–106, March 2015. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Gajdusek.pdf;
 https://www.jstor.org/stable/24640173.

Rostow:2000:MJF

[Ros00] W. W. Rostow. Modern Japan’s fourth challenge: The political economy of a stagnant population. Proceedings of the American Philosophical Society, 144(4):384–396, December 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1515618.

Rose:2006:DDJ

[Ros06a] Kenneth Rose. The 11th Duke of Devonshire, 2 January 1920– 3 May 2004. Proceedings of the American Philosophical Society, 150(3):475–478, September 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599012.

Rosenberg:2006:FF

[Ros06b] Pierre Rosenberg. Franklin and Fragonard. Proceedings of the American Philosophical Society, 150(4):575–590, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/4599026.

Rosen:2007:SBP

[Ros07] Jeffrey Rosen. The silver bullet: Protecting privacy and security through law and technology. Proceedings of the American Philosophical Society, 151(3):291–299, September 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/4599072.

Roth:2009:ERA

[Rot09] Martha Roth. Erica Reiner: 4 August 1924–31 December 2005. Proceedings of the American Philosophical Society, 153(3):365–369, September 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541679.

Rotberg:2011:HSN

[Rot11] Robert I. Rotberg. Helen Suzman: 17 November 1917–1 January 2009. Proceedings of the American Philosophical Society, 155(1):115– 119, March 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23056854.

Rothblatt:2012:BAB

[Rot12] Martine Rothblatt. Borders are not always boundaries: Are genetically engineered XenoLungs from pigs feasible before the end of this decade? Proceedings of the American Philosophical Society, 156(4):379–387, December 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558231.

Rubin:2020:DPS

[Rub20] David A. Rubin. Design problem-solving in the Anthropocene. Proceedings of the American Philosophical Society, 164(1):??, March 2020. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2021-12/attachments/Rubin.pdf.

Ruderman:2005:GJF

[Rud05a] Malvin A. Ruderman. Gordon James Fraser MacDonald, 30 July 1929–14 May 2002. Proceedings of the American Philosophical Society, 149(3):427–429, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598947.

Rudwick:2005:PNA

[Rud05b] Martin Rudwick. Picturing nature in the age of enlightenment. Proceedings of the American Philosophical Society, 149(3):279–303, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598936.

Rudenstine:2014:ABH

[Rud14] Neil L. Rudenstine. Albert Barnes and his foundation: Three paradoxes. Proceedings of the American Philosophical Society, 158(2):120– 127, June 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL http://live-amphil.pantheonsite.io/
sites/default/files/2017-10/attachments/Rudenstine.pdf;
https://www.jstor.org/stable/24640199.

Ruff:2006:JRJ

[Ruf06] Willie Ruff. John Rodgers, 11 July 1914–7 March 2004. Proceedings of the American Philosophical Society, 150(2):397–402, June 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/4599000.

Russell:2016:ALH

[Rus16] James S. Russell. Ada Louise Huxtable 14 March 1921–7 January 2013. Proceedings of the American Philosophical Society, 160(1): 97–101, March 2016. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Huxtable.pdf.

Ruttan:2005:STC

[Rut05] Vernon W. Ruttan. Scientific and technical constraints on agriculture production: Prospects for the future. Proceedings of the American Philosophical Society, 149(4):453–468, December 2005. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/4598955.

Rutishauser:2015:GEJ

[Rut15] Urs Rutishauser. Gerald Edelman: 1 July 1929–17 May 2014. Proceedings of the American Philosophical Society, 159(4):461–466, December 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Edelman.pdf;   https://
www.jstor.org/stable/26159198.

Singer:2009:JFM

[SA09] Maxine Singer and Sidney Altman. Joseph S. Fruton, 14 May 1912– 29 July 2007. Proceedings of the American Philosophical Society, 153(4):447–456, December 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/20721513.

Saaty:2016:SMN

[Saa16] Thomas Saaty. Seven is the magic number in nature. Proceedings of the American Philosophical Society, 160(4):335–360, December 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2017-07/attachments/Saaty.pdf.

Sabloff:2004:GRW

[Sab04] Jeremy A. Sabloff. Gordon Randolph Willey, 7 March 1913–28 April 2002. Proceedings of the American Philosophical Society, 148 (3):405–410, September 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558237.

Sabloff:2007:IDH

[Sab07] Jeremy A. Sabloff. It depends on how we look at things: New perspectives on the postclassic period in the Northern Maya Lowlands. Proceedings of the American Philosophical Society, 151(1):11–26, March 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/4599041.

Sakmar:2017:HGK

[Sak17] Thomas P. Sakmar. Har Gobind Khorana 9 January 1922–9 November 2011. Proceedings of the American Philosophical Society, 161(3):267– 275, September 2017. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-12/attachments/Khorana.pdf.

Salpeter:2006:TGM

[Sal06] Edwin E. Salpeter. Thomas Gold, 22 May 1920–22 June 2004. Proceedings of the American Philosophical Society, 150(3):479–486, September 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4599013.

Sandage:2006:HWB

[San06] Allan Sandage. Horace Welcome Babcock, 13 September 1912– 29 August 2003. Proceedings of the American Philosophical Society, 150(1):151–160, March 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598977.

Satava:2003:BEM

[Sat03] Richard M. Satava. Biomedical, ethical, and moral issues being forced by advanced medical technologies. Proceedings of the American Philosophical Society, 147(3):246–258, September 2003. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/1558213.

Starzl:2011:STO

[SB11] Thomas E. Starzl and Clyde Barker. The shared trail of organ, limb, and face transplantation. Proceedings of the American Philosophical Society, 155(1):1–12, March 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056844.

Scirica:2014:LHP

[SBC^{+}14] Anthony J. Scirica, Stephen B. Burbank, Guido Calabresi, Stewart Dalzell, Michael A. Fitts, Alfred W. Putnam, Jr., and Catherine Struve. Louis H. Pollak: 7 December 1922–8 May 2012. Proceedings of the American Philosophical Society, 158(1):85–99, March 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Pollak.pdf;   https://
www.jstor.org/stable/24640146.

Scadden:2009:SCB

[Sca09] David T. Scadden. Stem cell-based medicine. Proceedings of the American Philosophical Society, 153(2):185–192, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/40541657.

Shapiro:2012:PDA

[SCG^{+}12] Judith R. Shapiro, Jonathan R. Cole, Hanna H. Gray, Neil J. Smelser, and Harriet Zuckerman. Panel discussion: The American Research University in crisis? Proceedings of the American Philosophical Society, 156(2):188–204, June 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558191.

Schwartz:2002:GMP

[Sch02] Seymour I. Schwartz. The greatest misnomer on planet Earth. Proceedings of the American Philosophical Society, 146(3):264–281, September 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558168.

Schiebinger:2005:AEA

[Sch05] Londa Schiebinger. Agnotology and exotic abortifacients: The cultural production of ignorance in the Eighteenth-Century Atlantic world. Proceedings of the American Philosophical Society, 149 (3):316–343, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598938.

Schmidt:2011:DQ

[Sch11] Maarten Schmidt. The discovery of quasars. Proceedings of the American Philosophical Society, 155(2):142–146, June 2011. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/23056831.

Schramm:2013:EGP

[Sch13] Carl J. Schramm. Entrepreneurship and global prosperity: Implications for poverty and peace. Proceedings of the American Philosophical Society, 157(1):92–106, March 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/23558142.

Schmitt:2014:MMS

[Sch14] Rüdiger Schmitt. Manfred Mayrhofer: 26 September 1926–31 October 2011. Proceedings of the American Philosophical Society, 158(3):323– 327, September 2014. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Mayrhofer.pdf;
 https://www.jstor.org/stable/24640136.

Smith:2010:WJB

[SCL^{+}10] Susan Bell Smith, Ellen R. Cohn, Martin L. Levitt, Carl F. Miller, Richard S. Dunn, and Arlin M. Adams. Whitfield J. Bell Jr. 3 December 1914–2 January 2009. Proceedings of the American Philosophical Society, 154(2):261–278, June 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/41000100.

Scott:2006:SJC

[Sco06] John Scott. Sir John Carew Eccles, 27 January 1903–2 May 1997. Proceedings of the American Philosophical Society, 150(4):673–678, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4599034.

Sedley:2007:LGS

[Sed07] Stephen Sedley. Leslie George Scarman, Baron Scarman of Quatt, 29 July 1911–8 December 2004. Proceedings of the American Philosophical Society, 151(1):125–129, March 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/4599051.

Segal:2017:TRT

[Seg17] Nancy L. Segal. Twins reared together and apart: The science behind the fascination. Proceedings of the American Philosophical Society, 161(1):1–17, March 2017. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.org/

sites/default/files/2017-07/attachments/Segal.pdf.

Seitz:2001:JCM

[Sei01] Frederick Seitz. James Clerk Maxwell (1831–1879); member APS 1875. Proceedings of the American Philosophical Society, 145(1):1–44, March 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558323.

Seitz:2002:CNS

[Sei02] Frederick Seitz. China and natural science: A conundrum. In remembrance of Professor Ta You Wu (1907–2000). Proceedings of the American Philosophical Society, 146(1):1–17, March 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1558153.

Seitz:2004:HCC

[Sei04] Frederick Seitz. Henry Cavendish: The catalyst for the chemical revolution: To the memory of Glenn T. Seaborg (1912–1999). Proceedings of the American Philosophical Society, 148(2):151–179, June 2004.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/1558282.

Seitz:2005:PAS

[Sei05] Frederick Seitz. Pierre Aigrain, 28 September 1924–30 October 2002. Proceedings of the American Philosophical Society, 149(1):83– 87, March 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598911.

Seitz:2008:WOB

[Sei08] Frederick Seitz. William Oliver Baker, 15 July 1915–31 October 2005. Proceedings of the American Philosophical Society, 152(1):109–119, March 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/25478472.

Selden:2005:TBB

[Sel05] Steven Selden. Transforming better babies into fitter families: Archival resources and the history of the American eugenics movement, 1908–1930. Proceedings of the American Philosophical Society, 149(2):199–225, June 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598925.

Sela:2011:EKK

[Sel11] Michael Sela. Ephraim Katzir-Katchalski: 16 May 1916–30 May 2009. Proceedings of the American Philosophical Society, 155(2):211–215, June 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23056836.

Szlezak:2018:WBF

[SG18] Thomas Alexander Szlezák and M. Laura Gemelli. Walter Burkert 2 February 1931–11 March 2015. Proceedings of the American Philosophical Society, 162(1):115–124, March 2018. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2018-
08/attachments/Burkert.pdf.

Souter:2010:SCU

[SGTG10] David Souter, Ruth Bader Ginsburg, David S. Tatel, and Linda Greenhouse. The supreme court and useful knowledge: Panel discussion. Proceedings of the American Philosophical Society, 154 (3):294–306, September 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/41000082.

Shavelson:2003:LJC

[Sha03] Richard J. Shavelson. Lee J. Cronbach, 22 April 1916–1 October 2001. Proceedings of the American Philosophical Society, 147(4):379–385, December 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558300.

Shaw:2005:MWC

[Sha05] Gwendolyn DuBois Shaw. “Moses Williams, cutter of profiles”: Silhouettes and African American identity in the Early Republic. Proceedings of the American Philosophical Society, 149(1):22–39, March 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/4598906.

Shepard:2014:GAM

[She14] Roger N. Shepard. George Armitage Miller: 3 February 1920–22 July 2012. Proceedings of the American Philosophical Society, 158(2): 167–172, June 2014. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Miller_0.pdf;
 https://www.jstor.org/stable/24640205.

Shelemay:2015:SS

[She15] Kay Kaufman Shelemay. Songs that speak. Proceedings of the American Philosophical Society, 159(3):296–305, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Shelemay.pdf;   https:/
/www.jstor.org/stable/26159138.

Shultz:2006:JDJ

[Shu06] George P. Shultz. John Dunlop, 5 July 1914–2 October 2003. Proceedings of the American Philosophical Society, 150(1):187–190, March 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/4598982.

Shweder:2010:CGA

[Shw10] Richard A. Shweder. Clifford Geertz, 23 August 1926–30 October 2006. Proceedings of the American Philosophical Society, 154(1):87– 93, March 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/20721529.

Siegel:2014:EIF

[Sie14] Jeremy J. Siegel. Earnings, inflation, and future stock and bond returns. Proceedings of the American Philosophical Society, 158(3):222– 228, September 2014. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Siegel.pdf;
https://www.jstor.org/stable/24640127.

Silverman:2004:MEW

[Sil04] Robert A. Silverman. Marvin Eugene Wolfgang, 14 November 1924– 12 April 1998. Proceedings of the American Philosophical Society, 148(4):547–554, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558150.

Silverstein:2009:JHG

[Sil09] Michael Silverstein. Joseph H. Greenberg: 28 May 1915–7 May 2001. Proceedings of the American Philosophical Society, 153(3):341–349, September 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541676.

Silverstein:2016:FVA

[Sil16] Michael Silverstein. Fabrication, verification, authentication: An introduction. Proceedings of the American Philosophical Society, 160(1): 1–5, March 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL http://live-amphil.pantheonsite.io/
sites/default/files/2017-10/attachments/Silverstein.pdf.

Simons:2009:PIS

[Sim09] Thomas W. Simons, Jr. Pakistan and Islam sixty years further on. Proceedings of the American Philosophical Society, 153(3):279–284, September 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541672.

Simons:2011:FCH

[Sim11] Elwyn L. Simons. Francis Clark Howell: 27 November 1925–10 March 2007. Proceedings of the American Philosophical Society, 155(2):205– 209, June 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23056835.

Simon:2016:MBK

[Sim16] Elaine L. Simon. Michael B. Katz 13 April 1939–23 August 2014. Proceedings of the American Philosophical Society, 160(2):219–225, June 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Katz.pdf.

Singer:2004:JDE

[Sin04] Maxine Singer. James David Ebert, 11 December 1921–22 May 2001. Proceedings of the American Philosophical Society, 148(1):123–127, March 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558246.

Schreiner:2010:HGB

[SK10] Peter Schreiner and Holger A. Klein. Hans-Georg Beck: 18 February 1910–25 May 1999. Proceedings of the American Philosophical Society, 154(4):453–459, December 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056863.

Slichter:2004:HGD

[Sli04] Charles P. Slichter. Harry G. Drickamer, 19 November 1918–6 May 2002. Proceedings of the American Philosophical Society, 148 (3):375–381, September 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558233.

Slichter:2010:FSJ

[Sli10] Charles P. Slichter. Frederick Seitz: 4 July 1911–2 March 2008. Proceedings of the American Philosophical Society, 154(3):359–384, September 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/41000089.

Sloan:2012:WHG

[Slo12] Kay Sloan. William H. Goetzmann: 20 July 1930–7 September 2010. Proceedings of the American Philosophical Society, 156(2):229–235, June 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23558195.

Smelser:2011:RDM

[Sme11] Neil J. Smelser. Ralf Dahrendorf: 1 May 1929–17 June 2009. Proceedings of the American Philosophical Society, 155(4):465–469, December 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23208785.

Smith:2001:WIG

[Smi01] Richard M. Smith. Welfare of the individual and the group: Malthus and externalities. Proceedings of the American Philosophical Society, 145(4):402–414, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558181.

Smith:2004:WHM

[Smi04] A. Mark Smith. What is the history of medieval optics really about? Proceedings of the American Philosophical Society, 148(2):180–194, June 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558283.

Smith:2012:AIM

[Smi12] David Horton Smith. Alex Inkeles: 4 March 1920–9 July 2010. Proceedings of the American Philosophical Society, 156(2):237–243, June 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/23558196.

Snyder:2005:APS

[Sny05] Solomon H. Snyder. The audacity principle in science. Proceedings of the American Philosophical Society, 149(2):141–158, June 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/4598923.

Snyder:2007:JAM

[Sny07] Solomon H. Snyder. Julius Axelrod, 30 May 1912–29 December 2004. Proceedings of the American Philosophical Society, 151(1):81– 90, March 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4599045.

Sokoloff:2002:SKA

[Sok02] Louis Sokoloff. Seymour S. Kety, 25 August 1915–25 May 2000. Proceedings of the American Philosophical Society, 146(1):101–110, March 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558159.

Solow:2004:JTM

[Sol04] Robert M. Solow. James Tobin, 5 March 1918–11 March 2002. Proceedings of the American Philosophical Society, 148(3):399–404, September 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558236.

Sorenson:2015:JMB

[Sor15] Georgia Sorenson. James Macgregor Burns: 3 August 1918–15 July 2014. Proceedings of the American Philosophical Society, 159 (4):475–481, December 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Burns.pdf;   https://www.jstor.org/stable/26159200.

Spande:2012:BWM

[Spa12] Thomas F. Spande. Bernhard Witkop: 9 May 1917–22 November 2010. Proceedings of the American Philosophical Society, 156 (4):463–473, December 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23558240.

Spero:2018:OPT

[Spe18] Patrick Spero. The other presidency: Thomas Jefferson and the 3 American Philosophical Society. Proceedings of the American Philosophical Society, 162(4):321–360, ???? 2018. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2019-
06/attachments/Spero.pdf.

Suttkus:2003:SIV

[SPF03] Royal D. Suttkus, Brady A. Porter, and Byron J. Freeman. The status and infraspecific variation of Notropis stonei Fowler. Proceedings of the American Philosophical Society, 147(4):354–376, December 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/1558299.

Straus:2000:JRD

[SR00] Donald B. Straus and David Rockefeller. J. Richardson Dil Worth, 9 June 1916–29 December 1997. Proceedings of the American Philosophical Society, 144(3):301–307, September 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/1515593.

Susskind:2018:FP

[SS18] Daniel Susskind and Richard Susskind. The future of the professions. Proceedings of the American Philosophical Society, 162(2):125–138, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.amphilsoc.org/sites/default/
files/2018-11/attachments/Susskind%20and%20Susskind.pdf.

Starzl:2003:OTP

[Sta03] Thomas E. Starzl. Organ transplantation: A practical triumph and epistemologic collapse. Proceedings of the American Philosophical Society, 147(3):226–245, September 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/1558212.

Starzl:2005:FDM

[Sta05] Thomas E. Starzl. Francis D. Moore, 17 August 1913–24 November 2001. Proceedings of the American Philosophical Society, 149 (4):605–613, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598963.

Stallybrass:2006:BFP

[Sta06] Peter Stallybrass. Benjamin Franklin: Printed corrections and erasable writing. Proceedings of the American Philosophical Society, 150(4):553–567, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599024.

Starna:2008:ROL

[Sta08] William A. Starna. Retrospecting the origins of the League of the Iroquois. Proceedings of the American Philosophical Society, 152(3):279–321, September 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541589.

Stevenson:2003:CPE

[Ste03] Bryan Stevenson. Crime, punishment, and executions in the twentyfirst century. Proceedings of the American Philosophical Society, 147(1):24–29, March 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558125.

Steinhardt:2004:EUB

[Ste04a] Paul J. Steinhardt. The endless universe: A brief introduction. Proceedings of the American Philosophical Society, 148(4):464–470, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558141.

Stent:2004:PFW

[Ste04b] Gunther S. Stent. Paradoxes of free will and the limits of human reason. Proceedings of the American Philosophical Society, 148(2):205– 212, June 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558285.

Steward:2005:FJH

[Ste05] James Christen Steward. Francis James Herbert Haskell, 7 April 1928–18 January 2000. Proceedings of the American Philosophical Society, 149(4):593–597, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/4598961.

Stent:2006:FCJ

[Ste06] Gunther S. Stent. Francis Crick, 8 June 1916–28 July 2004. Proceedings of the American Philosophical Society, 150(3):467–474, September 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4599011.

Stern:2008:GCN

[Ste08] Fritz Stern. Gordon A. Craig, 26 November 1913–30 October 2005. Proceedings of the American Philosophical Society, 152(1):129–134, March 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/25478474.

Stedman:2009:RTM

[Ste09] Hansell Stedman. Room for thought: Of mutations, muscle, “fitness,” and the human condition. Proceedings of the American Philosophical Society, 153(3):271–278, September 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/40541671.

Sternberg:2011:HCR

[Ste11] Jonathan Sternberg. H. C. Robbins Landon: 6 March 1926–20 November 2009. Proceedings of the American Philosophical Society, 155(4):471–475, December 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23208786.

Stowe:2006:WMA

[Sto06] William W. Stowe. Writing Mount Auburn: Language, landscape, and place. Proceedings of the American Philosophical Society, 150(2):296–317, June 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598991.

Stoner:2011:AWJ

[Sto11] Joyce Hill Stoner. Andrew Wyeth: 12 July 1917–16 January 2009. Proceedings of the American Philosophical Society, 155(1):121–128, March 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23056855.

Stout:2014:CL

[Sto14] Lynn Stout. The corporation and the law. Proceedings of the American Philosophical Society, 158(4):364–371, December 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Stout.pdf;   https://www.
jstor.org/stable/24640181.

Storchmann:2015:EOW

[Sto15] Karl Storchmann. Expert opinion on the wine market. Proceedings of the American Philosophical Society, 159(2):129–140, June 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL http://live-amphil.pantheonsite.io/
sites/default/files/2017-10/attachments/Storchmann.pdf;
https://www.jstor.org/stable/24640210.

Sudarshan:2000:REM

[Sud00] E. C. G. Sudarshan. Robert Eugene Marshak, 11 October 1916–23 December 1992. Proceedings of the American Philosophical Society, 144(3):319–321, September 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515596.

Sugerman:2011:BSS

[Sug11] Harvey J. Sugerman. Bariatric surgery for severe obesity. Proceedings of the American Philosophical Society, 155(3):263–275, September 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23208684.

Suppes:2006:DDM

[Sup06] Patrick Suppes. Donald Davidson, 6 March 1917–31 August 2003. Proceedings of the American Philosophical Society, 150(2):353–359, June 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4598996.

Suppes:2014:RDL

[Sup14] Patrick Suppes. R. Duncan Luce: 16 May 1925–11 August 2012. Proceedings of the American Philosophical Society, 158 (4):441–448, December 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Luce.pdf;   https://www.jstor.org/stable/24640189.

Sirigu:2011:NHR

[SVD11] Angela Sirigu, Claudia Vargas, and Jean-Michel Dubernard. Novel hand representations in the human motor cortex after hand allograft. Proceedings of the American Philosophical Society, 155(1):29– 40, March 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23056847.

Swerdlow:2004:ETK

[Swe04] N. M. Swerdlow. An essay on Thomas Kuhn’s first scientific revolution, “The Copernican Revolution”. Proceedings of the American Philosophical Society, 148(1):64–120, March 2004. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/1558245.

Sylla:2014:HAC

[Syl14] Richard Sylla. How the American corporation evolved over two centuries. Proceedings of the American Philosophical Society, 158 (4):354–363, December 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Sylla.pdf;   https://www.jstor.org/stable/24640180.

Tabak:2003:EUI

[Tab03] Ronald J. Tabak. The egregiously unfair implementation of capital punishment in the United States: “super due process” or super lack of due process? Proceedings of the American Philosophical Society, 147(1):13–23, March 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558124.

Talbott:2006:HFB

[Tal06] Page Talbott. On how to fit Benjamin Franklin into eight thousand square feet. Proceedings of the American Philosophical Society, 150(4):643–646, December 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599031.

Tatel:2010:IR

[Tat10] David S. Tatel. Introductory remarks. Proceedings of the American Philosophical Society, 154(3):281–286, September 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/41000080.

Tatel:2015:RV

[Tat15] David S. Tatel. The right to vote. Proceedings of the American Philosophical Society, 159(1):1–4, March 2015. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Tatel.pdf;   https://www.jstor.org/stable/24640166.

Tattersall:2016:TPE

[Tat16] Ian Tattersall. The thinking primate: Establishing a context for the emergence of modern human cognition. Proceedings of the American Philosophical Society, 160(3):254–265, September 2016. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2017-07/attachments/Tattersall.pdf.

Tatel:2019:SPS

[Tat19] David S. Tatel. Separation of powers and statutory interpretation: a battle hidden in plain sight. Proceedings of the American Philosophical Society, 163(3):??, September 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.

org/sites/default/files/2020-08/attachments/Tatel.pdf.

Taubenberger:2006:OVS

[Tau06] Jeffery K. Taubenberger. The origin and virulence of the 1918 “Spanish” influenza virus. Proceedings of the American Philosophical Society, 150(1):86–112, March 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598974.

Taubenberger:2019:SII

[Tau19] Jeffery K. Taubenberger. Scientific insights into the 1918 influenza pandemic and their implications for the future. Proceedings of the American Philosophical Society, 163(4):312–324, December 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.amphilsoc.org/sites/default/files/
2021-01/attachments/Taubenberger.pdf.

Telegdi:2002:GOD

[Tel02] Valentine L. Telegdi. Giuseppe Occhialini, 5 December 1907–30 December 1993. Proceedings of the American Philosophical Society, 146(2):217–222, June 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558205.

Terzian:2010:ESD

[Ter10] Yervant Terzian. Edwin E. Salpeter: 3 December 1924–26 November 2008. Proceedings of the American Philosophical Society, 154 (4):499–502, December 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056869.

Thatcher:2017:HBJ

[Tha17] Sanford G. Thatcher. Herbert S. Bailey, Jr. 12 July 1921–28 June 2011. Proceedings of the American Philosophical Society, 161(1): 69–83, March 2017. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-07/attachments/Bailey.pdf.

Thomas:2002:HIJ

[Tho02] John Meurig Thomas. Hiroshi Inose, 5 January 1927–11 October 2000. Proceedings of the American Philosophical Society, 146(2):211–215, June 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558204.

Thompson:2003:SRT

[Tho03] James C. Thompson. Surgical research: Trepan to transplant. Proceedings of the American Philosophical Society, 147(3):204–225, September 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558211.

Thomas:2004:MPM

[Tho04] John Meurig Thomas. Max Perutz, 19 May 1914–6 February 2002. Proceedings of the American Philosophical Society, 148(2):235–241, June 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558289.

Thomas:2006:FF

[Tho06] John Meurig Thomas. Faraday and Franklin. Proceedings of the American Philosophical Society, 150(4):523–541, December 2006. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/4599022.

Thompson:2007:GCM

[Tho07] Kenneth W. Thompson. George C. McGhee, 10 March 1912– 4 July 2005. Proceedings of the American Philosophical Society, 151(4):467–472, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478462.

Thomas:2008:NJL

[Tho08] John Meurig Thomas. Nelson J. Leonard 1 September 1916–9 October 2006. Proceedings of the American Philosophical Society, 152(4):553–557, December 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541609.

Thompson:2009:DMF

[Tho09] Mark Thompson. Digital media and the future of quality broadcasting. Proceedings of the American Philosophical Society, 153(2):140– 146, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541654.

Thompson:2010:UGC

[Tho10] Lonnie G. Thompson. Understanding global climate change: Paleoclimate perspective from the world’s highest mountains. Proceedings of the American Philosophical Society, 154(2):133–157, June 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/41000095.

Thomson:2011:GCS

[Tho11] Keith Thomson. The “great-claw” and the science of Thomas Jefferson. Proceedings of the American Philosophical Society, 155 (4):394–403, December 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23208781.

Thomas:2013:SHD

[Tho13] John Meurig Thomas. Sir Humphry Davy: Natural philosopher, discoverer, inventor, poet, and man of action. Proceedings of the American Philosophical Society, 157(2):143–163, June 2013. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/24640238.

Thompson:2017:PPF

[Tho17] Lonnie G. Thompson. Past, present, and future of glacier archives from the world’s highest mountains. Proceedings of the American Philosophical Society, 161(3):226–243, September 2017. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2017-
12/attachments/Thompson.pdf.

Thomas:2018:AHZ

[Tho18] John Meurig Thomas. Ahmed H. Zewail 26 February 1946–2 August 2016. Proceedings of the American Philosophical Society, 162(2): 193–206, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2018-11/attachments/Zewail.pdf.

Tipton:2015:RBF

[Tip15] Steven M. Tipton. Robert Bellah: 23 February 1927–30 July 2013. Proceedings of the American Philosophical Society, 159(2):215–220, June 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Bellah.pdf;   https://
www.jstor.org/stable/24640217.

Tomlinson:2015:JKA

[Tom15] Gary Tomlinson. Joseph Kerman: 3 April 1924–17 March 2014. Proceedings of the American Philosophical Society, 159(4):455–459, December 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Kerman.pdf;   https://
www.jstor.org/stable/26159197.

Tracy:2000:DAI

[Tra00] Stephen Tracy. Dating Athenian inscriptions: A new approach. Proceedings of the American Philosophical Society, 144(1):67–76, March 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/1515605.

Tremblay:2009:AHL

[Tre09] Marc-Adélard Tremblay. Alexander Hamilton Leighton, 17 July 1908– 11 August 2007. Proceedings of the American Philosophical Society, 153(4):477–488, December 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/20721516.

Treuer:2019:IVL

[Tre19] Anton Treuer. Indigenous voice: Language and identity in America’s Native Nations. Proceedings of the American Philosophical Society, 163(3):??, September 2019. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.org/

sites/default/files/2020-08/attachments/Treuer.pdf.

Trimble:2005:JGO

[Tri05] Virginia Trimble. Jesse L. Greenstein, 15 October 1909–21 October 2002. Proceedings of the American Philosophical Society, 149(1):93– 98, March 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598913.

Tribe:2014:RDD

[Tri14] Laurence H. Tribe. Ronald Dworkin: 11 December 1931–14 February 2013. Proceedings of the American Philosophical Society, 158(2): 155–160, June 2014. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Dworkin.pdf;
https://www.jstor.org/stable/24640203.

Tribe:2017:RNB

[Tri17] Laurence H. Tribe. Reflections on the “natural born citizen” clause as illuminated by the Cruz candidacy. Proceedings of the American Philosophical Society, 161(2):111–131, June 2017. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Tribe.pdf.

Thomas:2014:JHS

[TS14] Sir John Meurig Thomas and Charles Pence Slichter. John H. Sinfelt: 18 February 1931–28 May 2011. Proceedings of the American Philosophical Society, 158(3):305–309, September 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Sinfelt_0.pdf;   https:
//www.jstor.org/stable/24640133.

Tsipis:2007:PMN

[Tsi07] Kosta Tsipis. Philip Morrison, 7 November 1915–22 April 2005. Proceedings of the American Philosophical Society, 151(4):473–477, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/25478463.

Ulrich:2020:WSW

[Ulr20] Laurel Thatcher Ulrich. Woman suffrage West to East: The view from Puck, 1915. Proceedings of the American Philosophical Society, 164(1):??, March 2020. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2021-12/attachments/Ulrich.pdf.

Vacanti:2007:JLT

[Vac07] Joseph P. Vacanti. The Jayne Lecture: Tissue engineering and regenerative medicine. Proceedings of the American Philosophical Society, 151(4):395–402, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478450.

Vacanti:2009:MJF

[Vac09] Joseph P. Vacanti. M. Judah Folkman, 24 February 1933–14 January 2008. Proceedings of the American Philosophical Society, 153 (4):443–446, December 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/20721512.

Vagelos:2001:SBS

[Vag01] P. Roy Vagelos. Social benefits of a successful biomedical research company: Merck. Proceedings of the American Philosophical Society, 145(4):575–578, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558194.

VanEekelen:2007:EVA

[Van07] Willem Frederik Van Eekelen. A European view of American foreign policy. Proceedings of the American Philosophical Society, 151(4):412–418, December 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478453.

Vansina:2011:PDC

[Van11] Jan Vansina. Philip D. Curtin: 22 May 1922–4 June 2009. Proceedings of the American Philosophical Society, 155(3):313–317, September 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/23208688.

VanHorne:2018:TCS

[Van18] John C. Van Horne. Two chips off the same block: Benjamin Franklin’s Library Company and Philosophical Society and the saga of their 275-year relationship. Proceedings of the American Philosophical Society, 162(4):361–386, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.

org/sites/default/files/2019-06/attachments/Van%20Horne.
pdf.

Varnedoe:2003:WMA

[Var03] Kirk Varnedoe. Why modern art matters now. Proceedings of the American Philosophical Society, 147(2):128–133, June 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1558254.

Varsano:2007:EP

[Var07] Paula Varsano. Enduring pleasures. Proceedings of the American Philosophical Society, 151(2):211–217, June 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/4599059.

Varmus:2009:SPS

[Var09] Harold Varmus. Science publishing and science libraries in the Internet Age. Proceedings of the American Philosophical Society, 153(2):160–184, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541656.

Vendler:2001:MVL

[Ven01] Helen Vendler. Matter on various length scales: Poetry. Proceedings of the American Philosophical Society, 145(4):389–401, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/1558180.

Vendler:2015:SJH

[Ven15] Helen Vendler. Seamus Justin Heaney: 13 April 1939–30 August 2013. Proceedings of the American Philosophical Society, 159(2): 233–240, June 2015. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Heaney.pdf;
https://www.jstor.org/stable/24640220.

Vita-Finzi:2006:IS

[VF06] Claudio Vita-Finzi. The inconstant Sun. Proceedings of the American Philosophical Society, 150(1):113–120, March 2006. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/4598975.

Vita-Finzi:2008:LBL

[VF08] Claudio Vita-Finzi. Luna Bergere Leopold, 8 October 1915–23 February 2006. Proceedings of the American Philosophical Society, 152(2):249–253, June 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/25478494.

Vita-Finzi:2010:IAW

[VF10] Claudio Vita-Finzi. Introduction: Action not words. Proceedings of the American Philosophical Society, 154(2):131–132, June 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/41000094.

Vita-Finzi:2015:PSS

[VF15] Claudio Vita-Finzi. The perfect solar storm. Proceedings of the American Philosophical Society, 159(4):415–420, December 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL http://live-amphil.pantheonsite.io/
sites/default/files/2017-10/attachments/Vita_Finzi.pdf;
https://www.jstor.org/stable/26159194.

Villarreal:2004:CVM

[Vil04] Luis P. Villarreal. Can viruses make us human? Proceedings of the American Philosophical Society, 148(3):296–323, September 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/1558228.

Vircillofranklin:2002:PCD

[Vir02] Carmela Vircillofranklin. “Pro communi doctorum virorum comodo”: The Vatican Library and its service to scholarship. Proceedings of the American Philosophical Society, 146(4):363–384, December 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/1558312.

Vogt:2014:RBO

[Vog14] Peter K. Vogt. Richard Bing: 12 October 1909–8 November 2010. Proceedings of the American Philosophical Society, 158 (3):287–291, September 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Bing.pdf;   https://www.jstor.org/stable/24640131.

VonValtier:2011:EAD

[Von11] William F. Von Valtier. “An extravagant assumption”: The demographic numbers behind Benjamin Franklin’s twenty-five-year doubling period. Proceedings of the American Philosophical Society, 155(2):158–188, June 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056833.

Verba:2011:LWP

[VV11] Sidney Verba and Ezra F. Vogel. Lucian W. Pye: 21 October 1921– 5 September 2008. Proceedings of the American Philosophical Society, 155(1):101–108, March 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056852.

Vest:2012:HWJ

[VW12] Charles M. Vest and Kathryn A. Willmore. Howard Wesley Johnson: 2 July 1922–12 December 2009. Proceedings of the American Philosophical Society, 156(1):89–94, March 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/23558082.

Walker:2002:HBW

[Wal02] George E. Walker. Herman B Wells, 7 June 1902–18 March 2000. Proceedings of the American Philosophical Society, 146(2):223–228, June 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558206.

Wald:2006:ICC

[Wal06] Patricia M. Wald. International criminal courts: Some kudos and concerns. Proceedings of the American Philosophical Society, 150(2):241– 260, June 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598989.

Walsh:2007:RHA

[Wal07] Christopher T. Walsh. Robert Heinz Abeles, 14 January 1926– 18 June 2000. Proceedings of the American Philosophical Society, 151(3):331–335, September 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599075.

Wallace:2009:BMM

[Wal09] Daniel B. Wallace. Bruce Manning Metzger, 9 February 1914–13 February 2007. Proceedings of the American Philosophical Society, 153(4):489–493, December 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/20721517.

Walker:2013:ELO

[Wal13] James M. Walker. Elinor “Lin” Ostrom: 7 August 1933–12 June 2012. Proceedings of the American Philosophical Society, 157(3):365–368, September 2013. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/24640402.

Warner:2005:TNW

[War05] Deborah Jean Warner. True North: And why it mattered in Eighteenth-Century America. Proceedings of the American Philosophical Society, 149(3):372–385, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/4598941.

Warnock:2012:SIC

[War12] John E. Warnock. Simple ideas that changed printing and publishing. Proceedings of the American Philosophical Society, 156(4):363–378, December 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23558230.

Warren:2015:A

[War15a] Rosanna Warren. Aftermath. Proceedings of the American Philosophical Society, 159(3):317, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/26159147.

Warren:2015:Ib

[War15b] Rosanna Warren. As if. Proceedings of the American Philosophical Society, 159(3):324, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/26159153.

Warren:2015:C

[War15c] Rosanna Warren. Cassandra. Proceedings of the American Philosophical Society, 159(3):311, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/26159142.

Warren:2015:F

[War15d] Rosanna Warren. Fire. Proceedings of the American Philosophical Society, 159(3):319, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/26159149.

Warren:2015:Gb

[War15e] Rosanna Warren. Glaucoma. Proceedings of the American Philosophical Society, 159(3):330–331, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/26159159.

Warren:2015:Ga

[War15f] Rosanna Warren. Graffiti. Proceedings of the American Philosophical Society, 159(3):328, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/26159157.

Warren:2015:Ia

[War15g] Rosanna Warren. Introduction. Proceedings of the American Philosophical Society, 159(3):306, September 2015. CODEN PAPCAA.

ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/26159139.

Warren:2015:J

[War15h] Rosanna Warren. July. Proceedings of the American Philosophical Society, 159(3):326, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/26159155.

Warren:2015:MS

[War15i] Rosanna Warren. Man in stream. Proceedings of the American Philosophical Society, 159(3):321, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/26159151.

Warren:2015:MJS

[War15j] Rosanna Warren. Max Jacob at Saint Benoît. Proceedings of the American Philosophical Society, 159(3):307, September 2015. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/26159140.

Warren:2015:Mc

[War15k] Rosanna Warren. Mediterranean. Proceedings of the American Philosophical Society, 159(3):315, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/26159145.

Warren:2015:Md

[War15l] Rosanna Warren. Mistral I. Proceedings of the American Philosophical Society, 159(3):316, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/26159146.

Warren:2015:Mb

[War15m] Rosanna Warren. Moment. Proceedings of the American Philosophical Society, 159(3):314, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/26159144.

Warren:2015:Ma

[War15n] Rosanna Warren. Mud. Proceedings of the American Philosophical Society, 159(3):312–313, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/26159143.

Warren:2015:O

[War15o] Rosanna Warren. Ocular. Proceedings of the American Philosophical Society, 159(3):320, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/26159150.

Warren:2015:PR

[War15p] Rosanna Warren. Poetry reading. Proceedings of the American Philosophical Society, 159(3):??, ???? 2015. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Warren.pdf.

Warren:2015:PP

[War15q] Rosanna Warren. Porta Portese. Proceedings of the American Philosophical Society, 159(3):318, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.

jstor.org/stable/26159148.

Warren:2015:R

[War15r] Rosanna Warren. Rats. Proceedings of the American Philosophical Society, 159(3):325, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/26159154.

Warren:2015:SDF

[War15s] Rosanna Warren. Sans domicile fixe. Proceedings of the American Philosophical Society, 159(3):327, September 2015. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/26159156.

Warren:2015:T

[War15t] Rosanna Warren. Tashlich. Proceedings of the American Philosophical Society, 159(3):322–323, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/26159152.

Warren:2015:TD

[War15u] Rosanna Warren. The twelfth day. Proceedings of the American Philosophical Society, 159(3):308–310, September 2015. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/26159141.

Warren:2015:W

[War15v] Rosanna Warren. A way. Proceedings of the American Philosophical Society, 159(3):329, September 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/26159158.

Washington:2006:CMT

[Was06] Warren M. Washington. Computer modeling the Twentieth- and Twenty-First-Century climate. Proceedings of the American Philosophical Society, 150(3):414–427, September 2006. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/4599005.

Washington:2017:ISO

[Was17] Warren M. Washington. Introduction to the symposium on observed climate change. Proceedings of the American Philosophical Society, 161(3):202–207, September 2017. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.amphilsoc.

org/sites/default/files/2017-12/attachments/Washington.pdf.

Watson:2005:RJB

[Wat05] Patty Jo Watson. Robert John Braidwood, 29 July 1907–15 January 2003. Proceedings of the American Philosophical Society, 149(2):233– 241, June 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4598927.

Weaire:2001:PLF

[Wea01] Denis Weaire. A philomorph looks at foam. Proceedings of the American Philosophical Society, 145(4):564–574, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/1558193.

Weeks:2004:RWP

[Wee04] Theodore R. Weeks. Russification: Word and practice 1863–1914. Proceedings of the American Philosophical Society, 148(4):471–489, December 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558142.

Weiss:2004:CCT

[Wei04] Robin A. Weiss. Circe, Cassandra, and the Trojan pigs: Xenotransplantation. Proceedings of the American Philosophical Society, 148(3):281–295, September 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558227.

Weissman:2006:SCB

[Wei06] Irving L. Weissman. Stem cells: Biology, transplantation, and political ethics. Proceedings of the American Philosophical Society, 150(1):121–147, March 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598976.

West:2011:HQT

[Wes11] Martin West. The homeric question today. Proceedings of the American Philosophical Society, 155(4):383–393, December 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

https://www.jstor.org/stable/23208780.

Whitman:2014:LPR

[WH14] Leanna Lee Whitman and Michael Hayes. Lou Pollak: The road to “Brown v. Board of Education” and beyond. Proceedings of the American Philosophical Society, 158(1):31–60, March 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Whitman.pdf;   https://
www.jstor.org/stable/24640143.

Wharton:2003:TWS

[Wha03] Clifton R. Wharton, Jr. Theodore W. Schultz, 30 April 1902–26 February 1998. Proceedings of the American Philosophical Society, 147(3):311–315, September 2003. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558221.

Whitehead:2000:PUA

[Whi00] John C. Whitehead. Not-for-profits: An underdeveloped asset for America. Proceedings of the American Philosophical Society, 144(1):77–82, March 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1515606.

Wilson:2005:HWA

[Wil05a] David B. Wilson. Harry Woolf, 12 August 1923–6 January 2003. Proceedings of the American Philosophical Society, 149(2):273–278, June 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4598933.

Wilson:2005:LEP

[Wil05b] Edward O. Wilson. The Linnaean enterprise: Past, present, and future. Proceedings of the American Philosophical Society, 149 (3):344–348, September 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598939.

Williamson:2009:ADC

[Wil09] Oliver E. Williamson. Alfred D. Chandler Jr. 15 September 1918– 9 May 2007. Proceedings of the American Philosophical Society, 153(2):225–228, June 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541662.

Wilding:2016:FM

[Wil16] Nick Wilding. Forging the Moon. Proceedings of the American Philosophical Society, 160(1):37–72, March 2016. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:/

/live-amphil.pantheonsite.io/sites/default/files/2017-10/
attachments/Wilding.pdf.

Will:2017:DEG

[Wil17] Clifford M. Will. Did Einstein get it right? A centennial assessment. Proceedings of the American Philosophical Society, 161(1): 18–30, March 2017. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2017-07/attachments/Will.pdf.

Winch:2001:DFA

[Win01] Donald Winch. Darwin fallen among political economists. Proceedings of the American Philosophical Society, 145(4):415–437, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/1558182.

Winston:2009:FHB

[Win09a] Michael R. Winston. Frederick H. Burkhardt 13 September 1912– 23 September 2007. Proceedings of the American Philosophical Society, 153(1):75–81, March 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541631.

Winter:2009:WWP

[Win09b] Irene J. Winter. What/when is a portrait? Royal images of the Ancient Near East. Proceedings of the American Philosophical Society, 153(3):254–270, September 2009. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/40541670.

Winkler:2015:TWW

[Win15] Jonathan Reed Winkler. Telecommunications in World War I. Proceedings of the American Philosophical Society, 159(2):162–168, June 2015. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL http://live-amphil.pantheonsite.io/sites/
default/files/2017-10/attachments/Winkler.pdf;   https://
www.jstor.org/stable/24640213.

Winner:2019:HAW

[Win19] Ellen Winner. How art works: a conversation between philosophy and psychology. Proceedings of the American Philosophical Society, 163(2):??, June 2019. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2020-03/attachments/Winner.pdf.

Witten:2020:MFA

[Wit20] Edward Witten. Michael Francis Atiyah. Proceedings of the American Philosophical Society, 164(1):??, March 2020. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.amphilsoc.org/sites/default/files/2021-
12/attachments/Atiyah.pdf.

Wolfensohn:2000:RFP

[Wol00] James D. Wolfensohn. The Richard A. F. Penrose, Jr., Memorial Lecture: Development choices in a changing world. Proceedings of the American Philosophical Society, 144(3):237–244, September 2000. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/1515587.

Wolff:2001:DGE

[Wol01a] Christoph Wolff. Defining genius: Early reflections of J. S. Bach’s self-image. Proceedings of the American Philosophical Society, 145 (4):474–481, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558186.

Wolynes:2001:LFG

[Wol01b] Peter G. Wolynes. Landscapes, funnels, glasses, and folding: From metaphor to software. Proceedings of the American Philosophical Society, 145(4):555–563, December 2001. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.

org/stable/1558192.

Wolman:2002:HIS

[Wol02] M. Gordon Wolman. The human impact: Some observations. Proceedings of the American Philosophical Society, 146(1):81–98, March 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (elec-

tronic). URL https://www.jstor.org/stable/1558158.

Woloch:2004:RRP

[Wol04] Isser Woloch. Robert R. Palmer, 11 January 1909–11 June 2002. Proceedings of the American Philosophical Society, 148(3):393–398, September 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/1558235.

Wolman:2005:ISI

[Wol05] M. Gordon Wolman. Introduction to symposium on issues in agriculture. Proceedings of the American Philosophical Society, 149 (4):451–452, December 2005. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4598954.

Wolpert:2011:SBO

[Wol11] Lewis Wolpert. Seymour Benzer: 15 October 1921–30 November 2007. Proceedings of the American Philosophical Society, 155(3):297–300, September 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23208686.

Wolf:2014:HMP

[Wol14a] Susan Wolf. Happiness and meaning: A plurality of values rather than a conflict of norms. Proceedings of the American Philosophical Society, 158(1):18–24, March 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Wolf.pdf;   https://www.jstor.org/stable/24640141.

Wolff:2014:WOS

[Wol14b] Edward N. Wolff. Who owns stock in American corporations? Proceedings of the American Philosophical Society, 158 (4):372–391, December 2014. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://live-amphil.

pantheonsite.io/sites/default/files/2017-10/attachments/
Wolff.pdf;   https://www.jstor.org/stable/24640182.

Wood:2006:ESN

[Woo06a] Michael Wood. Edward Said, 1 November 1935–25 September 2003. Proceedings of the American Philosophical Society, 150(1):197–200, March 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4598984.

Wood:2006:FTT

[Woo06b] Michael Wood. Found in translation: Travels between languages. Proceedings of the American Philosophical Society, 150(3):443–448, September 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4599007.

Worley:2018:PFM

[Wor18] Richard Worley. Paul F. Miller, Jr. 19 October 1927–9 September 2017. Proceedings of the American Philosophical Society, 162(4): 389–393, ???? 2018. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2019-06/attachments/Miller.pdf.

Wallace:2015:HBC

[WP15] Anthony F. C. Wallace and Timothy B. Powell. How to buy a continent: The protocol of Indian treaties as developed by Benjamin Franklin and other members of the American Philosophical Society. Proceedings of the American Philosophical Society, 159(3):251– 281, September 2015. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://live-amphil.pantheonsite.
io/sites/default/files/2017-10/attachments/Wallace.pdf;
https://www.jstor.org/stable/26159136.

Wright:2007:WNP

[Wri07] Gavin Wright. William Nelson Parker, 14 June 1919–29 April 2000. Proceedings of the American Philosophical Society, 151(2):267–270, June 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243

(electronic). URL https://www.jstor.org/stable/4599066.

Wu:2007:SSC

[Wu07] Hung-Hsi Wu. Shiing-Shen Chern, 26 October 1911–3 December 2004. Proceedings of the American Philosophical Society, 151(1):97– 101, March 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/4599047.

Wylam:2011:PBM

[Wyl11] William B. Wylam. Paul B. Maccready: 25 September 1925– 28 August 2007. Proceedings of the American Philosophical Society, 155(2):217–224, June 2011. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/23056837.

Wyngaarden:2004:DSF

[Wyn04] James B. Wyngaarden. Donald Sharp Fredrickson, 8 August 1924– 7 June 2002. Proceedings of the American Philosophical Society, 148(3):383–391, September 2004. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558234.

Yakovenko:2012:VIA

[Yak12] Sergei Yakovenko. Vladimir Igorevich Arnold: 12 June 1937–3 June 2010. Proceedings of the American Philosophical Society, 156(1):61– 65, March 2012. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/23558078.

Yoder:2002:GSC

[Yod02] Hatten S. Yoder. Geology: Significant component of new multidisciplinary sciences. Proceedings of the American Philosophical Society, 146(1):37–55, March 2002. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/1558155.

Yu:2007:TCC

[Yu07] Pauline Yu. Travels of a culture: Chinese poetry and the European imagination. Proceedings of the American Philosophical Society, 151(2):218–229, June 2007. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https://www.jstor.org/

stable/4599060.

Zemon:2002:FRM

[Zem02] Vance Zemon. Floyd Ratliff: 1 May 1919–13 June 1999. Proceedings of the American Philosophical Society, 146(4):395–401, December 2002.

CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL https://www.jstor.org/stable/1558314.

Zewail:2006:FV

[Zew06] Ahmed H. Zewail. Franklin’s vision. Proceedings of the American Philosophical Society, 150(4):542–552, December 2006. CODEN PA-PCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL https:

//www.jstor.org/stable/4599023.

Ziolkowski:2020:JND

[Zio20] Jan M. Ziolkowski. The juggler of Notre Dame and the medievalizing of modernity. Proceedings of the American Philosophical Society, 164(2):??, ???? 2020. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL https://www.amphilsoc.org/sites/
default/files/2022-06/attachments/Ziolkowski.pdf.

Zucker:2008:AMW

[Zuc08] Alexander Zucker. Alvin M. Weinberg 20 April 1915–18 October 2006. Proceedings of the American Philosophical Society, 152(4):571–576, December 2008. CODEN PAPCAA. ISSN 0003-049X (print), 2326-

9243 (electronic). URL https://www.jstor.org/stable/40541612.